### v1

# Active Learning for Control-Oriented Identification of Nonlinear Systems

Abstract
Model-based reinforcement learning is an effective approach for controlling an unknown system. It is based on a longstanding pipeline familiar to the control community in which one performs experiments on the environment to collect a dataset, uses the resulting dataset to identify a model of the system, and finally performs control synthesis using the identified model. As interacting with the system may be costly and time consuming, targeted exploration is crucial for developing an effective control-oriented model with minimal experimentation. Motivated by this challenge, recent work has begun to study finite sample data requirements and sample efficient algorithms for the problem of optimal exploration in model-based reinforcement learning. However, existing theory and algorithms are limited to model classes which are linear in the parameters. Our work instead focuses on models with nonlinear parameter dependencies, and presents the first finite sample analysis of an active learning algorithm suitable for a general class of nonlinear dynamics. In certain settings, the excess control cost of our algorithm achieves the optimal rate, up to logarithmic factors. We validate our approach in simulation, showcasing the advantage of active, control-oriented exploration for controlling nonlinear systems.
1 Introduction
In recent years, model-based reinforcement learning has been successfully applied to various application domains including robotics, healthcare, and autonomous driving (Levine et al., 2020; Moerland et al., 2023). These approaches often proceed by performing experiments on a system to collect data, and then using the data to fit models for the dynamics. In the specified application domains, performing experiments requires interaction with the physical world, which can be both costly and time-consuming. It is therefore important to design the experimentation and identification procedures to efficiently extract the most information relevant to control. In particular, experiments must be designed with the downstream control objective in mind. This fact is well-established in classical controls and identification literature (Ljung, 1998; Gevers, 1993; Hjalmarsson et al., 1996; Pukelsheim, 2006). While these works provide some guidance for experiment design, they mostly focus on linear systems, and supply only asymptotic guarantees.
Driven by the empirical success of machine and deep learning in solving classes of complex control problems, the learning and control communities have recently begun revisiting the classical pipeline of identification to control, proposing new algorithms, and analyzing them from a non-asymptotic viewpoint. Early efforts focused on end-to-end control guarantees for unknown linear system under naive exploration (injecting white noise inputs) (Dean et al., 2020; Mania et al., 2019). These methods have also been refined by using active learning to collect better data for control synthesis (Wagenmaker et al., 2021). This approach has been extended to nonlinear systems with a linear dependence on the unknown parameters (Wagenmaker et al., 2023). Other works studying model-based control of nonlinear systems also assume linear dependence on the unknown parameters, or consider related simplifying assumptions in settings including tabular or low-rank Markov Decision Processes (Uehara and Sun, 2021; Song and Sun, 2021). Model-based reinforcement learning for a general class of nonlinear systems has also been considered (Sukhija et al., 2023). However, their guarantees focus on the worst case uncertainty of any control policy rather than end-to-end control costs for a particular objective.
There is a significant gap in that there are no algorithms with strong guarantees (achieving the optimal rates) for model-based reinforcement learning of general nonlinear dynamical systems. We leverage recently developed machinery for non-asymptotic analysis of nonlinear system identification to tackle this problem (Ziemann and Tu, 2022).
1.1 Contribution
We introduce and analyze the Active Learning for Control-Oriented Identification (ALCOI) algorithm. This algorithm extends an approach for model-based reinforcement learning proposed by Wagenmaker et al. (2023) for dynamical systems with a linear dependence on the unknown parameter to general nonlinear dynamics that satisfy some smoothness assumptions. The algorithm is inspired by a reduction of the excess control cost to the system identification error, which may then be controlled using novel finite sample system identification error bounds for smooth nonlinear systems.
Leveraging the aforementioned reduction of the excess control cost and system identification error bounds, we derive finite sample bounds for the excess cost of our algorithm.
Theorem 1.1 (Main Result, Informal).
Let the ALCOI algorithm interact with an unknown nonlinear dynamical system for some number of exploration rounds before proposing a control policy designed to optimize some objective. The excess cost of the proposed policy on the objective satisfies
$\mathsf{excess\,cost}\!\leq\!\frac{\mathsf{hardness\,of\,control}\!\times\!\mathsf{hardness\,of\,identification}}{\mathsf{\#\,exploration\,rounds}}.$
The “hardness of control” captures how the error in estimation of the dynamics translates to error in control, while the “hardness of identification” captures how challenging it is to identify the parameters under the best possible exploration policy. Moreover, our analysis reveals how the respective hardness quantities interact. Wagenmaker et al. (2023) provide upper and lower bounds for this problem in a setting where the dynamics model is linear in the unknown parameters. Our upper bound is tight up to logarithmic factors in this setting, and we conjecture that it is also tight up to logarithmic factors in the setting where the model is nonlinear in the parameters.
The non-asymptotic system identification result may be of independent interest. It derives from invoking recently developed machinery for the analysis of nonlinear system identification along with the delta method, a classical approach from statistics. These bounds provide rates that match the asymptotic limit up to logarithmic factors.
1.2 Related Work
Additional Work Analyzing Identification & Control
Finite sample guarantees for active exploration of pure system identification have been studied in linear (Wagenmaker and Jamieson, 2020), and nonlinear (with linear dependence on the unknown parameters) settings (Mania et al., 2020). Lower bounds complementing the upper bounds for the end-to-end control are also present (Wagenmaker et al., 2021, 2023), and have been specialized to the linear-quadratic regulator setting to characterize systems which are hard to learn to control (Lee et al., 2023). Recent literature considers gradient-based approaches for experiment design in linear-quadratic control (Anderson and Hespanha, 2024). For more details on finite sample analysis of learning to control, see the survey by Tsiamis et al. (2023). The aforementioned results do not focus on general nonlinear systems. Such analysis exists for identification; however, in the absence of end-to-end control error bounds (Sattar and Oymak, 2022; Ziemann and Tu, 2022). In contrast, we achieve end-to-end control error bounds for active learning applied for learning to control general nonlinear systems.
Dual Control
A related paradigm to the “identify then control” scheme studied in this work is that of dual control, in which the learner must interact with an unknown system while simultaneously optimizing a control objective (Feldbaum, 1960). Åström and Wittenmark (1973) study a version of this problem known as the self-tuning regulator, providing asymptotic guarantees of convergence. Non-asymptotic guarantees for the self tuning regulator have been studied more recently from the online learning perspective of regret (Abbasi-Yadkori et al., 2011). Subsequent work provides matching upper and lower bounds for the regret of the self-tuning regulator problem (Simchowitz and Foster, 2020). Lower bounds refining the dependence on system-theoretic constants have also been established Ziemann and Sandberg (2022). The regret of learning to control nonlinear dynamical systems (with linear dependence on the unknown parameter) has also been studied (Kakade et al., 2020; Boffi et al., 2021). As in the “identify then control” setting, prior work in dual control has not provided finite sample analysis of the end-to-end control error for systems with nonlinear dependence on the unknown parameters.
Notation: Expectation (respectively probability) with respect to all the randomness of the underlying probability space is denoted by $\operatorname{\mathbb{E}}$ (respectively $\operatorname{\mathbb{P}}$). The Euclidean norm of a vector $x$ is denoted $\left\|x\right\|$. For a matrix $A$, the spectral norm is denoted $\left\|A\right\|$, and the Frobenius norm is denoted $\left\|A\right\|_{F}$. A symmetric, positive semi-definite matrix $A=A^{\top}$ is denoted $A\succeq 0$. $A\succeq B$ denotes that $A-B$ is positive semi-definite. Similarly, a symmetric, positive definite matrix $A$ is denoted $A\succ 0$. The minimum eigenvalue of a symmetric, positive semi-definite matrix $A$ is denoted $\lambda_{\min}(A)$. For a positive definite matrix $A$, we define the $A$-norm as $\left\|x\right\|_{A}^{2}=x^{\top}Ax$. The gradient of a scalar valued function $f:\mathbb{R}^{n}\to\mathbb{R}$ is denoted $\nabla f$, and the Hessian is denoted $\nabla^{2}f$. The Jacobian of a vector-valued function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is denoted $Dg$, and follows the convention for any $x\in\mathbb{R}^{n}$, the rows of $Dg(x)$ are the transposed gradients of $g_{i}(x)$. The $p^{th}$ order derivative of $g$ is denoted by $D^{p}g$. Note that for $p\geq 2$, $D^{p}g(x)$ is a tensor for any $x\in\mathbb{R}^{n}$. The operator norm of such a tensor is denoted by $\left\|D^{p}g(x)\right\|_{\mathsf{op}}$. For a function $f:\mathsf{X}\to\mathbb{R}^{d_{\mathsf{Y}}}$, we define $\left\|f\right\|_{\infty}\triangleq\sup_{x\in\mathsf{X}}\left\|f(x)\right\|$. A Euclidean norm ball of radius $r$ centered at $x$ is denoted $\mathcal{B}(x,r)$.
2 Problem Formulation
We consider a nonlinear dynamical system evolving according to
$X_{t+1}=f(X_{t},U_{t};\phi^{\star})+W_{t},\quad t=1,\dots T,$ (1)
with state $X_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{X}}}$, input $U_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{U}}}$, and $d_{\mathsf{X}}$-dimensional noise $W_{t}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{w}^{2}I)$. For simplicity, we assume $X_{1}=0$. Here, $f$ is the dynamics function, which depends on an unknown parameter $\phi^{\star}\in\mathbb{R}^{d_{\phi}}$. We assume that there exists some positive $B$ such that $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$
We study a learner whose objective is to determine a policy $\pi=\mathopen{}\left\{\pi_{t}\right\}\mathclose{}_{t=1}^{T}$ from a policy class $\Pi^{\star}$ to minimize the cost $\mathcal{J}(\pi,\phi^{\star})$, where
$\mathcal{J}(\pi,\phi)=\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right].$ (2)
The superscript on the expectation denotes that the dynamics (1) are rolled out under parameter $\phi$, while the subscript denotes that the system is played in closed-loop under the feedback control policy $U_{t}=\pi_{t}(X_{1},U_{1},\dots,X_{t-1},U_{t-1},X_{t})$ for $t=1,\dots,T$. The learner follows a two step interaction protocol with an exploration phase, and an evaluation phase. In the exploration phase, the learner interacts with the system for a total of $N$ episodes, each consisting of $T$ timesteps, by playing exploration policies $\pi\in\Pi_{\exp}$. The policy class $\Pi_{\exp}$ is an exploration policy class, described in more detail below. The learner does not incur any cost during the exploration episodes, and seeks only to gain information about the system. After the $N$ interaction episodes, it uses the collected data to propose a policy $\hat{\pi}\in\Pi^{\star}$. The learner is then evaluated on the expected cost of the proposed policy on a new evaluation episode. In particular, it incurs cost $\mathcal{J}(\hat{\pi},\phi^{\star})$.
The policy classes $\Pi^{\star}$ and $\Pi_{\mathsf{exp}}$ are known; $\Pi^{\star}$ consists of deterministic policies, but $\Pi_{\mathsf{exp}}$ may be random. We do not assume $\Pi^{\star}=\Pi_{\mathsf{exp}}$. We assume that the policy class $\Pi^{\star}$ has the parametric form:
$\Pi^{\star}=\mathopen{}\left\{\pi^{\theta}|\theta\in\mathbb{R}^{d_{\theta}}\right\}\mathclose{}.$
No such parametric assumption is made on the exploration class $\Pi_{\exp}$. It instead consists of whatever experimental procedures are available. For instance, it may be the class of policies with power or energy bounded inputs. Policies belonging to $\Pi_{\exp}$ must be history dependent (causal). We assume that the learner is allowed to randomly select choices of policies in $\Pi_{\exp}$.111i.e. for any $\pi^{1},\pi^{2}\in\Pi_{\exp}$ and any $b\in[0,1]$, the policy $\pi_{\mathsf{mix}}$ which at the start of a new episode plays $\pi^{1}$ for the duration of the episode with probability $b$ and $\pi^{2}$ for the duration of the episode with probability $1-b$ also belongs to $\Pi_{\exp}$. Given these policy classes, the learning procedure should seek to identify the best exploitation policy belonging to $\Pi^{\star}$ by playing the most informative exploration policy in the class $\Pi_{\exp}$.
2.1 Certainty Equivalent Control
Figure 1: Identfication to control pipeline.
We focus on learners which follow a model-based approach to synthesize a control policy from interaction with the system, outlined in Figure 1. In this section, we discuss the learner’s procedure for the last two steps: system identification and control synthesis. In Section 3, we return to the question of which experiments the learner should perform.
Given the data collected during the experimentation phase, the learner finds an estimate for the dynamics by solving a nonlinear least squares problem. In particular, suppose that during the $N$ experimentation episodes of length $T$, the learner collects data $\mathopen{}\left\{(U_{t}^{n},X_{t}^{n},X_{t+1}^{n})\right\}\mathclose{}_{n,t=1}^{N,T+1}$. The subscript denotes the time index within each episode, while the superscript denotes the episode index. Using this dataset, the learner may identify the dynamics of the system by solving
$\hat{\phi}\in\operatorname*{argmin}_{\phi\in\mathbb{R}^{d_{\phi}}}\sum_{n=1}^{N}\sum_{t=1}^{T}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi)\right\|^{2}.$ (3)
Solving this problem provides a parameter estimate which is an effective predictor under the distribution of states and inputs seen during the experimentation. This notion can be captured via the prediction error.
Definition 2.1.
We define $\mathbf{Err}_{\pi}^{\phi^{\star}}(\phi)$ as the prediction error for a parameter $\phi$ under policy $\pi$:
$\mathbf{Err}_{\pi}^{\phi^{\star}}(\phi)=\mathbb{E}_{\pi}^{\phi^{\star}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|f(X_{t},U_{t};\phi)-f(X_{t},U_{t};\phi^{\star})\right\|^{2}\right].$
Once the learner estimates $\hat{\phi}$, the controller parameters are determined from the dynamics parameters by solving the policy optimization problem as
$\theta^{\star}(\hat{\phi})\in\operatorname*{argmin}_{\theta\in\operatorname{\mathbb{R}}^{d_{\theta}}}J(\pi^{\theta},\hat{\phi}).$ (4)
The certainty equivalent policy may then be expressed as a function of the estimated dynamics parameters $\hat{\phi}$ as
$\pi^{\star}(\hat{\phi})\triangleq\pi^{\theta^{\star}(\hat{\phi})}.$ (5)
We note that both the nonlinear least squares problem (3) and the certainty equivalent control synthesis procedure of (5) may be computationally challenging. The focus of this work is to understand the statistical complexity of the problem rather than the computational complexity. In the episodic setting we consider, both of these problems are solved offline. Therefore, given sufficient time and compute, it is often possible to determine good approximations to the optimal solutions using non-convex optimization solvers and approaches for policy optimization from the model-based reinforcement learning literature (Levine et al., 2020).
2.2 Assumptions
By (5), the optimal policy for the objective (2) under the true parameter $\phi^{\star}$ defining the dynamics (1) is thus given by $\pi^{\star}(\phi^{\star})$, and the corresponding objective value is $\mathcal{J}(\pi^{\star}(\phi^{\star}),\phi^{\star})$. Meanwhile, the objective value attained under an estimate $\hat{\phi}$ is $\mathcal{J}(\pi^{\star}(\hat{\phi}),\phi^{\star})$. We abuse notation and define the shorthand
$\mathcal{J}_{\tilde{\phi}}(\phi)\triangleq\mathcal{J}(\pi^{\star}(\phi),\tilde{\phi})$ (6)
to describe the control cost of applying a certainty equivalence policy synthesized using parameter $\phi$ on a system with dynamics described by $\tilde{\phi}$. It has been shown by Wagenmaker et al. (2021, 2023) that for models which are linear in the parameters, the gap $\mathcal{J}_{\phi^{\star}}(\phi)-\mathcal{J}_{\phi^{\star}}(\phi^{\star})$ is characterized by the squared parameter error weighted by the model-task Hessian, defined below.
Definition 2.2.
The model-task Hessian for objective (2) and dynamics (1) is given by
$\mathcal{H}(\tilde{\phi})=\nabla_{\phi}^{2}\mathcal{J}_{\tilde{\phi}}(\phi)|_{\phi=\tilde{\phi}},$
where $\mathcal{J}_{\tilde{\phi}}$ is defined in (6).
To express the excess cost achieved by a certainty equivalent controller synthesized using the estimated model parameters $\hat{\phi}$, we operate under the following smoothness assumption on the dynamics.
Assumption 2.1.
[…]
$\left\|D_{\phi}^{(i)}D_{u}^{(j)}f(x,u;\phi)\right\|_{\mathsf{op}}\leq L_{f}.$
The above assumption is satisfied for, e.g., control-affine dynamics which depend smoothly on $\phi$: $f(x_{t},u_{t};\phi)=g_{1}(x_{t};\phi)+g_{2}(u_{t};\phi)u$, with $g_{1}$ and $g_{2}$ each three time differentiable with respect to $\phi$. In this example, differentiability with respect to $u$ is immediate from the affine dependence.
We also require that the policy class $\Pi^{\star}$ is smooth.
Assumption 2.2.
(Smooth Policy Class) For $t=1,\dots,T$, $x\in\mathcal{X}$, and any policy $\pi\in\Pi^{\star}$, the function $\pi_{t}^{\theta}(x)$ is four-times differentiable in $\theta$. Furthermore, $\left\|D_{\theta}^{(i)}\pi_{t}^{\theta}(x)\right\|_{\mathsf{op}}\leq L_{\theta}$ for $i=1,\dots,4$, $\theta\in\operatorname{\mathbb{R}}^{d_{\theta}}$, $x\in\mathcal{X}$, and $t=1,\dots,T$.
Note that such smoothness conditions are not imposed for the exploration policy class $\Pi_{\exp}$. The exploration policy class could, for instance, consist of model predictive controllers with constraints on the injected input energy, which do not satisfy such smoothness assumptions.
We additionally require that the costs are bounded for policies in the class $\Pi_{\star}$ and all dynamics parameters in a neighborhood of the true parameter.
Assumption 2.3.
(Bounded Costs) There exists some $r_{\mathsf{cost}}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, and all $\pi\in\Pi^{\star}$, we have $\mathbb{E}_{\pi,\phi}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]\leq L_{\mathsf{cost}}$.
Due to the smoothness assumptions on the dynamics (Assumption 2.1) and the policy class (Assumption 2.2), the above assumption holds if the stage costs $c_{t}$ are smooth.
We additionally suppose that $\Pi^{\star}$ is not overparametrized.
Assumption 2.4.
The global minimum of $\mathcal{J}(\pi^{\theta},\phi^{\star})$, $\theta^{\star}(\phi^{\star})$, is unique and satisfies $\nabla^{2}_{\theta}\mathcal{J}(\pi^{\theta},\phi^{\star})|_{\theta=\theta^{\star}(\phi^{\star})}\succ 0$.
The above condition requires that there are a unique set of optimal control parameters for the true system. [^2] To gain an intuition for this condition, observe that it holds for linear quadratic regulators as long as the input penalty matrix is positive definite.
Under the above assumptions, we may show that the certainty equivalent policy is smooth near the optimal solution.
Lemma 2.1 (Proposition 6 of Wagenmaker et al. (2023)).
There exists some $r_{\theta}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\theta}(\phi^{\star}))$,
  • $\nabla_{\theta}\mathcal{J}(\pi^{\theta},\phi)|_{\theta=\theta^{\star}(\phi)}=0$
  • $\theta^{\star}(\phi)$ is three times differentiable in $\phi$ and $\left\|D_{\phi}^{i}\theta^{\star}(\phi)\right\|_{\mathsf{op}}\!\leq\!L_{\pi^{\star}}$ for some $L_{\pi^{\star}}>0$ and $i\in\mathopen{}\left\{1,2,3\right\}\mathclose{}$.
In order to bound the parameter recovery error in terms of the prediction error, additional identifiability conditions are needed. The following definition of a Lojasiewicz exploration policy is determined from a Lojasiewicz condition that arises in the optimization literature that measures the sharpness of an objective near its optimizer (Roulet and d’Aspremont, 2017). In our setting, it quantifies the degree of identifiability from using a particular exploration policy. It does so by bounding the growth of identification error as a polynomial of prediction error.
Definition 2.3 (Lojasiewicz condition, Roulet and d’Aspremont (2017)).
For positive numbers $C_{\mathsf{Loja}}$ and $\alpha$, we say that a policy $\pi\in\Pi_{\exp}$ is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz if
$\left\|\hat{\phi}-\phi^{\star}\right\|\leq C_{\mathsf{Loja}}\mathbf{Err}_{\pi}^{\phi^{\star}}(\hat{\phi})^{\alpha}.$
To ensure parameter recovery is possible for the learner, we make the following assumption regarding identifiability.
Assumption 2.5.
Fix some positive constant $C_{\mathsf{Loja}}$ and $\alpha\in(\frac{1}{4},\frac{1}{2}]$. The learner has access to a policy $\pi^{0}\in\Pi_{\exp}$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz.
While the Lojasiewicz assumption ensures that the data collected via the exploration policy $\pi^{0}$ is sufficient to identify the parameters, the rate of recovery may be slow. To bypass this limitation, we assume that some policy in the exploration class satisfies a persistence of excitation condition. Persistence of excitation for a nonlinear dynamical system involves the positive definiteness of the following matrix:
$\Sigma^{\pi}\triangleq\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right],$ (7)
where $Df(X_{t},U_{t},\phi^{\star})$ is the Jacobian of $f$ with respect to $\phi$ evaluated at $\phi^{\star}$.
Assumption 2.6.
There exists a policy $\pi\in\Pi_{\exp}$ for which
$\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}>0,$
Finally, we make an assumption that all models parameters are relevant to perfect control.
Assumption 2.7.
The model-task Hessian is positive definite: $\mathcal{H}(\phi^{\star})\succ 0$.
The above assumption is quite strong, and does not hold for common baseline problems such as the linear-quadratic regulator with fully unknown dynamics matrices. However, this assumption is made primarily for ease of exposition. It can be removed in the linear setting at the expense of an algorithm which is more complicated to express concisely (Wagenmaker et al., 2023). The same approach used to bypass this condition in the linear setting can be easily extended to the nonlinear setting considered in this paper. We do not formalize this here, although we will in future work.
3 Proposed Algorithm and Main Result
The above smoothness assumptions allow us to characterize the excess control cost of a policy synthesized via certainty equivalence applied to a parameter estimate $\hat{\phi}$, (5).
Lemma 3.1 (Thm. 1 of Wagenmaker et al. (2023)).
 [^3] Thm. 1 of Wagenmaker et al. (2023) holds for systems with linear dependence on the unknown parameters; however it generalizes with minimal modification to the proof. See Appendix B.
Suppose Assumptions 2.1-2.4 hold. Let $r_{\theta}(\phi^{\star})$ be as defined in Lemma 2.1. Then for $\hat{\phi}\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
$\mathcal{J}_{\phi^{\star}}(\hat{\phi})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\!\leq\!\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})}+C_{\mathsf{cost}}\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{3},$ (8)
where $\mathcal{J}_{\phi^{\star}}(\phi)$ is as defined in (6) and
$C_{\mathsf{cost}}=\mathsf{poly}(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}).$
Footnote 3 informs us that the leading term of the excess cost is given by the parameter estimation error weighted by the model-task Hessian, $\left\|\hat{\phi}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}$.
Asymptotically, the distribution of the parameter estimation error is normally distributed with mean zero, and covariance given by the inverse Fisher Information matrix under the data collection policy $\pi$ evaluated at the true parameter value (cf. Theorem 1 of Ljung and Caines (1980)):
$\lim_{N\to\infty}\sqrt{N}\mathbf{FI}^{\pi}(\phi^{\star})^{1/2}(\hat{\phi}-\phi_{\star})\sim\mathcal{N}(0,I),$
where the Fisher information for the parameter $\phi$ under the policy $\pi$ for the model (1) is given as
$\mathbf{FI}^{\pi}(\phi)\!=\!\frac{\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}\!Df(X_{t},U_{t},\phi)^{\top}\!Df(X_{t},U_{t},\phi)\right]}{\sigma_{w}^{2}},$ (9)
and measures the signal-to-noise ratio of the data collected from an episode of interaction with the system under the exploration policy $\pi$.
We provide a novel non-asymptotic result which characterizes the $H$-norm of the parameter error for a positive definite matrix $H$ in terms of the Fisher Information matrix.
Theorem 3.1.
Suppose Assumption 2.1 holds. Let $H$ be a positive definite matrix and let $\delta\in(0,\frac{1}{4}]$. Consider the least squares estimate $\hat{\phi}$ determined from (3) using data collected from $N$ episodes via an exploration policy $\pi$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz for some $\alpha\in(\frac{1}{4},\frac{1}{2}]$, and satisfies $\lambda_{\min}\left(\Sigma^{\pi}\right)>0$, with $\Sigma^{\pi}$ as defined in (7). There exists a polynomial $\texttt{poly}_{\alpha}$ depending only on $\alpha$ such that the following condition holds. With probability at least $1-\delta$,
$\left\|\hat{\phi}-\phi^{\star}\right\|_{H}^{2}$ $\leq 40\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H\mathbf{FI}^{\pi}(\phi^{\star})^{-1})}{N}$ (10)
as long as
$N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},d_{\phi},d_{\mathsf{X}},\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
By substituting the inequality (10) into the leading term of (8), one would expect that the the excess cost of deploying the certainty equivalent policy synthesized on a least squares estimate determined from data collected using policy $\pi$ is characterized by
$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}).$ (11)
In light of this, we would like to choose the exploration policy $\pi$ which minimizes this upper bound:
$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}).$ (12)
Our main result shows that the above intuition can be made rigorous through careful analysis and design of the exploration policy. It then proceeds to show that we can find an exploration policy approximately solving (12), even though the parameter $\phi^{\star}$ defining the exploration objective is unknown prior to experimentation.
To circumvent the issue of the unknown parameter $\phi^{\star}$, we consider a two step approach in which we first obtain a crude parameter estimate, and then refine it by playing a targeted exploration policy. Denote the crude estimate by $\hat{\phi}^{-}$. This parameter can be used to search for a policy that approximately solves the optimization problem in (12). The straightforward approach to do so is to solve the problem under the estimated parameter:
$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}(H(\hat{\phi}^{-})\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}).$ (13)
While such a strategy may be effective if the coarse estimate is sufficiently close to the ground truth, it contends with the challenge that there is a distribution shift between the setting that the exploration policy is optimized for, and that which it is deployed in. In particular, $\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-1})$ pertains to the statistics from rolling out the dynamics under a parameter value different from the ground truth. We therefore instead consider the following problem:
$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\right)\mbox{ where }$ (14)
$\Lambda_{\tilde{\pi}}\!=\!\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\!\left[\sum_{t=1}^{T}Df(X_{t},U_{t};\hat{\phi}^{-})^{\top}Df(X_{t},U_{t};\hat{\phi}^{-})\right]$ (15)
Note the difference between (13) and (14). In (13), the Fisher information is determined using an expectation under the the dynamics parameter estimate $\hat{\phi}^{-}$ (see (9)). In contrast, the expectation in (14) is evaluated under the true parameter $\phi^{\star}$. If (14) can be optimized over the exploration policies, it alleviates the aforementioned distribution shift issue.
A minimizer for the above objective (14) may be approximated via an online learning procedure, as was recently proposed by Wagenmaker et al. (2023) for models that are linear in the parameter. In particular, we leverage the dynamic optimal experiment design (DOED) procedure proposed in the aforementioned work. For the reader’s convenience, we state the algorithm and highlight several modifications in Section C.1. We refer to our modified version of the algorithm as $\texttt{DOED}_{+}$.
The above discussion motivates Algorithm 1, named Active Learning for Control-Oriented Identification (ALCOI). The algorithm takes as input an initial policy satisfying the Lojasiewicz condition (Assumption 2.5), the exploration policy class, the target policy, the number of exploration rounds, and the desired confidence parameter $\delta$. Given these components, the algorithm proceeds in three stages. The first stage begins in 2 by playing the initial policy for one fourth of the exploration budget. In 3, it uses the collected data to derive a coarse estimate $\hat{\phi}^{-}$ for the unknown parameters by solving a least squares problem. Then, in 4 it constructs a confidence ball $\mathcal{B}^{-}$ centered at $\hat{\phi}^{-}$:
$\mathcal{B}^{-}$ =$=\bigg{\{}\phi:\left\|\phi-\hat{\phi}^{-}\right\|\leq\frac{R}{N^{\alpha}}\bigg{\}},$ (16)
where
$R=\left(\frac{2048\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(L_{f}TN\right)+\log\frac{1}{\delta}\right)\right)^{\alpha}.$ (17)
Next, the estimate $\hat{\phi}^{-}$ is used to construct the model-task Hessian at $\hat{\phi}^{-}$ as $\mathcal{H}(\hat{\phi}^{-})$. This is in turn used to define an exploration objective, $\Phi:\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}$ to be passed to the $\texttt{DOED}_{+}$ algorithm. This exploration objective is a regularized version of that in (14) which takes as input a Gram matrix $\Lambda$, and outputs $\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$. Here $\lambda>0$ is a regularizer chosen from the estimate of the model-task Hessian in 5. The quantity $B_{\Phi}$ is then defined from the model-task Hessian and the regularizer to characterize the smoothness of the objective $\Phi$. The exploration objective is passed to $\texttt{DOED}_{+}$ along with the coarse estimate $\hat{\phi}^{-}$, the ball $\mathcal{B}^{-}$ and the smoothness bound $B_{\Phi}$ in 8. This call to $\texttt{DOED}_{+}$ returns a collection of $\tilde{K}$ policies, where $\tilde{K}$ is a polynomial function of the supplied budget; see Section C.1 for further details. The mixture of the returned policies approximately minimizes the exploration objective $\Phi$ over Gram matrices of the form (15), thereby approximating a solution to (14). We therefore use them to design a mixture policy which plays any of the returned policies with equal probability, and plays the initial exploration policy $\pi^{0}$ with probability $\frac{1}{2}$. [^4] The initial policy $\pi^{0}$ included in $\pi_{\mathsf{mix}}$ to ensure that $\pi_{\mathsf{mix}}$ is a $(2C_{\mathsf{Loja}},\alpha)$-Lojasiewicz policy. This policy is run to collect data from the system, and obtain a fresh estimate $\hat{\phi}^{+}$ for $\phi^{\star}$. [^5] In practice, one would use all of the the data to estimate $\hat{\phi}^{+}$. However, using only the data from an independent rollout contributes only a constant factor to the rate, and simplifies the analysis. Finally, the estimate is used to synthesize the certainty equivalent policy as in (5).
Algorithm 1 $\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$
1:Input: Initial policy $\pi^{0}$, exploration policy class $\Pi_{\exp}$, target policy class $\Pi^{\star}$, episodes $N$, confidence $\delta$.
2: Play $\pi^{0}$ for $\lfloor\frac{N}{4}\rfloor$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},x_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{4}\rfloor}$.
3: Fit $\hat{\phi}^{-}$ from the collected data by solving (3).
4: Set $\mathcal{B}^{-}$ according to (16).
5: Set $\lambda\leftarrow\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}$.
6: Define objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$.
7: Set objective smoothness bound $B_{\Phi}\leftarrow\frac{TL_{f}^{2}\left\|\mathcal{H}(\hat{\phi}^{-})\right\|}{\lambda^{2}}$.
8: Determine exploration policies as $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}\leftarrow\texttt{DOED}_{+}(\Phi,\lfloor\frac{N}{4}\rfloor,\pi^{0},\hat{\phi}^{-},\mathcal{B}^{-},B_{\Phi})$.
9: Define $\pi_{\mathsf{mix}}$ which at the start of each episode plays $\pi^{0}$ with probability $\frac{1}{2}$, and for $i=1,\dots,\tilde{K}$ plays $\pi^{i}$ with probability $\frac{1}{2\tilde{K}}$.
10:Play $\pi_{\mathsf{mix}}$ for $\lfloor N/2\rfloor$ episodes, collecting data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
11: Fit $\hat{\phi}^{+}$ by solving (3) with the data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
12:Return: certainty equivalent policy $\hat{\pi}=\pi^{\star}(\hat{\phi}^{+})$
Our main result is a finite sample bound characterizing the excess cost of the policy return by Algorithm 1.
Theorem 3.2 (Main Result).
Suppose $f$, $\pi^{0}$, $\Pi_{\exp}$, $\Pi^{\star}$ satisfy Assumptions 2.1-2.7. Let $\delta\in(0,\frac{1}{4}]$. Consider running Algorithm 1 to generate a control policy $\hat{\pi}$ as $\hat{\pi}=\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$. There exists a universal positive constant $c$ and polynomial function $\texttt{poly}_{\alpha}$ depending only on the Lojasiewicz parameter $\alpha$ such that the following holds true. With probability at least $1-\delta$ we have that
$\mathcal{J}(\hat{\pi},\phi^{\star})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\leq c\frac{\log\frac{d_{\phi}}{\delta}}{N}\!\min_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\right),$
as long as
$N\!$ $\geq\!\mathsf{poly}_{\alpha}\bigg{(}\!T,L_{f},L_{\mathsf{cost}},L_{\pi^{\star}},d_{\phi},d_{\mathsf{X}},\frac{1}{\lambda_{\min}^{\star}},r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})$
$\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))},\left\|\mathcal{H}(\phi^{\star})\right\|,\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
The above result characterizes the excess control cost in terms of three key quantities. First, the term $\mathcal{H}(\phi^{\star})$ is the model-task Hessian, which describes how error in identification of the dynamics model parameters translates to the control cost. Second, is the inverse Fisher Information of the optimal exploration policy term, which measures a signal-to-noise ratio quantifying the hardness of parameter identification. Finally, the number of exploration episodes $N$ on the denominator captures the rate of decay from increasing the experimental budget.
We note that in the setting where the dynamics model has linear dependence on the parameters, Wagenmaker et al. (2023) present a lower bound on the excess control cost achieved by any learner following the model based interaction protocol described in Section 2. The upper bound of Theorem 3.2 matches this lower bound up to universal constants, and the log term $\log\left(d_{\phi}/\delta\right)$. Future work will pursue general lower bounds that hold for dynamics models with a nonlinear dependence on the unknown parameter. We conjecture that in general, the upper bounds of Theorem 3.2 are tight up to universal constants and the term $\log\left(d_{\phi}/\delta\right)$.
We note that the burn-in time is currently polynomial in the relevant system parameters; however, we do not pursue tight burn-in times in this work. It may be possible to improve the dependence of the burn-in on various system quantities, e.g. by leveraging stability or reachability to obtain optimal dependence of the burn-in on $T$.
4 Proof Sketch
Full proof details may be found in the appendix. Here, we present a sketch. Our main result proceeds by demonstrating the following sub-steps. In these sub-steps, let $C$ be a polynomial of the problem parameters, as in the burn-in requirement of Theorem 3.2.
  1. 1.
    The coarse parameter estimation error decays gracefully with the total amount of data:
    $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{C}{(TN)^{\alpha}},$ (18)
    as long as $N$ exceeds some polynomial burn-in time. This result is derived from recent results characterizing non-asymptotic bounds for identification (Ziemann and Tu, 2022), and takes the place of the estimator consistency requirements in classical asymptotic identification literature (Ljung and Caines, 1980). By making the number of episodes $N$ sufficiently large, we can make this error arbitrarily small. It thus characterizes a type of “consistency burn-in”.
  2. 2.
    As long as the coarse estimation error of (18) is sufficiently small, the ideal optimal exploration objective of (12) is well-approximated by the objective (14). In particular, for any exploration policy $\pi\in\Pi_{\exp}$,
    $\bigg{|}\!\operatorname{\mathrm{tr}}\left(\!\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi}^{-1}\!\right)\!-\!\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})\bigg{|}\!\leq\!C\!\left\|\hat{\phi}^{-}\!\!-\!\phi^{\star}\right\|,$ (19)
    with $\Lambda_{\pi}$ as in (15).
  3. 3.
    The call to the $\texttt{DOED}_{+}$ algorithm in 8 of Algorithm 1 returns a collection of policies whose mixture approximately minimizes the exploration objective (14):
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi_{\mathsf{mix}}}^{-1}\right)\leq\frac{C}{N^{\alpha}}\!+\!\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}\mathcal{H}(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\Bigg{)},$ (20)
    for $\Lambda_{\pi}$ as in (15). Demonstrating this result involves specializing the dynamic optimal exploration algorithm and its proof of correctness to our setting from Wagenmaker et al. (2023).
  4. 4.
    For $N$ sufficiently large, we may use the consistency guarantee (18) to prove Theorem 3.1. The proof of this fact follows by revisiting the delta method (Van der Vaart, 2000) through the lens of concentration inequalities. Doing so results in the near sharp [^6] Rates matching the asymptotic limit up to logarithmic factors. rates we obtain.
Using the above results, our argument proceeds according to the following series of inequalities applied to the excess cost:
$\mathcal{J}_{\phi^{\star}}(\hat{\phi}^{+})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\leq\left\|\hat{\phi}^{+}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}\!+\!C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{3}$ (21)
$\leq 40\log\left(\frac{24d_{\phi}}{\delta}\right)\frac{\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})}{N}+\frac{C}{N^{3/2}},$
where the first inequality follows by Footnote 3, and the second inequality follows by applying Theorem 3.1 with $H=\mathcal{H}(\phi^{\star})$ for the first term, and $H=I$ for the second term [^7] The condition that $\Sigma^{\pi_{\mathsf{mix}}}\succ 0$ in Theorem 3.1 can be shown from the success of DOED in (20)). . Next, it follows that
$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})$ $\overset{(i)}{\leq}\frac{C}{N^{\alpha}}+\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\right)$
$\overset{(ii)}{\leq}2\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\Bigg{)}$
$\overset{(iii)}{\leq}3\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\Bigg{)},$
where inequality $(i)$ follows from (19) and (18), inequality $(ii)$ follows from (20), and inequality $(iii)$ follows from (19) and (18). The main result then follows by substituting the above bound into (21), and taking $N$ to exceed a polynomial burn-in time so the higher order terms become negligible.
5 Numerical Validation
We deploy ALCOI on an illustrative example to illustrate the benefits of active control-oriented exploration. For more experiments, and further details, see Appendix E. Consider the two dimensional system
$X_{t+1}=X_{t}+U_{t}+W_{t}+\sum_{i=1}^{4}\psi(X_{t}-\phi_{\star}^{(i)})$
with $X_{t},U_{t},W_{t}$ and $\phi_{\star}^{(i)}$ assuming values in $\mathbb{R}^{2}$. Here $\psi:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is defined by $\psi(x)=5\frac{x}{\left\|x\right\|}\exp(-x^{2})$. The noise is distributed according to a standard normal distribution. The parameters $\phi_{\star}^{(1)}$, $\phi_{\star}^{(2)}$, $\phi_{\star}^{(3)}$, $\phi_{\star}^{(4)}$ are set as $\begin{bmatrix}5\\ 0\end{bmatrix},\begin{bmatrix}-5\\ 0\end{bmatrix},\begin{bmatrix}0\\ 5\end{bmatrix}$ and $\begin{bmatrix}0\\ -5\end{bmatrix}$, respectively.
We consider model-based reinforcement learning with a horizon $T=10$ and quadratic cost functions: for all $t\in[T]$,
$c_{t}(x,u)=\left\|x-\begin{bmatrix}5.5\\ 0\end{bmatrix}\right\|^{2},\,c_{T+1}(x)=\left\|x-\begin{bmatrix}5.5\\ 0\end{bmatrix}\right\|^{2}.$
The policy class $\Pi^{\star}$ consists of feedback linearization controllers defined by parameters $\theta=(K,\hat{\phi}^{(1)},\dots,\hat{\phi}^{(4)})$, with $K\in\mathbb{R}^{2\times 2}$ and $\hat{\phi}^{(i)}\in\mathbb{R}^{2}$ for $i=1,\dots,4$:
$\pi^{\theta}(X_{t})=K\left(X_{t}-\begin{bmatrix}-5.5\\ 0\end{bmatrix}\right)-\sum_{i=1}^{4}\psi(X_{t}-\hat{\phi}^{(i)}).$
The exploration class $\Pi_{\exp}$ consists of policies with input energy bounded by $T$: $\sum_{t=1}^{T}\left\|U_{t}\right\|^{2}\leq T$.
We compare ALCOI with random exploration and approximate $A$-optimal experiment design. For random exploration, the learner injects isotropic Gaussian noise which is normalized such that $\sum_{t=1}^{T}\left\|U_{t}\right\|^{2}=T$. For approximate $A$-optimal experiment design, the learner runs the ALCOI, but with the model-task Hessian estimate, $\mathcal{H}(\hat{\phi}^{-})$, replaced by $I$.
Figure 2: Comparison of the proposed control-oriented identification procedure with approximate $A$-optimal design, and random experiment design. The mean over $100$ runs is shown, with the standard error shaded.
Figure 2 illustrates that ALCOI achieves a lower excess control cost than the alternatives at all iterations. To understand why this is the case, note that in order to regulate the system to the position $X_{t}=\begin{bmatrix}5.5&0\end{bmatrix}^{\top}$, the parameter $\phi_{\star}^{(1)}$ must be identified accurately. However, due to the Gaussian kernel, accurately estimating $\phi_{\star}^{(1)}$ requires that the experiment data consists of trajectories where the state is near $\phi_{\star}^{(1)}$. Random exploration clearly fails to collect such trajectories. Approximate $A$-optimal experiment design does collect such trajectories; however, it also collects trajectories steering the state to $\begin{bmatrix}-5&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&5\end{bmatrix}^{\top}$, and $\begin{bmatrix}0&-5\end{bmatrix}^{\top}$ in order to identify the parameters $\phi_{\star}^{(2)},\phi_{\star}^{(3)}$ and $\phi_{\star}^{(3)}$. ALCOI, in contrast, designs experiments that are effective for identifying the parameters most relevant for control. For the chosen objective, this means that the algorithm invests the most exploration energy in collecting data in the neighborhood of $\phi_{\star}^{(1)}$. This illustrative example hints at the practical benefit of the proposed approach.
6 Conclusions
We have introduced and analyzed the Active Learning for Control-Oriented Identification (ALCOI) algorithm, marking a significant step towards understanding active exploration in model-based reinforcement learning for a general class of nonlinear dynamical systems. We provide finite sample bounds on the excess control cost achieved by the algorithm which offer insight into the interaction between the hardness of control and identification. Our bounds are known to be sharp up to logarithmic factors in the setting of nonlinear dynamical systems with linear dependence on the parameters, and we conjecture that they are sharp in general. Future work will attempt to verify that this is the case. It would also be interesting for future work to consider learning partially observed dynamics using general prediction error methods, rather than assuming a noiseless state observation.
ACKNOWLEDGMENT
Ingvar Ziemann is supported by a Swedish Research Council international postdoc grant. George J. Pappas is supported in part by NSF Award SLES-2331880. Nikolai Matni is supported in part by NSF Award SLES-2331880 and NSF CAREER award ECCS-2045834.
Appendix A System Identification Results
In this section we consider $\mathopen{}\left\{X_{t}^{k},U_{t}^{k},X_{t+1}^{k}\right\}\mathclose{}_{t,k=1}^{T,K}$ with joint probability distribution $\mathsf{P}^{K}=\mathsf{P}\times\dots\times\mathsf{P}$ induced playing an exploration policy $\pi\in\Pi_{\mathsf{exp}}$ on the system dynamics (1). As we consider a single dynamics parameter $\phi^{\star}$, and a single data collection policy $\pi$, we drop the subscripts $\pi$ and $\phi^{\star}$ from the expectations and probabilities.
The identification results are organized as follows. First, in Section A.1, we show that under a policy satisfying a Lojasiewicz condition Definition 2.3, the least squares parameter estimation error decays with the number of episodes $N$. Section A.2 leverages the smoothness of the dynamics and the consistency result from Section A.1 to express the parameter error in terms of a self-normalized martingale. This self-normalized martingale is then bounded in Section A.3 to prove Theorem 3.1.
A.1 Consistency of Least Squares Parameter Estimation
Recent developments Ziemann and Tu (2022); Ziemann et al. (2023) have provided finite sample bounds for the $L_{2}$ prediction error of nonlinear least squares estimators in the presence of time dependent data with rates that do not depend on the mixing time of the data. Our first estimation bound instantiates these results in our setting under the assumption that the data collection policy is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz. To express these bounds we define $\mathcal{G}=\mathopen{}\left\{f(\cdot,\cdot,\phi):\phi\in\mathbb{R}^{d_{\phi}}\right\}\mathclose{}$ and $\mathcal{G}_{\star}=\mathcal{G}-\mathopen{}\left\{f(\cdot,\cdot,\phi^{\star})\right\}\mathclose{}$. Using the Lipschitz bound on $f$, we find that the size $\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)$ of the the minimal cardinality $\varepsilon$-net covering $\mathcal{G}_{\star}$ is bounded as
$\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)\leq d_{\phi}\log\left(\frac{2L_{f}}{\varepsilon}+1\right).$ (22)
We will present an estimation error bound that holds as long as the number of episodes exceeds some value (a burn-in time) depending on the confidence with which we want it to hold. To state this requirement, define
$\tau_{\mathbf{Err}}(\delta)=\max\bigg{\{}\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{4\alpha-1}},$ (23)
$\Bigg{(}4(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}\left(\log\left(4BK^{2\alpha}\right)+d_{\phi}\log\left(2L_{f}K^{2\alpha}+1\right)\right)+\log\frac{1}{\delta}\Bigg{)}^{\frac{1}{1-8\alpha+16\alpha^{2}}}\bigg{\}}.$
Lemma A.1.
Suppose Assumption 2.1 holds, and $\delta\in(0,\frac{1}{2}]$. Let $\hat{\phi}$ be the least squares estimate from (3) using $K$ episodes of data collected with a $(C_{\mathsf{Loja}},\alpha)$ policy, for $K\geq\tau_{\mathsf{ER}}(\delta)$. Then with probability at least $1-\delta$
$\mathbf{Err}_{\pi,\phi^{\star}}(\hat{\phi})\leq\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(\frac{3L_{f}TK}{\delta}\right)\right).$
To prove this result, we first state several preliminary results. In order to do so, we introduce additional notation. Let $\mathcal{G}_{\star}^{s}=\mathopen{}\left\{bg:b\in[0,1],g\in\mathcal{G}_{\star}\right\}\mathclose{}$ be the star-shaped hull of $\mathcal{G}_{\star}$. [^8] A function class $\mathcal{G}$ is star-shaped if for all $b\in[0,1]$, $g\in\mathcal{G}\implies bg\in\mathcal{G}$. Since all elements of $\mathcal{G}_{\star}^{s}$ satisfy $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$, we have from Lemma 4.5 of Mendelson (2002) that the size $\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)$ of a minimum cardinality $\varepsilon$-net of $\mathcal{G}_{\star}^{s}$ in the $\left\|\cdot\right\|_{\infty}$-norm satisfies
$\log\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)\leq\log\left(4B/\varepsilon\right)+\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon).$ (24)
Define $B(r)=\mathopen{}\left\{g\in\mathcal{G}_{\star}^{s}:\sqrt{\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]}\leq r\right\}\mathclose{}$, and denote the boundary of this set as $\partial B(r)$.
We begin by stating the definition of hypercontractivity from Ziemann and Tu (2022).
Definition A.1 (Trajectory $(C_{\mathsf{hc}},a)$-hypercontractivity).
Fix constants $C_{\mathsf{hc}}>0$ and $a\in[1,2]$. The tuple $(\mathcal{G},\mathsf{P})$ consisting of the function class $\mathcal{G}$ and the distribution $\mathsf{P}$ satisfies $(C_{\mathsf{hc}},a)$-hypercontractivity if
$\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{4}\right]\leq C_{\mathsf{hc}}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{a}\quad\forall g\in\mathcal{G},$
where the expectation is with respect to $\mathsf{P}$, the joint probability distribution of $\mathopen{}\left\{X_{t}^{1},U_{t}^{1},X_{t+1}^{1}\right\}\mathclose{}_{t=1}^{T}$.
We now state an exponential inequality which we will leverage along with hypercontractivity to bound the population risk by the empirical risk for any $g$ in the boundary of $\mathcal{G}_{\star}^{s}$.
Lemma A.2.
Let $g$ belong to $\mathcal{G}_{\star}^{s}$. For every $\lambda\in\mathbb{R}_{+}$ we have
$\operatorname{\mathbb{E}}\exp\left(-\lambda\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\right)\leq\exp\left(-\lambda K\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]+\frac{\lambda^{2}K}{2}\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\right]\right).$
Proof.
Observe that for $x\geq 0$, the inequality $e^{-x}\leq 1-x+x^{2}/2\leq e^{-x+x^{2}/2}$ holds. Using this and the fact that the trajectories are independent and identically distributed, we find
$\operatorname{\mathbb{E}}\exp\left(-\lambda\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\right)$ $=\prod_{j\in[K]}\operatorname{\mathbb{E}}\exp\left(-\lambda\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)$
$\leq\prod_{j\in[K]}\operatorname{\mathbb{E}}\left(1-\lambda\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}+\frac{\lambda^{2}}{2}\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|\right)^{2}\right)$
$\leq\exp\left(-\lambda K\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]+\frac{\lambda^{2}K}{2}\operatorname{\mathbb{E}}\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\right).$
Using the above exponential inequality along with a Chernoff bound provides the following high probability bound on the empirical prediction error by the population counterpart.
Lemma A.3.
Suppose $(\partial B(r),\mathsf{P})$ is $(C_{\mathsf{hc}},a)$-hypercontractive. For every $g\in\partial B(r)$, we have
$\operatorname{\mathbb{P}}\left[\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}<\frac{1}{2}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\right]$ $\leq\exp\left(-\frac{K}{4C_{\mathsf{hc}}}r^{4-2a}\right).$
Proof.
A Chernoff bound along with Lemma A.2 yields the estimate
$\operatorname{\mathbb{P}}\left[\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}<\frac{1}{2}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\right]$
$\leq\inf_{\lambda\geq 0}\exp\left(-\lambda K\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]+\frac{\lambda^{2}K}{2}\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\right]\right)$
$=\exp\left(-\frac{K\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{2}}{4\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\right]}\right).$
By Cauchy-Schwarz,
$\operatorname{\mathbb{E}}\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\leq T\operatorname{\mathbb{E}}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{4}=T^{2}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{4}\right].$
Therefore, by the fact that $(\partial B(r),\pi)$ is $(C_{\mathsf{hc}},a)$-hypercontractive, we have that
$\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{4}\right]\leq C_{\mathsf{hc}}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{a}.$
As a result, we may bound
$\exp\left(-\frac{K\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{2}}{4\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right)^{2}\right]}\right)\leq\exp\left(-\frac{K\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{2-a}}{4C_{\mathsf{hc}}}\right).$
To conclude, we note that $\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t},U_{t})\right\|^{2}\right]=r^{2}$ by the fact that $g\in\partial B(r)$. ∎
The above bound for a single member of the function class may be extended to a uniform bound over function in $\mathcal{G}_{\star}^{s}$, but outside the ball $B(r)$ using a covering argument.
Lemma A.4 (Lower isometry).
Suppose there exists a $r/\sqrt{8}$-net $\mathcal{G}_{r}$ of $\partial B(r)$ in the norm $\left\|\cdot\right\|_{\infty}$ such that $(\mathcal{G}_{r},\pi)$ satisfies $(C_{\mathsf{hc}},a)$-hypercontractivity. Then the following lower isometry holds
$\operatorname{\mathbb{P}}\left[\sup_{g\in\mathcal{G}_{\star}^{s}\setminus B(r)}\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
Proof.
The fact that $\mathcal{G}_{\star}^{s}$ is shaped allows us to the restrict attention to $g\in\partial B(r)$. By the fact that $\mathcal{G}_{r}$ is an $r/\sqrt{8}$-net of $\partial B(r)$, we have by the reverse triangle inequality that for any $g\in\partial B(r)$, there exists some $g_{i}\in\mathcal{G}_{r}$ such that
$\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\geq\frac{1}{2TK}\sum_{t,k=1}^{T,K}\left\|g_{i}(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{r^{2}}{8}.$ (25)
Define the event
$\mathcal{E}=\bigcup_{g\in\mathcal{G}_{r}}\mathopen{}\left\{\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\leq\frac{1}{2}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\right\}\mathclose{}.$
It follows from a union bound along with Lemma A.3 that
$\operatorname{\mathbb{P}}\left[\mathcal{E}\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
Fix an arbitrary $g\in\partial B(r)$. On the event $\mathcal{E}^{c}$, we have
$\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}$ $\geq\frac{1}{2TK}\sum_{t,k=1}^{T,K}\left\|g_{i}(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{r^{2}}{8}$ (We may find such a $g_{i}$ by (25))
$\geq\frac{1}{4}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g_{i}(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]-\frac{r^{2}}{8}$ (By definition of $\mathcal{E}$)
$=\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g_{i}(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]$ $(g_{i}\in\partial B(r)).$
Since $g\in\partial B(r)$ was arbitrary, we have that
$\operatorname{\mathbb{P}}\left[\sup_{g\in\partial B(r)}\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
To extend the inequality in the expectation to hold for the set $\mathcal{G}_{\star}^{s}\setminus B(r)$, we observe that by the fact that we may rescale $g$ to lie in $\partial B(r)$ since $\mathcal{G}_{\star}^{s}$ is star-shaped. ∎
We now proceed with our proof of Lemma A.1.
Proof.
We first verify that $\mathcal{G}_{\star}^{s}$ is hypercontractive. By proposition 4.2 of Ziemann et al. (2024), the fact that $\pi$ is a $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz policy and $\left\|Df(x,u;\phi)\right\|\leq L_{f}$ (Assumption 2.1) implies that for all $g\in\mathcal{G}_{\star}$,
$\lim_{p\to\infty}\left(\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{p}\right]\right)^{1/p}\leq L_{f}C_{\mathsf{Loja}}\alpha^{-1}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]^{\alpha/2}.$
We additionally have that
$\left(\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{4}\right]\right)^{1/4}\leq\lim_{p\to\infty}\left(\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{p}\right]\right)^{1/p}.$
Combining these inequalities implies that $(\mathcal{G}_{\star},\mathsf{P})$ is $(C_{\mathsf{hc}},a)$ hypercontractive, where $C_{\mathsf{hc}}=(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}$ and $a=4\alpha$. By the following elementary inequality, the star-shaped hull $\mathcal{G}_{\star}^{s}$ satisfies hypercontractivity with the same parameters: for $b\in[0,1]$ and $a\leq 2$, $b^{4}<b^{2a}$.
Bounding the population risk by the empirical risk:
To bound the worst-case difference between the excess risk and its empirical counterpart over all functions in the class $\mathcal{G}_{\star}$, we split the problem into two cases: one in which $g$ belongs to the ball $B(r)$, and one in which it does not.:
$\sup_{g\in\mathcal{G}_{\star}}\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]-\frac{8}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}$
$\leq\sup_{g\in\mathcal{G}_{\star}\cap B(r)}\left(\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]-\frac{8}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\right)$
$+\sup_{g\in\mathcal{G}_{\star}\setminus B(r)}\left(\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]-\frac{8}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\right).$
The first term is bounded by $r^{2}$ by the definition of $B(r)$ and the fact that $-\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\leq 0.$ By Lemma A.4, the second term is less than zero with probability at least $1-{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right)$. Then under this event, the following bound holds for all $g\in\mathcal{G}_{\star}$.
$\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 8\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}+r^{2}.$
We may bound $\log{\left|\mathcal{G}_{r}\right|}$ using (24) and (22) as $\log{\left|\mathcal{G}_{r}\right|}\leq\log\left(\frac{4B}{r}\right)+d_{\mathsf{X}}\log\left(\frac{2L_{f}}{r}+1\right).$ Then choosing $r^{2}=\frac{1}{K^{a}}$, we find that by the requirement that
$K\geq\left(4C_{\mathsf{hc}}\left(\log\left(4BK^{a/2}\right)+d_{\phi}\log\left(2L_{f}K^{a/2}+1\right)\right)+\log\frac{1}{\delta}\right)^{\frac{1}{1-2a+a^{2}}},$
the inequality
$\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 8\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}+\frac{1}{K^{a}}$ (26)
is satisfied for all $g\in\mathcal{G}_{\star}$ with probability at least $1-\delta$.
Bounding the empirical risk by the offset complexity: We now seek to bound the empirical risk. By the fact that $\hat{\phi}$ solves (3), it follows that
$\sum_{t,k=1}^{T,K}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\hat{\phi})\right\|^{2}\leq\sum_{t,k=1}^{T,K}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi^{\star})\right\|^{2}.$
Substituting in $X_{t+1}^{n}=f(X_{t}^{n},U_{t}^{n};\phi^{\star})+W_{t}^{k}$ and rearranging, we find
$\sum_{t,k=1}^{T,K}\left\|f(X_{t}^{n},U_{t}^{n},\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star})\right\|^{2}\leq 2\sum_{t,k=1}^{T,K}\langle f(X_{t}^{n},U_{t}^{n};\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star}),W_{t}^{k}\rangle.$
Multiplying the above inequality by two, and subtracting $\sum_{t,k=1}^{T,K}\left\|f(X_{t}^{n},U_{t}^{n},\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star})\right\|^{2}$ from each side leads to the following offset basic inequality:
$\sum_{t,k=1}^{T,K}\left\|f(X_{t}^{n},U_{t}^{n},\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star})\right\|^{2}$
$\leq 4\sum_{t,k=1}^{T,K}\langle f(X_{t}^{n},U_{t}^{n};\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star}),W_{t}^{k}\rangle-\sum_{t,k=1}^{T,K}\left\|f(X_{t}^{n},U_{t}^{n},\hat{\phi})-f(X_{t}^{n},U_{t}^{n};\phi^{\star})\right\|^{2}.$
For $g\in\mathcal{G}_{\star}$, define the offset martingale complexity $M_{T,K}(g)$ as
$M_{T,K}(g)\triangleq\frac{1}{TK}\left(4\sum_{t,k=1}^{T,K}\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right).$
Then the offset basic inequality allows us to bound the empirical risk by the offset martingale complexity as
$\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}\leq M_{T,K}(g).$ (27)
Bounding the offset complexity: We bound the offset complexity using a covering argument.
Let $\mathcal{G}_{\gamma}$ be a minimal cardinality $\gamma$-cover for $\mathcal{G}_{\star}$ with cardinality $\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\gamma)$. Also let $\Pi$ denote the projection of an element $g\in\mathcal{G}_{\star}$ onto $\mathcal{G}_{\gamma}$. We have that,
$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)$ $=\sup_{g\in\mathcal{G}_{\star}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$
$=\sup_{g\in\mathcal{G}_{\star}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n})-g_{\Pi g}(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle+4\langle g_{\Pi g}(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right).$
For any $g\in\mathcal{G}_{\star}$, we may apply the Cauchy-Schwarz inequality and the bound $\left\|g-\Pi g\right\|_{\infty}\leq\gamma$ to bound
$\sup_{g\in\mathcal{G}_{\star}}\frac{1}{TK}\sum_{t,k=1}^{T,K}4\langle g(X_{t}^{n},U_{t}^{n})-g_{\Pi g}(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle$ $\leq 4\gamma\sqrt{\frac{\sum_{t,k=1}^{T,K}\left\|W_{t}^{k}\right\|^{2}}{TK}}\leq 12\gamma\sigma_{w}d_{\mathsf{X}}$
with probability at least $1-\delta$ using the concentration of chi-squared random variables, and the fact that $K\geq 3\log\left(1/\delta\right)$. Additionally note that
$\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\geq\frac{1}{2}\left\|g_{\Pi g}(X_{t}^{n},U_{t}^{n})\right\|^{2}-\left\|g(X_{t}^{n},U_{t}^{n})-g_{\Pi g}(X_{t}^{n},U_{t}^{n})\right\|^{2}\geq\frac{1}{2}\left\|g_{\Pi g}(X_{t}^{n},U_{t}^{n})\right\|^{2}-\gamma^{2}.$
Combining the above inequalities allows us to bound the offset complexity over all parameters in terms of the offset complexity of the finite set of parameters determined by the cover $\mathcal{G}_{\gamma}$.
$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)\leq 12\gamma\sigma_{w}d_{\mathsf{X}}+\gamma^{2}+\frac{\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)}{TK}.$ (28)
We may then in turn bound the offset complexity over a finite hypothesis class using Proposition F.2 of Ziemann et al. (2023). With probability at least $1-\delta$,
$\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$ $\leq 16\sigma_{w}^{2}\log\left(\frac{{\left|\mathcal{G}_{\gamma}\right|}}{\delta}\right)\leq 16\sigma_{w}^{2}\left(d_{\phi}\log\left(\frac{2L_{f}}{\gamma}+1\right)+\log\frac{1}{\delta}\right),$ (29)
where the final inequality followed from (22). Setting $\gamma=\frac{\sigma_{w}}{TK}$, we may combine (28) with (29) to conclude
$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)\leq\frac{12\sigma_{w}^{2}d_{\mathsf{X}}}{TK}+\frac{\sigma_{w}^{2}}{(TK)^{2}}+16\sigma_{w}^{2}\left(d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\leq\frac{32\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right).$
By the burn-in requirement
$K\geq\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{a-1}},$
we may combing the above inequality with (26) and (27) to achieve the desired result. ∎
With this result in hand, we may immediately determine a high probability bound on the parameter estimation error as it appears in the bound in Footnote 3.
Corollary A.1.
Under the setting of Lemma A.1, it follows that with probability at least $1-\delta$,
$\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}\!\leq\!\left(\frac{512\sigma_{w}^{2}}{TN}\left(d_{\mathsf{X}}\!+\!d_{\phi}\log\left(\frac{3L_{f}TN}{\delta}\right)\right)\right)^{2\alpha}.$
For $\alpha<\frac{1}{2}$, the above bound decays at a sub-optimal rate. However, the dependence can be improved by leveraging the Delta method, outlined in the subsequent section.
A.2 Improved Rates via the Delta Method
We can improve upon the estimation error bounds from the previous section by leveraging the smoothness assumptions on the dynamics. In particular, the loss function is locally quadratic in the parameter error once when the parameter error is sufficiently small. Therefore, once we have sufficient data available, the parameter error should scale with the rate of $\frac{1}{TK}$.
We will show that this is the case using an appeal to Taylor’s theorem. In particular, we apply Taylor’s theorem to the optimality conditions of the least squares problem, an approach known as the delta method (Van der Vaart, 2000). Doing so allows us to to show that the parameter error may be expressed in terms of a self-normalized martingale. To express this result, we define the shorthand $Df_{t}^{k}\triangleq D_{\phi}f(X_{t}^{k},U_{t}^{k};\phi^{\star})$. In differentiating the optimality conditions, we encounter an empirical covariance for the covariates defined by the Jacobian of $f$ evaluated at the experiment data:
$\hat{\Sigma}_{K}^{\pi}=\frac{1}{K}\sum_{k=1}^{K}\sum_{t=1}^{T}(Df_{t}^{k})^{\top}Df_{t}^{k}.$ (30)
The superscript $\pi$ denotes the policy that is rolled out to generate $U_{t}^{k}$. Given sufficiently many trajectories, this estimate converges to the expected value $\Sigma^{\pi}$ (defined in (7)). As long as this covariance matrix is positive definite, $\Sigma^{\pi}\succ 0$, we can show that the delta method is successful in improving the estimation rate.
In order to do so, we require sufficiently many trajectories. In addition to the burn-in for the excess risk bound, we require that $N$ is large enough that Corollary A.1 guarantees $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ with high probability. We refer to the burn-in required for this to hold as $K\geq\tau_{\mathsf{small\,error}}(\delta)$, where
$\tau_{\mathsf{small\,error}}(\delta)=\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right).$
Furthermore, we require that the higher order terms ($\mathsf{h.o.t.}$) from the taylor expansion involved in the Delta method are dominated by the lower order terms. For downstream use of the delta method, we allow this to depend on the norm of an arbitrary positive definite matrix $H$. We refer to the required burn-in as $K\geq\tau_{\mathsf{h.o.t.}}(\delta,H)$, where
$\tau_{\mathsf{h.o.t}}(\delta)$ $=\Bigg{(}\frac{8T^{1-\alpha}L_{f}^{2}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
$+\frac{8d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\sqrt{32T\log\frac{4}{\delta}}\Bigg{)}^{\frac{1}{\alpha}}.$
With these burn-in times defined, we may state the result which expresses the parameter estimation error
Lemma A.5.
Let Assumption 2.1 hold and fix $\delta\in(0,\frac{1}{2}]$. Let $H$ be a positive semi-definite matrix. Fix some $\delta\in(0,\frac{1}{4}]$. Consider setting $\hat{\phi}$ as the least squares solution from $K$ episodes of data collected with policy $\pi$ for
$K$ $\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/2),\tau_{\mathsf{small\,error}}(\delta),\tau_{\mathsf{h.o.t}}(\delta,H)\right\}\mathclose{}$
Suppose $\hat{\Sigma}_{K}^{\pi}\succeq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})>0$. Then with probability at least $1-\delta$,
$\hat{\phi}-\phi^{\star}$ $=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k},$
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}$.
As long as $\Sigma^{\pi}\succ 0$ and $K$ is sufficiently large that $\Sigma_{K}^{\pi}$ concentrates to $\Sigma^{\pi}$, we may use the above result to show that $\hat{\phi}_{K}-\phi^{\star}$ behaves like a self-normalized martingale. We may then invoke Theorem 7.2 of Wagenmaker et al. (2021) to bound this self-normalized martingale, leading to Theorem 3.1. Before proving the above lemma, we introduce several preliminary results.
Lemma A.6.
Let $A$ be a symmetric, positive definite matrix. Let $B$ be such that $\left\|B\right\|\leq s\lambda_{\min}(A)$ for $s\leq\frac{1}{2}$. Then $(A+B)^{-1}C=(I+G)A^{-1}C$ for some $G$ satisfying $\left\|G\right\|\leq 2s.$
Proof.
Let $B=U\Sigma V^{\top}$ be the thin singular value decomposition of $B$. By the woodbury matrix idenity
$(A+U\Sigma V^{\top})^{-1}C=\left(I-A^{-1}U(\Sigma^{-1}+V^{\top}A^{-1}U)^{-1}V\right)A^{-1}C.$
Let $G=-A^{-1}U(\Sigma^{-1}+V^{\top}A^{-1}U)^{-1}V$. By submultiiplicativity, $\left\|G\right\|_{\mathsf{op}}\leq\frac{1}{\lambda_{\min}(A)}\left(\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)\right)^{-1}.$ We have that
$\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)$ $\geq\sigma_{\min}(\Sigma^{-1})-\left\|V^{\top}A^{-1}U\right\|=\sigma_{\min}(\Sigma^{-1})-\frac{1}{\lambda_{\min}(A)}.$
Then by noting that $\sigma_{\min}(\Sigma^{-1})=\left\|B\right\|$, we have $\left\|G\right\|_{\mathsf{op}}\leq\left(\frac{\lambda_{\min}(A)}{\left\|B\right\|}-1\right)^{-1}$ The result follows by substituting the bound on $\left\|B\right\|$. ∎
Lemma A.7 (Martingale concentration).
Suppose $S_{1},\dots,S_{T}$ is a real-valued martingale process adapted to the filtration $\mathcal{F}_{1},\dots,\mathcal{F}_{T}$ with conditionally $\sigma^{2}$-sub-Gaussian increments. Then
$\operatorname{\mathbb{P}}\left[S_{T}\geq\rho\right]\leq\exp\left(-\frac{\rho^{2}}{2T\sigma^{2}}\right).$
Proof.
By a Chernoff bound, $\operatorname{\mathbb{P}}\left[S_{T}\geq\rho\right]\leq\inf_{\lambda\geq 0}\exp(-\rho\lambda)\operatorname{\mathbb{E}}\left[\exp\left(\lambda S_{T}\right)\right].$ From the fact that the increments are sub-Gaussian, we have
$\operatorname{\mathbb{E}}\left[\exp\left(\lambda S_{T}\right)\right]$ $=\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}\left[\exp\left(\lambda S_{T}\right)|\mathcal{F}_{T-1}\right]=\operatorname{\mathbb{E}}\left[\exp\left(\lambda S_{T-1}\right)\operatorname{\mathbb{E}}\left[\exp\left(\lambda(S_{T}-S_{T-1})\right)|\mathcal{F}_{T-1}\right]\right]$
[…]
Substituting into the Chernoff bound and optimizing over $\lambda$ provides the result. ∎
We now proceed to prove Lemma A.5
Proof.
By the fact that $\hat{\phi}_{K}$ solves (3),
00 0 $=\left(\nabla_{\phi}\sum_{k=1}^{K}\sum_{t=1}^{T}\left\|X_{t+1}^{k}-f(X_{t}^{k},U_{t}^{k};\phi)\right\|^{2}\right)|_{\phi=\hat{\phi}_{K}}\triangleq\psi(\hat{\phi}_{K}).$
A second order taylor expansion of $\psi(\hat{\phi}_{K})$ about $\phi^{\star}$ provides
$\psi(\hat{\phi}_{K})$ $=\psi(\phi^{\star})+D_{\phi}\psi(\phi)|_{\phi=\phi_{\star}}(\hat{\phi}_{K}-\phi^{\star})+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}],$
for some $\tilde{\phi}$ which is a convex combination of $\phi^{\star}$ and $\hat{\phi}_{K}$. Here $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot,\cdot]$ is a tensor mapping $\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}^{d_{\phi}}$. We write $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot]$ as the tensor mapping $\mathbb{R}^{d_{\phi}}\to\mathbb{R}^{d_{\phi}\times d_{\phi}}$ such that
$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}]=D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}](\hat{\phi}_{K}-\phi^{\star}).$
As long as $D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ is non-singular, then the above two equations imply
$\hat{\phi}_{K}$ $-\phi^{\star}=\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star}).$
Evaluating the derivatives provides
$\psi(\phi^{\star})$ $=\sum_{k=1}^{K}\sum_{t=1}^{T}(Df_{t}^{k})^{\top}W_{t}^{k}$
$D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}$ $=K\hat{\Sigma}_{K}^{\pi}+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{2}f_{t}^{k})[W_{t}^{k}]}_{E_{1}}$
$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ $=\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}]Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})}_{E_{2}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})^{\top}(D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}])^{\top}}_{E_{3}}$
$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{3}f(X_{t}^{k},U_{t}^{k};\tilde{\phi}))[\hat{\phi}_{K}-\phi^{\star},W_{t}^{k}]}_{E_{4}}$
We will bound the operator norms of $E_{1},E_{2},E_{3}$ and $E_{4}$ in order to apply Lemma A.6. We may bound $E_{2}$ and $E_{3}$ directly using Assumption 2.1:
$\left\|E_{2}\right\|_{\mathsf{op}}$ $\leq TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|\mbox{ and }\left\|E_{3}\right\|_{\mathsf{op}}\leq TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|.$
For $E_{1}$ and $E_{4}$, we use a concentration argument. First note that by expressing the $i^{th}$ entry of $W_{t}^{k}$ as $W_{t}^{k}[i]$ and the $i^{th}$ entry of $f_{t}^{k}$ as $f_{t}^{k}[i]$, we have
$E_{1}=\sum_{i=1}^{d_{\mathsf{X}}}\sum_{k=1}^{K}\sum_{t=1}^{T}W_{t}^{k}[i]D^{2}(f_{t}^{k}[i])$
Appealing to a covering argument (cf. Lemma 2.4 of Ziemann et al. (2023)), we may construct a $\frac{1}{2}$-net $\mathcal{N}$ for the unit sphere in $\mathbb{R}^{d_{\mathsf{X}}}$. Then we have $\operatorname{\mathbb{P}}\left[\left\|E_{1}\right\|_{\mathsf{op}}\geq\rho\right]\leq 5^{d_{\mathsf{X}}^{2}}\sup_{u,v\in\mathcal{N}}\operatorname{\mathbb{P}}\left[u^{\top}E_{1}v\geq\frac{1}{4}\rho\right]$. For any $u,v\in\mathbb{R}^{d_{\mathsf{X}}}$,
$u^{\top}E_{1}v=\sum_{i=1}^{d_{\mathsf{X}}}\sum_{k=1}^{K}\sum_{t=1}^{T}W_{t}^{k}[i]u^{\top}D^{2}(f_{t}^{k}[i])v.$
The using boundedness, Assumption 2.1, the quantity $\sum_{t=1}^{T}W_{t}^{k}[i]u^{\top}D^{2}(f_{t}^{k}[i])v$ is a martingale with conditionally $\sigma_{w}^{2}L_{f}^{2}$-sub-Gaussian increments. Then by Lemma A.7,
$\operatorname{\mathbb{P}}\left[\sum_{t=1}^{T}W_{t}^{k}[i]u^{\top}D^{2}(f_{t}^{k}[i])v\geq\frac{\rho}{4}\right]\leq\exp\left(-\frac{\rho^{2}}{32TK\sigma_{w}^{2}L_{f}^{2}}\right).$
Using this bound in the covering argument above, we find
$\operatorname{\mathbb{P}}\left[\left\|E_{1}\right\|_{\mathsf{op}}\geq\rho\right]\leq 5^{d_{\mathsf{X}}^{2}}\exp\left(-\frac{\rho^{2}}{32TK\sigma_{w}^{2}L_{f}^{2}}\right).$
As a result, with probability at least $41-\delta/$ 4,
$\left\|E_{1}\right\|_{\mathsf{op}}\leq d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}.$
Similarly, we may show that with probability at least $41-\delta/$ 4
$\left\|E_{4}\right\|_{\mathsf{op}}\leq d_{\mathsf{X}}\sigma_{w}L_{f}\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\sqrt{32TK\log\frac{4}{\delta}}.$
Union bounding, we find that with probability at least $21-\delta/$ 2,
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq d_{\mathsf{X}}\sigma_{w}L_{f}(1+\left\|\hat{\phi}_{K}-\phi^{\star}\right\|)\sqrt{32TK\log\frac{4}{\delta}}+2TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|.$
By the fact that $K\geq\tau_{\mathbf{Err}}(\delta/2)$, the conditions on Corollary A.1 are satisifed such that with probability at least $21-\delta/$ 2,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}.$
Using the burn-in $K\geq\tau_{\mathsf{small\,error}}(\delta)$, we have that this is at most $1$. Therefore, with probability at least $1-\delta$,
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq 2d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}+2(TK)^{1-\alpha}L_{f}^{2}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
As $K\geq\tau_{\mathsf{h.o.t}}(\delta,H)$, we have that
$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left((H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}K\lambda_{\min}\left(\hat{\Sigma}_{K}^{\pi}\right).$
We may therefore invoke Lemma A.6 with $s=\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}$ to show that
$\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star})=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k},$
for some $G$ satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}.$
A.3 Proof of Theorem 3.1
To proceed with our proof of the main result, we will leverage Lemma A.5. In particular, we require a bound on the self-normalized martingale that arises. We state one such bound due to Wagenmaker et al. (2021) for this purpose. The bound is stated in order to characterize the norm of a self-normalized martingale with respect to an arbitrary positive definite matrix $H$.
Lemma A.8 (Theorem 7.2 of Wagenmaker et al. (2021)).
Suppose $\Sigma^{\pi}\succ 0$ and $H\succeq 0$. Define the event $\mathcal{E}\triangleq\mathopen{}\left\{\left\|\Sigma_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}\right\}\mathclose{})$ for some $\beta\in(0,\frac{1}{4})$. As long as $\mathcal{E}$ holds, the following holds with probability at least $1-\delta$,
$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 5(1+\alpha)\sigma_{w}^{2}\log\frac{6d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma^{\pi}\right)^{-1}\right)}{K},$
where $\alpha=26\beta^{2}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right))$.
Proof.
The proof follows from the proof of Theorem 7.2 of Wagenmaker et al. (2021) by making the following substitutions for $\bm{Z}$ and $\bm{w}$ in the proof:
$\bm{Z}=\begin{bmatrix}Df_{1}^{1}\\ \vdots\\ Df_{T}^{1}\\ \vdots\\ Df_{1}^{K}\\ \vdots\\ Df_{T}^{K}\end{bmatrix}\quad\mbox{ and }\quad\bm{w}=\begin{bmatrix}w_{1}^{1}\\ \vdots\\ W_{t}^{1}\\ \vdots\\ w_{1}^{K}\\ \vdots\\ W_{t}^{K}\end{bmatrix}.$
To guarantee concentration of the empirical covariance (30) to the true covariance (7), we state the following covariance concentration result.
Lemma A.9.
Let $\beta\in(0,\frac{1}{4})$. Suppose
$K\geq\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta)}{\beta^{2}\lambda_{\min}(\Sigma^{\pi})^{2}}.$
Then with probability at least $1-\delta$,
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}).$
Proof.
Let $\mathcal{N}$ be an $\varepsilon$-net of $\mathopen{}\left\{v\in\mathbb{R}^{d_{\phi}}:\left\|v\right\|=1\right\}\mathclose{}$ with $\varepsilon=\frac{1}{4}$. By a covering argument (Lemma 2.5 of Ziemann et al. (2023)),
$\operatorname{\mathbb{P}}\left[\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\geq\rho\right]\leq 9^{d_{\phi}}\max_{v\in\mathcal{N}}\operatorname{\mathbb{P}}\left[{\left|v^{\top}(\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi})v\geq\frac{\rho}{2}\right|}\right].$
Let $\hat{\Sigma}^{\pi,k}=\sum_{t=1}^{T}(Df_{t}^{k})^{\top}Df_{t}^{k}$ and observe that $\hat{\Sigma}_{K}^{\pi}=\sum_{k=1}^{K}\hat{\Sigma}^{\pi,k}$. Observe that for any unit vector $v$, $\operatorname{\mathbb{E}}\left[v^{\top}(\hat{\Sigma}_{K}^{\pi,k}-\Sigma^{\pi})v\right]=0$. Furthermore, using the bound on $\left\|Df_{t}^{n}\right\|_{\mathsf{op}}$ in Assumption 2.1, we have $0\leq v^{\top}\hat{\Sigma}^{\pi,k}v\leq TL_{f}^{2}$. Then by Hoeffding’s inequality (Theorem 2.2.6 of Vershynin (2018)), we have
$\operatorname{\mathbb{P}}\left[{\left|v^{\top}(\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi})v\geq\frac{\rho}{2}\right|}\right]\leq\exp\left(\frac{-\rho^{2}}{2KT^{2}L_{f}^{4}}\right).$
Inverting the bound, we find that with probability at least $1-\delta$,
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\frac{\sqrt{2}TL_{f}^{2}\sqrt{3d_{\phi}+\log(1/\delta)}}{\sqrt{K}}.$
The result follows from the lower bound on $K$. ∎
Using the above concentration result and self-normalized martingale bound, we may prove our main bound on the system identification error, Theorem 3.1.
Proof.
We will invoke the covariance concentration lemma above with the value of $\beta$ chosen according to Lemma A.8. In particular, we define
$\tau_{\mathsf{cov\,conc}}(\delta)=\frac{52T^{2}L_{f}^{4}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)(3d_{\phi}+\log(1/\delta))}{\lambda_{\min}(\Sigma^{\pi})^{2}}.$
If $K\geq\tau_{\mathsf{cov\,conc}}(\delta)$, we may invoke Lemma A.9 to show that with probability at least $41-\delta/$ 4,
$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}),$
where $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}\leq\frac{1}{2}$. As a result, we have that $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$. Given this condition and the burn-in conditions $K\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/16),\tau_{\mathsf{small\,error}}(\delta/4),\tau_{\mathsf{h.o.t.}}(\delta/4)\right\}\mathclose{}$, we satisfy the conditions of Lemma A.5. Therefore, conditioned on the event of covariance concentration, we have that with probability at least $41-\delta/$ 4,
$\hat{\phi}_{K}-\phi^{\star}$ $=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k},$
where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}$. Then by a triangle inequality and submultiplicativity, we have that under the success events of covariance concentration and the delta method,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 2\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}+2\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}$ (31)
In particular, we have reduced the parameter recovery errort to the sum of weighted norms of a self-normalized martingale. Calling on self-normalized martingale machinary Abbasi-Yadkori et al. (2011), Wagenmaker et al. (2021) we may bound each of these norms with high probability. In particular, invoking Lemma A.7 with $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}$, we find that with probability at least $21-\delta/$ 2, the following bounds hold:
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
Substituting these bounds into (31), we find that by union bounding over the success events, we have that with probability at least $1-\delta$,
$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 40\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1}}{K}.$
Appendix B Results for smooth nonlinear systems
We begin by modifying the result of Lemma D.1 of Wagenmaker et al. (2023) that bounds the derivatives of the cost with respect to both the controller and the dynamics parameters.
Lemma B.1.
Under Assumptions 2.1, 2.3, and 2.2 we have that for any $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, the controller cost $\mathcal{J}(\pi^{\theta},\phi)$ is four times differentiable in $\theta$ and $\phi$, and the derivatives satisfy the bound
$\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}(L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}).$
for all $i,j\in\mathopen{}\left\{0,1,2,3\right\}\mathclose{}$, $1\leq i+j\leq 4$.
Proof.
The proof proceeds as in that of Theorem D.1 of Wagenmaker et al. (2023).
In particular, we first define $p_{w}$ to denote the density of the noise, $W_{t}$. Let $p_{\phi,\theta}(\cdot)$ denote the density over the trajectories $(X_{1:T+1},U_{1:T})$ induced by playing controller $\pi^{\theta}$ on the system (1) with parameter $\phi$. Observe that
$p_{\theta,\phi}(X_{1:T+1},U_{1:T})=\prod_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi^{\theta}_{t}(X_{t});\phi^{\star})).$ (32)
We may express our average controller cost as
$J(\pi^{\theta},\phi)=\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
Note that by our noise assumptions and the smoothness assumptions, we may interchange limits and integrals by appealing to the dominated convergence theorem.
Let $\phi_{\bm{m}}=\phi+m_{1}\Delta_{1}^{\phi}+m_{2}\Delta_{2}^{\phi}+m_{3}\Delta_{3}^{\phi}$ and $\theta_{\bm{s}}=\theta+s_{1}\Delta_{1}^{\theta}+s_{2}\Delta_{2}^{\theta}+s_{3}\Delta_{3}^{\theta}$ for $\Delta_{i}^{\theta}$ and $\Delta_{j}^{\phi}$ satisfying $\left\|\Delta_{i}^{\theta}\right\|=1$ and $\left\|\Delta_{j}^{\phi}\right\|=1$ for $i,j=1,2,3$. We will consider the directional derivatives of $\mathcal{J}(\pi^{\theta},\phi)$ by considering the derivatives of $\mathcal{J}(\pi^{\theta_{\bm{s}}},\phi_{\bm{m}})$ with respect to $x_{1},x_{2},x_{3},x_{4}\in\mathopen{}\left\{m_{1},m_{2},m_{3},s_{1},s_{2},s_{3}\right\}\mathclose{}$. By showing that $\mathcal{J}(\pi^{\theta_{\bm{s}}},\phi_{\bm{m}})$ is differentiable with respect to these quantities for any unit vectors $\Delta_{i}^{\theta}$ and $\Delta_{i}^{\phi}$, then $\mathcal{J}(\pi^{\theta},\phi)$ is four times differentiable with respect to $\theta$ and $\phi$. Additionally, we can bound the operator norm of the derivatives by bounding the directional derivatives for all unit vectors $\Delta_{i}^{\theta}$ and $\Delta_{i}^{\phi}$.
To show differentiability, note that
$\frac{d}{dx_{1}}J(\pi^{\theta_{\bm{s}}},\phi_{\bm{m}})=\int\frac{d}{dx_{1}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
$=\int\underbrace{\left(\frac{d}{dx_{1}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)\right)}_{\mbox{Derivative of log density}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
Proceeding to differentiate the above quantity with respect to $x_{2}$, $x_{3}$, and $x_{4}$, provides from the product rule a sum of terms which are of the form above, but with the derivative of the log density replaced by the product of higher order derivatives of the log density.
Recalling the form of the density in (32), we have that
$\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)=\sum_{t=1}^{T}-\frac{1}{2\sigma_{w}^{2}}\left\|X_{t+1}-f(X_{t},\pi^{\theta_{\bm{s}}};\phi_{\bm{m}})\right\|^{2}+T\log\frac{1}{\sqrt{2\pi\sigma_{w}^{2d_{\mathsf{X}}}}}.$
By the differentiability of $f$ with respect to $\phi$ and $u$ from Assumption 2.1 and the differentiability of $\pi^{\theta}$ with respect to $\theta$ from Assumption 2.2, we have that the log density above is four times differentiable in $x_{1},\dots x_{4}$ for any values of $\Delta_{i}^{\theta}$ and $\Delta_{j}^{\phi}$.
To bound the norm of the gradients, observe that
$\frac{d}{dm_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{\phi}f(X_{t},\pi^{\theta;\phi})\Delta_{i}^{\phi}$
$\frac{d}{ds_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{u}f(X_{t},\pi^{\theta};\phi)D_{\theta}\pi^{\theta}\Delta_{i}^{\theta}$
Differentiating further with respect to $m_{i}$ and $s_{i}$ results only in higher order derivatives. Using the operator norm bounds on the derivative of the policy and the dynamics, we may conclude that for $i,j\in\mathopen{}\left\{0,1,2,3\right\}\mathclose{}$ such that $1\leq i+j\leq 4$,
$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)$
We may then in turn bound $\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}$ as
$\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)$
×$\times\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{4}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})$
$\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
×$\times\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})},$
by Cauchy-Schwarz. We have from Assumption 2.3 that
$\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
$=\sqrt{\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]}\leq L_{\mathsf{cost}}.$
[…]
$\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}\leq\mathsf{poly}(T,d_{\mathsf{X}}).$
Therefore our derivative is bounded as
$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1},L_{\mathsf{cost}},d_{\mathsf{X}}\right).$
Given Lemma B.1, the proof of Lemma 2.1 follows from an argument identical to that of Proposition 6 in Wagenmaker et al. (2023), with Lemma B.1 invoked in place of their Lemma D.1. We may also use Lemma B.1 to show the following two results from Wagenmaker et al. (2023).
Lemma B.2 (Lemma D.3 of Wagenmaker et al. (2023)).
Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
$\left\|\mathcal{H}(\phi^{\star})-\mathcal{H}(\phi)\right\|_{\mathsf{op}}\leq C_{\mathsf{Hpert}}\left\|\phi-\phi^{\star}\right\|.$
where
$C_{\mathsf{Hpert}}=\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right).$
Lemma B.3 (Lemma D.4 of Wagenmaker et al. (2023)).
Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
$\left\|\mathcal{H}(\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right).$
The proofs of Footnote 3, Lemma B.2, and Lemma B.3 from the same arguments as the proofs of Lemma D.2, Lemma D.3 and Lemma D.4 of Wagenmaker et al. (2023) with Lemma B.1 taking place of their Lemma D.1, and the conclusion of Lemma 2.1 taking place of their Assumption 13.
Appendix C Modifications to Dynamic Optimal Experiment Design
C.1 Dynamic Optimal Experiment Design
We consider an experiment design procedure, Algorithm 2, which is a modified version of the algorithm from Wagenmaker et al. (2023). The goal of this algorithm is to collect a dataset consisting of $N$ episodes playing exploration policies belonging to $\Pi_{\exp}$ whose Gram matrix $\Lambda$ (15) which approximately minimizes the objective $\Phi(\Lambda)$. It takes as input the objective $\Phi$, an episode budget $N$, an initial exploration policy $\pi^{0}$, a parameter estimate $\hat{\phi}$, and a confidence set for the parameters $\mathcal{B}$. The algorithm plays a sequence of $N^{1/3}$ epochs, each with $N^{2/3}$ episodes. For the first epoch, it plays the initial policy to set an initial value of the Gram matrix $\Lambda_{0}$. Starting from this value, the algorithm performs a variant of gradient descent on the objective $\Phi$, where each epoch is a gradient step. At each iteration, it computes the gradient $\Xi_{n}$ of this objective at the current iterate $\Lambda_{n}$. Rather than taking a step in the direction $\Xi_{n}$, it takes a step in the descent direction constrained to the set of Gram matrices achievable under the dynamics evolve according to (1) when the input is chosen according to some exploration policy $\pi\in\Pi_{\exp}$. It does so by estimating the solution to
$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right],$
where $c^{n}$ is defined in terms of the gradient $\Xi_{n}$ as
$c^{n}(x,u)=\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{n}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}.$
In the above optimization problem, the expectation is taken with respect to the true parameter, which is unknown. However, the confidence ball $\mathcal{B}$ centered at the estimate $\hat{\phi}$ can be used to approximate the solution. In particular, the algorithm applies optimism in the face of uncertainty (Abbasi-Yadkori et al., 2011; Kakade et al., 2020) to select the exploration policy which solves the problem
$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\min_{\phi\in\mathcal{B}}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right].$
Doing so allows us to achieve high probability bounds on the objective value obtained by the dynamic optimal experiment design algorithm (20). The formal version of this is modified from Theorem 3 of Wagenmaker et al. (2023), and is presented in Theorem C.1.
[…]
1:Input: Objective $\Phi$, episodes $N$, initial exploration policy $\pi^{0}$, parameter estimate $\hat{\phi}$, parameter uncertainty set $\mathcal{B}$, smoothness bound $B_{\Phi}$
2:Set $\tilde{N}\leftarrow\lfloor N^{2/3}\rfloor$, $\tilde{K}\leftarrow\lfloor N^{1/3}\rfloor-1$, $\gamma_{n}\leftarrow\frac{1}{n+1}$.
3:Play $\pi^{0}$ for $\tilde{N}$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$,
4:Store $\mathcal{D}_{0}\leftarrow\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,k=1}^{T+1,\tilde{N}}$
5:Form the Gram matrix $\Delta_{0}=\frac{1}{\tilde{N}}\sum_{{t,n=1}}^{T,\tilde{N}}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})$
6:for $k=1,\dots,\tilde{K}$ do
7:     Define $\Xi_{k}\leftarrow\nabla_{\Lambda}\Phi(\Lambda)|_{\Lambda=\Lambda_{k-1}}$ and $c^{k}(x,u)\!\leftarrow\!\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{k}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}$
8:     Set $\pi^{k}\leftarrow\underset{\tilde{\pi}\in\Pi_{\exp}}{\operatorname*{argmin}}\,\underset{\phi\in\mathcal{B}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{k}(X_{t},U_{t})\right]$
9:     Run $\pi^{k}$ for $\tilde{N}$ episodes, collecting $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T,\tilde{N}}$, and update the dataset and Gram matrix as
$\Lambda_{k}$ $\leftarrow(1-\gamma_{k})\Lambda_{k-1}+\frac{\gamma_{k}}{\tilde{N}}\sum_{t,n=1}^{T,\tilde{N}}Df(x,u;\hat{\phi})^{\top}\left(\Xi_{k}\right)Df(x,u;\hat{\phi}),\quad\mathcal{D}_{k}\leftarrow\mathcal{D}_{k-1}\cup\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$
10:Return: $\Lambda_{\tilde{K}}$, $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}$
C.2 Proof of correctness for $\texttt{DOED}_{+}$ Algorithm
The objective of Algorithm 2 is to collect data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{n,t=1}^{N,T}$ from the system to optimize the some objective $\Phi(\Lambda)$, where
$\Lambda=\sum_{n,t=1}^{N,T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}).$
In light of this, we define the set of all possible Gram matrices corresponding to the parameter parameter $\hat{\phi}$:
$\Omega(\phi)$ =$=\bigg{\{}\sum_{t=1}^{T}Df(x_{t},u_{t};\phi)^{\top}Df(x_{t},u_{t};\phi):x_{t}\in\mathbb{R}^{d_{\mathsf{X}}},u_{t}\in\mathbb{R}^{d_{\mathsf{U}}},\forall h\in[H]\bigg{\}}.$
The correctness of the algorithm can be guaranteed under the following regularity assumption on the objective.
Assumption C.1.
[…]
$\left\|\nabla\Phi(\Lambda)-\nabla\Phi(\Lambda^{\prime})\right\|_{\star}\leq\kappa\left\|\Lambda-\Lambda^{\prime}\right\|\,\,\,\,\forall\Lambda,\Lambda^{\prime}\in\Omega(\hat{\phi}).$
Furthermore, suppose there exists some $B_{\Phi}<\infty$ such that for all $\Lambda\in\Omega(\hat{\phi}),x\in\mathbb{R}^{d_{\mathsf{X}}},u\in\mathbb{R}^{d_{\mathsf{U}}}$,
$\left\|Df(x,u,\hat{\phi})\nabla\Phi(\Lambda)Df(x,u,\hat{\phi})^{\top}\right\|\leq B_{\Phi}.$
This leads us to the following result, characterizing the optimality of the data collected via the dynamic optimal experiment design procedure.
Theorem C.1 (Modification of Theorem 3 of Wagenmaker et al. (2023)).
Suppose Assumptions 2.1 and C.1 hold. Additionally, suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$. Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfies
$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$\leq\frac{2\kappa T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}\!+\!B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}\!+\!\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right).$
To prove the above result, we first present an extension of the simulation lemma from Kakade et al. (2020) to the setting of smooth nonlinear systems.
Lemma C.1 (Self-Bounding, Simulation Lemma: Modified from Kakade et al. (2020)).
Suppose Assumption 2.1 holds. For any policy $\pi$, model parameters $\phi$ and non-negative cost $c$ satisfying $c<1$, it holds that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi-\phi^{\star}\right\|.$
Proof.
The proof follows as in the proof of Lemma B.3 of Kakade et al. (2020). Consider a state/input pair $(X_{t},U_{t})$ from rolling out the policy $\pi$ on the system $\phi^{\star}$. Where Kakade et al. (2020) apply the bound on the difference in means between two Gaussian distributions, we use the Gaussian distributions describing the next state update from $X_{t},U_{t}$ under the model described by $\phi$ and by $\phi^{\star}$: $\mathcal{N}(f(X_{t},U_{t};\phi^{\star}),\sigma_{w}^{2})$ and $\mathcal{N}(f(X_{t},U_{t};\phi),\sigma_{w}^{2})$. This results in a bound on the quantity of interest as
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]$
$\leq\sqrt{T\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]}\sqrt{\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2},1\right\}\mathclose{}\right]}.$
Using the fact that the cost is bounded by one, we have
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]\leq T^{2}$
For the remaining quantity, we may bound
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi\right\|,1\right\}\mathclose{}\right]$ $\leq\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2}\right]$
$=\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|Df(X_{t},U_{t},\tilde{\phi})(\phi^{\star}-\phi)\right\|^{2}\right]$
$\leq\frac{TL_{f}^{2}}{\sigma_{w}^{2}}\left\|\phi^{\star}-\phi\right\|^{2}.$
The equality follows by a first order Taylor expansion of $f$ about $\phi^{\star}$. Here $\tilde{\phi}$ is a convex combination of $\phi$ and $\phi^{\star}$. The final inequality follows from Assumption 2.1. Combining these results concludes the proof. ∎
We now proceed to prove Theorem C.1.
Proof.
The proof follows by modifying the proof of Theorem 3 in Wagenmaker et al. (2023). In place of the regret minimization algorithm used by Wagenmaker et al. (2023) to estimate the solution to
$\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right],$
we use the policy synthesized directly via optimism over the parameter estimate set $\mathcal{B}$. Let $\phi_{m}$ be the parameter value under which the optimistic policy was synthesized and let $\pi^{m}$ be the optimistic policy. By our assumption that $\phi^{\star}\in\mathcal{B}$, the optimistic policy synthesis provides that
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right].$
By Lemma C.1, we have that
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi^{\star}-\phi^{m}\right\|.$
By the fact that $D(\mathcal{B})$ is the diameter of $\mathcal{B}$, it follows that the above quantity is bounded as
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}D(\mathcal{B}).$
Substituting this bound in place of the bound form the regret minimization step in the proof of Theorem 3 of Wagenmaker et al. (2023), and carrying out the remaining steps from the proof yields that with probability at least $1-\delta$,
$\Phi(\Delta_{\tilde{K}})-\operatorname*{argmin}_{\pi\in\Pi_{\exp}}\Phi(\Sigma^{\pi})\leq\frac{2\beta T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}+B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log(2\tilde{K}/\delta)}{\tilde{N}}}\right).$
We now consider the correctness of $\texttt{DOED}_{+}$ as it is invoked in 8 by verifying that the conditions to apply the above result are met.
Lemma C.2 (Modified from Corollary 1 of Wagenmaker et al. (2023)).
Let Assumption 2.1 hold. Suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$. Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$ for some $\lambda>0$. Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$. Then with probability at least $1-\delta$, the data collected at the final round of DOED satisfies
$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{N}+1)}{\lambda^{3}(\tilde{N}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{N}}{\delta}}{\tilde{K}}}\right).$
Proof.
The result follows from Theorem C.1. The smoothness parameters of Assumpton C.1 under the chosen objective are shown in Corollary 1 of Wagenmaker et al. (2023) to be bounded as $B_{\Phi}\leq\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}$ and $\kappa\leq\frac{2T^{2}L_{f}^{4}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{3}}$. ∎
Lemma C.3.
Suppose Assumptions 2.1 and 2.6 hold. Additionally suppose that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{4TL_{f}^{2}}$. Then there exists an exploration policy $\pi\in\Pi_{\exp}$ such that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]\geq\frac{\lambda_{\min}^{\star}}{2}.$
Proof.
The result follows by observing that there exists a policy such that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}.$
by Assumption 2.6. Then
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]-\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]$
$=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right].$
By a first order Taylor expansion of $Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})$ about $\phi^{\star}$, and using the boundedness conditions of Assumption 2.1, we can bound
$\left\|\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right]\right\|_{\mathsf{op}}\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{2}.$
Lemma C.4.
Assume that Assumptions 2.1 and 2.6 hold. Additionally, suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}}\left\|\phi_{1}-\phi_{2}\right\|$. Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$. Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$. Let $\delta\in(0,\frac{1}{2}]$. Suppose $\lambda\leq\frac{\lambda_{\min}(H)\lambda_{\min}^{\star}}{8\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$, $D(\mathcal{B})\leq\frac{R}{N^{\alpha}}$ and
$N\geq\max\mathopen{}\left\{\left\lceil\frac{4\lambda_{\min}^{\star}T^{4}L_{f}^{8}\log(N+1)}{\lambda^{3}d_{\phi}}\right\rceil^{3},\left\lceil\frac{8\lambda_{\min}^{\star}TL_{f}^{2}\sqrt{2\log(2N/\delta)}}{\lambda^{2}d_{\phi}}\right\rceil^{3},\left(\frac{4T^{3}L_{f}^{3}R\lambda_{\min}^{\star}}{\lambda^{2}\sigma_{w}d_{\phi}}\right)^{1/\alpha}\right\}\mathclose{}$
Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfy
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]\right)$ (33)
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
and
$\lambda_{\min}\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)\geq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}.$
Proof.
By Lemma C.2, we have that with probability at leat $1-\delta$,
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(N+1)}{\lambda^{3}(\lfloor N^{1/3}\rfloor+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}R}{N^{\alpha}\sigma_{w}}+\sqrt{\frac{8\log\frac{2N}{\delta}}{\lfloor N^{2/3}\rfloor}}\right)$
$\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}},$
where the final inequality follows from the lower bound on $N$. By Assumption 2.6 paired with Lemma C.3, we have that
$\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$=\operatorname{\mathrm{tr}}\left(\mathcal{H}\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]+\lambda I\right)^{-1}\right)\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}}.$
We may lower bound
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)$
$\geq\frac{\lambda_{\min}\left(\mathcal{H}\right)}{\lambda_{\min}\left(\frac{1}{\tilde{N}\tilde{K}}\sum_{t,n=1}^{T,(\tilde{N}\times\tilde{K})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)+\lambda}.$
We conclude by combining these inequalities, and using the assumption that $\lambda\leq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$. ∎
Lemma C.5.
Consider a policy $\pi\in\Pi_{\exp}$. Suppose $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\min\mathopen{}\left\{\frac{\tilde{\lambda}}{4TL_{f}^{2}},1\right\}\mathclose{}$ for some $\tilde{\lambda}>0$. Define
$\Sigma^{\pi}_{\hat{\phi}}=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right],$
and suppose that $\Sigma^{\pi}_{\hat{\phi}}\succeq\tilde{\lambda}I$. Then $\Sigma^{\pi}\succeq\frac{\tilde{\lambda}}{2}I$, and
  1. 1.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}\right)^{-1}_{\hat{\phi}}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
  2. 2.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
Proof.
To begin, we observe that by a first order Taylor expansion of $Df(x,u;\phi^{\star})$ about $\hat{\phi}$, we have that
$\Sigma^{\pi}$ $\succeq\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]+G,$
where $\left\|G\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$ by the boundedness assumptions on the derivatives of $f$ from Assumption 2.1. By the assumption that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\tilde{\lambda}}{4TL_{f}^{2}}$, it holds that $\left\|G\right\|\leq\frac{\tilde{\lambda}}{2}$. Then it follows that $\lambda_{\min}(\Sigma^{\pi})\geq\lambda_{\min}(\Sigma^{\pi}_{\hat{\phi}})-\left\|G\right\|\geq\frac{\tilde{\lambda}}{2}$.
We proceed to bound $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|$. We immediately have that $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\frac{1}{\lambda_{\min}\left(\Sigma^{\pi}\right)\lambda_{\min}\left(\Sigma^{\pi}_{\hat{\phi}}\right)}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq\frac{2}{\tilde{\lambda}^{2}}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|$. By Taylor expanding $Df(x,u;\hat{\phi})$ about $\phi^{\star}$, we have from the boundedness assumptions of Assumption 2.1 that $\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|+TL_{f}\left\|\hat{\phi}-\phi^{\star}\right\|^{2}\leq 3TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$. Furthermore, by Lemma B.2, $\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
To conclude, we bound
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\frac{6TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|}{\tilde{\lambda}^{2}},$
and
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\frac{d_{\phi}}{\tilde{\lambda}}\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|$
$\leq\frac{d_{\phi}}{\tilde{\lambda}}\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
Combining these inequalities with the triangle inequality provides the desired result. ∎
Appendix D Proof of Main Excess Cost Bound, Theorem 3.2
Armed with the above results, we are ready to prove Theorem 3.2.
See 3.2
Proof.
We have that $\frac{N}{4}\geq\tau_{\mathbf{Err}}(\delta/4)$. Then by Assumption 2.5, we may invoke Corollary A.1 to show that with probability at least $41-\delta/$ 4, $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$, for $R$ defined in (17). This in turn implies that ϕsuperscriptitalic-$\phi^{\star}\in\mathcal{B}^{-}$ for $B^{-}$ define in (16). The diameter of the ball is given as $D(\mathcal{B}^{-})=\frac{2R}{N^{\alpha}}$. Denote the event that ϕsuperscriptitalic-$\phi^{\star}\in\mathcal{B}^{-}$ as $\mathcal{E}_{\mathsf{coarse}}$. We condition on the event $\mathcal{E}_{\mathsf{coarse}}$ for the remainder of the proof.
Let $C_{\mathsf{Hpert}}$ be the system theoretic constant defined in Lemma B.2. As long as $\frac{N}{4}\geq\frac{2RC_{\mathsf{Hpert}}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}^{\frac{1}{\alpha}}$, it follows that
$\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}{2C_{\mathsf{Hpert}}}.$ (34)
Then by Lemma B.2, we have $\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\geq\frac{1}{2}\lambda_{\min}(\mathcal{H}(\phi^{\star}))$ and $\left\|\mathcal{H}(\hat{\phi}^{-})\right\|\leq 2\left\|\mathcal{H}(\phi^{\star})\right\|$.
Condition on the event $\mathcal{E}_{\mathsf{coarse}}$. By the choice of $\lambda$ in 5 in Algorithm 1, the assumption that
$\frac{N}{4}\geq\tau_{\mathsf{PD\,data}}(\delta/4),$
where
$\tau_{\mathsf{PD\,data}}(\delta)=\max\mathopen{}\left\{\left\lceil\frac{4\lambda_{\min}^{\star}T^{4}L_{f}^{8}\log(N+1)}{\lambda^{3}d_{\phi}}\right\rceil^{3},\left\lceil\frac{8\lambda_{\min}^{\star}TL_{f}^{2}\sqrt{2\log(2N/\delta)}}{\lambda^{2}d_{\phi}}\right\rceil^{3},\left(\frac{4T^{3}L_{f}^{3}R\lambda_{\min}^{\star}}{\lambda^{2}\sigma_{w}d_{\phi}}\right)^{1/\alpha}\right\}\mathclose{},$
and Assumption 2.6, we may guarantee that with probability at least $1-\frac{\delta}{4}$ the success event of Lemma C.4 holds. Denote this event $\mathcal{E}_{\mathsf{doed}}$. $\texttt{DOED}_{+}$ collects $\tilde{K}\tilde{N}$ episodes of data data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,k=1}^{T+1,(\tilde{N}\times\tilde{K})}$, where $\tilde{K}=\lfloor\frac{N}{4}^{1/3}\rfloor$ and $\tilde{N}=\lfloor\frac{N}{4}^{2/3}\rfloor$. Under the success event, this collected data satisfies
$\lambda_{\min}\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)\geq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{4\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}.$ (35)
By the law of total expectation, the mixture policy $\pi^{\mathsf{mix}}$ defined in 9 of Algorithm 1 satisfies
$\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ $=\frac{1}{2}\operatorname{\mathbb{E}}_{\pi^{0},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ (36)
$+\frac{1}{2\tilde{K}}\sum_{i=1}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$
$\succeq\frac{1}{2(\tilde{N}+1)}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right].$
It follows from Lemma C.10 of Wagenmaker et al. (2023) that the empirical covariance of the data collected from Algorithm 2 concentrates to the true covariance. In particular, with probability at least $41-\delta/$ 4,
$\left\|\frac{1}{\tilde{K}+1}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]-\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right\|$ (37)
$\leq d_{\phi}TL_{f}^{2}\sqrt{\frac{8d_{\phi}\log\left(1+8\sqrt{\tilde{K}\tilde{N}}\right)+8\log(4/\delta)}{\tilde{K}\tilde{N}}}.$
Denote this concentration event as $\mathcal{E}_{\mathsf{doed\,conc}}$. Under the burn-in requirement,
$\frac{N}{4}\geq\tau_{\mathsf{doed\,conc}}(\delta)=\left\lceil\left(\frac{16\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}\right)^{2}d_{\phi}^{4}T^{2}L_{f}^{4}\left(8d_{\phi}\log\left(1+8\sqrt{N/4}\right)+8\log(4/\delta)\right)\right\rceil^{3},$ (38)
this event implies
$\frac{1}{\tilde{K}+1}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ $\succeq\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})$
$-\frac{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}d_{\phi}^{2}}I.$
Combining this inequality with (36) and (35) and using the fact that $\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\geq\frac{1}{2}\lambda_{\min}(\mathcal{H}(\phi^{\star}))$ and that $\left\|\mathcal{H}(\hat{\phi}^{-})\right\|\leq 2\left\|\mathcal{H}(\phi^{\star})\right\|$, we find that
$\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]\succeq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}I.$ (39)
Denote the event that Algorithm 2 collects data satisfying (35) and (37) as $\mathcal{E}_{\mathsf{doed\,covariance}}$. As long as
$\frac{N}{4}\geq\left(\frac{64RTL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}\right)^{\frac{1}{\alpha}},$
it follows that
$\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)}{4TL_{f}^{2}}$ (40)
from the bound $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$. Then by Lemma C.5,
$\Sigma^{\pi_{\mathsf{mix}}}\succeq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{32\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}I\succ 0.$
Additionally, the policy $\pi_{\mathsf{mix}}$ is $(2C_{\mathsf{Loja}},\alpha)$-Lojasiewicz. This can be seen by noting that by the law of total expectation. In particular, for all $\phi\in\mathbb{R}^{d_{\phi}}$,
$\mathbf{Err}_{\pi_{\mathsf{mix}},\phi^{\star}}(\phi)\geq\frac{1}{2}\mathbf{Err}_{\pi^{0},\phi^{\star}}(\phi).$
Then under the events $\mathcal{E}_{\mathsf{coarse}}$, $\mathcal{E}_{\mathsf{doed}}$ and $\mathcal{E}_{\mathsf{doed\,covariance}}$ in conjunction with the burn-in requirement
$\frac{N}{4}\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/32),\tau_{\mathsf{small\,error}}(\delta/16),\tau_{\mathsf{h.o.t}}(\delta/16),\tau_{\mathsf{cov\,conc}}(\delta/16)\right\}\mathclose{},$
the conditions are set to apply Theorem 3.1 on the data collected by rerunning $\pi^{\mathsf{mix}}$ for $\frac{N}{4}$ episodes. Doing so ensures that with probability at least $41-\delta/$ 4,
$\left\|\hat{\phi}^{-}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}\leq 160\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})}{N}.$ (41)
Denote this event as $\mathcal{E}_{\mathsf{ID}}$.
Then under the event $\mathcal{E}_{\mathsf{coarse}}\cap\mathcal{E}_{\mathsf{doed\,conc}}\cap\mathcal{E}_{\mathsf{ID}}$, by using the covariance bound (39) and the estimation bound (40), it follows from Lemma C.5
$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})$ $\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)^{-1}\right)$ (42)
$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\left\|\hat{\phi}-\phi^{\star}\right\|.$
By the lower bound (39), and the choice of $\lambda$ in 5 in Algorithm 1, it follows that
$\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)^{-1}\right)$ (43)
$\leq 2\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$
$\leq 4\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})+\lambda I\right)^{-1}\right)$ ((37) combined with (38))
$\leq 4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$
$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)$ ( (33) holds under the event $\mathcal{E}_{\mathsf{doed}}$)
By invoking Lemma C.5,
$\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$ (44)
[…]
$\leq\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star})}\right)\left\|\hat{\phi}-\phi^{\star}\right\|,$
where the final inequality follows from Lemma C.5. Combining (42) with (43) and (44) and using the bound $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$ yields
$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})$ $\leq 4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)$ (45)
$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{R}{N^{\alpha}}$
$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)$
As a result of the above identfiication bounds, the excess cost of the returned policy $\pi^{\star}(\hat{\phi}^{+})$ is bounded via Footnote 3 as
$\mathcal{J}(\phi^{+})-\mathcal{J}(\phi^{\star})$ $\leq\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})}+C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{3}$
$\leq\left(1+\frac{C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})}$
[…]
$\leq\left(1+\frac{2C_{\mathsf{cost}}R}{N^{\alpha}\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{160\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}}{N}\Bigg{(}4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)$
$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{R}{N^{\alpha}}$
$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)\Bigg{)}.$
To conclude, we leverage the burn-in condition
$\frac{N}{4}\geq\max\Bigg{\{}\Bigg{\lceil}\left(\frac{4T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(N+1)\lambda^{-3}+2TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\lambda^{-2}\left(2T^{2}L_{f}R\sigma_{w}^{-1}+\sqrt{8\log\frac{2N}{\delta}}\right)+C_{\mathsf{H\,pert}}R}{4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)}\right)\bigg{\rceil}^{3/\alpha},$
$\left(2\frac{C_{\mathsf{cost}}R}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)^{1/\alpha}\Bigg{\}},$
where
$C_{\mathsf{H\,pert}}=\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)$
Under this condition, the excess cost bound simplifies to
$\mathcal{J}(\phi^{+})-\mathcal{J}(\phi^{\star})$ $\leq\frac{2560\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}}{N}\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right).$
Appendix E Aditional Experiments and Experiment Details
In Section E.1, we complement our illustrative example with an implementation of ALCOI on a toy physical system. In Section E.2 we provide further experimental details.
E.1 Cartpole Experiments
We additionally run our proposed algorithm on the cartpole system defined by the dynamics
$(M+m)(\ddot{p}+b_{p}\dot{p})+m\ell\cos(\theta)(\ddot{\theta}+b_{\theta}\dot{\theta})$ $=m\ell\dot{\theta}^{2}\sin(\theta)+u$
$m\cos(\theta)(\ddot{p}+b_{p}\dot{p})+m\ell(\ddot{\theta}+b_{\theta}\dot{\theta})$ $=mg\sin(\theta),$
where $p$ is the position of the cart, $\theta$ is the angle of the pole from the upright position, and $u$ is the control input. Here, $M$ represents the mass of the cart, $m$ the mass of the pole, $\ell$ the length of the pole, $g$ the acceleration due to gravity, $b_{x}$ is the friction coefficient for the cart, and $b_{\theta}$ is the friction coefficient for the pole. We discretize the system using the Euler approach using a timestep of $dt=0.1$. We simluate the discretized system under additive zero mean Gaussian noise with covariance $0.01I_{4}$. The unknown parameters are $\phi_{\star}=\begin{bmatrix}M,m,\ell,g,b_{x},b_{\theta}\end{bmatrix}=[1,0.1,1,10,0.5,0.5]$ For every episode, the system starts from the hanging position with state $\begin{bmatrix}0&0&\pi&0\end{bmatrix}$. The desired behavior is to swing the pole to the upright position with the cart positioned at the origin within the time horizon of $T=30$ timesteps. This is described by the cost functions $c_{t}(x,u)=0$, $c_{T+1}(x)=\left\|x\right\|^{2}$.
To simplify the policy synthesis, rather than solving for the certainty equivalent policy which directly minimize the cost, we deploy energy shaping controllers which switch to an LQR controller about the upright position (Tedrake, 2009). In particular, the certainty equivalent controller under estimates $\hat{\phi}=\begin{bmatrix}\hat{M},\hat{m},\hat{\ell},\hat{g},\hat{b}_{x},\hat{b}_{\theta}\end{bmatrix}$ selects the input as follows. If ${\left|\theta\right|}<\pi/4$, set $u$ according to the LQR controller synthesized by linearizing the dynamics under the estimated parameters about the upright equilibrium point with weights $Q=I$, $R=1$. Otherwise, set
$u$ $=(M+m-m\cos(\theta))\ddot{x}_{d}+mg\cos(\theta)\sin(\theta)-m\ell\dot{\theta}^{2}\sin(\theta)$ (46)
$\ddot{x}_{d}$ $=5\dot{\theta}\cos(\theta)\left(\frac{1}{2}m\ell^{2}\dot{\theta}^{2}2+mg\ell\cos(\theta)-mg\ell\right)-0.01p-0.01\dot{p}.$ (47)
During exploration, the input energy budget is restricted to $0.1T$.
The comparison between the three approaches, random exploration, approximate $A$-optimal exploration, and ALCOI is shown in Figure 3. As in the illustrative example, we see the benefit of control-oriented exploration over both random exploration.
Figure 3: Excess cost versus number of exploration episodes for the proposed control-oriented identification procedure, approximate $A$-optimal design, and random exploration for the cartpole swing-up task. The mean over $100$ runs is shown, and the standard error is shaded.
E.2 Implementation Details
For both the cartpole experiments and the illustrative example, the initial exploration policy used by ALCOI and approximate $A$-optimal exploration is set to the random exploration policy.
We deviate from the proposed algorithm in the following respects.
  • Rather than running the initial policy for $\lfloor N/4\rfloor$ episodes and the algorithm $\texttt{DOED}_{+}$ with a budget of $\lfloor N/4\rfloor$ before running the mixture policy for $\lfloor N/4\rfloor$ episodes, we run the initial policy for $\lfloor N/2\rfloor$ episodes and the algorithm $\texttt{DOED}_{+}$ with a budget of $\lfloor N/2\rfloor$ episodes, and use all of the collected data to fit the refined estimate.
  • We treat the regularization parameter $\lambda$ as a hyperparameter, which we fix to $0.001$ in all experiments, rather than setting it according to 5 from Algorithm 1.
  • The approach of optimism in the face of uncertainty from 8 of Algorithm 2 is replace by certainty equivalent synthesis of the exploration policy. As the policy class in both experiments consists of all trajectory dependent policies with bounded input energy, we approximate this certainty equivalent synthesis via a receding horizon control procedure. In particular, at time step $t$ if the current state is $x_{t}$, the learner solves the problem:
    $\underset{\tilde{x}_{t},\dots,\tilde{x}_{T+1},\tilde{u}_{t},\dots,\tilde{u}_{T}}{\mathsf{minimize:}}$ $\quad\sum_{k=t}^{T}c(\tilde{x}_{k},\tilde{u}_{k})$
    $\mathsf{s.t.:}$ $\quad\tilde{x}_{k+1}=f(\tilde{x}_{k},\tilde{u}_{k};\hat{\phi}),\quad\tilde{x}_{t}=x_{t}\quad\sum_{k=1}^{t}\left\|u_{k}\right\|^{2}+\sum_{k=t+1}^{T}\left\|\tilde{u}_{k}\right\|^{2}\leq\mathsf{total\,budget},$
    where $\hat{\phi}$ is the learner’s current esimate of the system parameters and $(\tilde{x}_{k}$, $\tilde{u}_{k})$ constitute the planned trajectory. The learner plays the input $\tilde{u}_{t}$, on the true system, and re-evaluates the optimal control after observing the next state. The solution to the optimization problem defining the above receding horizon controller is computed via random shooting with a warm start. In particular, before applying each control input, $100$ new control sequences are generated from a standard normal distribution, and are normalized to satisfy the energy constraints. The seqeuence which minimizes the objective is then selected. This procedure is warm-started with the solution at the previous timestep by setting that as an additional candidate solution.
  • The certainty equivalent policies intended to minimize the objective $\mathcal{J}(\pi,\phi^{\star})$ are replaced with feedback linearization policies defined in terms of the parameter estimate. In particular, for the illustrative example and cartpole example, the policies are synthesized as
    1. 1.
      Illustrative example: $\hat{\pi}(\hat{\phi})=K\left(X_{t}-\begin{bmatrix}-5.5\\ 0\end{bmatrix}\right)-\sum_{i=1}^{4}\psi(X_{t}-\hat{\phi}^{(i)})$, where $K$ is the LQR controller synthesized via a linearization of the dynamics under the parameter $\hat{\phi}$.
    2. 2.
      Cartpole: the swing up controller is defined using feedback linearization as in (46).
    Note that a consequence of this is that there may exist policies in the policy class $\Pi_{\star}$ which achieve lower cost than the policy defined in terms of the optimal parameter $\phi_{\star}$.
The first modification is made because discarding the half of the data is wasteful. It is included in the algorithm only in order to simplify the analysis, as the constant factor of $2$ is not a concern of this analysis. The second modification is made to avoid requiring knowledge of the parameter $\lambda_{\min}^{\star}$ by the algorithm. The third modification is made to avoid the computational intractability of OFU. The sampling based approach is used since the optimization problem defining the receding horizon controller is intractable. The final modification is made to avoid the computational burden of solving the policy optimization problem via a Bellman equation in both policy synthesis and in computation of the model-task Hessian. These modifications result in a gap between the theoretical guarantees, and the numerical validation. Future work could reduce this gap, e.g. by removing the approach of optimism in the face of uncertainty from the analysis.

### v2

@@
 $X_{t+1}=f(X_{t},U_{t};\phi^{\star})+W_{t},\quad t=1,\dots T,$ (1)
-with state $X_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{X}}}$, input $U_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{U}}}$, and $d_{\mathsf{X}}$-dimensional noise $W_{t}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{w}^{2}I)$.
+with state $X_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{X}}}$, input $U_{t}$ assuming values in $\mathbb{R}^{d_{\mathsf{U}}}$, and $d_{\mathsf{X}}$-dimensional noise $W_{t}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{w}^{2}I)$ for some $\sigma_{w}>0$.
 For simplicity, we assume $X_{1}=0$.
 Here, $f$ is the dynamics function, which depends on an unknown parameter $\phi^{\star}\in\mathbb{R}^{d_{\phi}}$.
-We assume that there exists some positive $B$ such that $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$
+We assume that there exists some positive $B$ such that $\left\|\phi^{\star}\right\|\leq B-1$ and $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$ satisfying $\left\|\phi\right\|\leq B$.
 We study a learner whose objective is to determine a policy $\pi=\mathopen{}\left\{\pi_{t}\right\}\mathclose{}_{t=1}^{T}$ from a policy class $\Pi^{\star}$ to minimize the cost $\mathcal{J}(\pi,\phi^{\star})$, where
 $\mathcal{J}(\pi,\phi)=\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right].$ (2)
+The functions $c_{t}$ are known stage costs.
 The superscript on the expectation denotes that the dynamics (1) are rolled out under parameter $\phi$, while the subscript denotes that the system is played in closed-loop under the feedback control policy $U_{t}=\pi_{t}(X_{1},U_{1},\dots,X_{t-1},U_{t-1},X_{t})$ for $t=1,\dots,T$.
@@
 Using this dataset, the learner may identify the dynamics of the system by solving
-$\hat{\phi}\in\operatorname*{argmin}_{\phi\in\mathbb{R}^{d_{\phi}}}\sum_{n=1}^{N}\sum_{t=1}^{T}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi)\right\|^{2}.$ (3)
+$\hat{\phi}\in\operatorname*{argmin}_{\phi\in\mathbb{R}^{d_{\phi}},\left\|\phi\right\|\leq B}\sum_{n=1}^{N}\sum_{t=1}^{T}\left\|X_{t+1}^{n}-f(X_{t}^{n},U_{t}^{n};\phi)\right\|^{2}.$ (3)
 Solving this problem provides a parameter estimate which is an effective predictor under the distribution of states and inputs seen during the experimentation.
@@
 Assumption 2.3.
-(Bounded Costs) There exists some $r_{\mathsf{cost}}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, and all $\pi\in\Pi^{\star}$, we have $\mathbb{E}_{\pi,\phi}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]\leq L_{\mathsf{cost}}$.
-Due to the smoothness assumptions on the dynamics (Assumption 2.1) and the policy class (Assumption 2.2), the above assumption holds if the stage costs $c_{t}$ are smooth.
-We additionally suppose that $\Pi^{\star}$ is not overparametrized.
+(Regular costs) The stage costs $c_{t}$ are three times differentiable and $\left\|D_{u}^{(i)}c_{t}(x,u)\right\|_{\mathsf{op}}\leq L_{\mathsf{cost}}$ for $i=1,\dots,3$, $\theta\in\mathbb{R}^{d_{\theta}}$, $(x,u)\in\mathbb{R}^{d_{\mathsf{X}}+d_{\mathsf{U}}}$.
+There exists some $r_{\mathsf{cost}}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\mathsf{cost}}(\phi^{\star}))$, and all $\pi\in\Pi^{\star}$, we have $\mathbb{E}_{\pi}^{\phi}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]\leq L_{\mathsf{cost}}$.
+We additionally assume that the certainty equivalent policy is a smooth function of the dynamics parameter in a neighborhood around the optimal parameter.
 Assumption 2.4.
-The global minimum of $\mathcal{J}(\pi^{\theta},\phi^{\star})$, $\theta^{\star}(\phi^{\star})$, is unique and satisfies $\nabla^{2}_{\theta}\mathcal{J}(\pi^{\theta},\phi^{\star})|_{\theta=\theta^{\star}(\phi^{\star})}\succ 0$.
-The above condition requires that there are a unique set of optimal control parameters for the true system.
-[^2] To gain an intuition for this condition, observe that it holds for linear quadratic regulators as long as the input penalty matrix is positive definite.
-Under the above assumptions, we may show that the certainty equivalent policy is smooth near the optimal solution.
-Lemma 2.1 (Proposition 6 of Wagenmaker et al. (2023)).
 There exists some $r_{\theta}(\phi^{\star})>0$ such that for all $\phi\in\mathcal{B}(\phi^{\star},r_{\theta}(\phi^{\star}))$,
@@
   • $\theta^{\star}(\phi)$ is three times differentiable in $\phi$ and $\left\|D_{\phi}^{i}\theta^{\star}(\phi)\right\|_{\mathsf{op}}\!\leq\!L_{\pi^{\star}}$ for some $L_{\pi^{\star}}>0$ and $i\in\mathopen{}\left\{1,2,3\right\}\mathclose{}$.
+[…]
 In order to bound the parameter recovery error in terms of the prediction error, additional identifiability conditions are needed.
@@
 To bypass this limitation, we assume that some policy in the exploration class satisfies a persistence of excitation condition.
-Persistence of excitation for a nonlinear dynamical system involves the positive definiteness of the following matrix:
-$\Sigma^{\pi}\triangleq\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right],$ (7)
-where $Df(X_{t},U_{t},\phi^{\star})$ is the Jacobian of $f$ with respect to $\phi$ evaluated at $\phi^{\star}$.
+This condition can be expressed by first defining the Fisher information matrix for a parameter $\phi$ and a policy $\pi$ as
+$\mathbf{FI}^{\pi}(\phi)\!\triangleq\!\frac{\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}\!Df(X_{t},U_{t},\phi)^{\top}\!Df(X_{t},U_{t},\phi)\right]}{\sigma_{w}^{2}},$ (7)
+where $Df(X_{t},U_{t},\phi)$ is the Jacobian of $f$ with respect to $\phi$.
+The Fisher information measures the signal-to-noise ratio of the data collected from an episode of interaction with the system under exploration policy $\pi$.
+With this definition, persistance of excitation is equivalent to the positive definiteness of the matrix $\mathbf{FI}^{\pi}(\phi^{\star})$.
 Assumption 2.6.
 There exists a policy $\pi\in\Pi_{\exp}$ for which
-$\operatorname{\mathbb{E}}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}>0,$
-Finally, we make an assumption that all models parameters are relevant to perfect control.
-Assumption 2.7.
-The model-task Hessian is positive definite: $\mathcal{H}(\phi^{\star})\succ 0$.
-The above assumption is quite strong, and does not hold for common baseline problems such as the linear-quadratic regulator with fully unknown dynamics matrices.
-However, this assumption is made primarily for ease of exposition.
-It can be removed in the linear setting at the expense of an algorithm which is more complicated to express concisely (Wagenmaker et al., 2023).
-The same approach used to bypass this condition in the linear setting can be easily extended to the nonlinear setting considered in this paper.
-We do not formalize this here, although we will in future work.
+$\mathbf{FI}^{\pi}(\phi^{\star})\succeq\mu I\succ 0.$
 3 Proposed Algorithm and Main Result
 The above smoothness assumptions allow us to characterize the excess control cost of a policy synthesized via certainty equivalence applied to a parameter estimate $\hat{\phi}$, (5).
+In particular, we extend a result from Wagenmaker et al.
+(2023) from the linear in parameters setting to the smooth nonlinear setting.
 Lemma 3.1 (Thm. 1 of Wagenmaker et al. (2023)).
- [^3] Thm. 1 of Wagenmaker et al. (2023) holds for systems with linear dependence on the unknown parameters; however it generalizes with minimal modification to the proof. See Appendix B.
 Suppose Assumptions 2.1-2.4 hold.
-Let $r_{\theta}(\phi^{\star})$ be as defined in Lemma 2.1.
+Let $r_{\theta}(\phi^{\star})$ be as defined in Assumption 2.4.
 Then for $\hat{\phi}\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
@@
 $C_{\mathsf{cost}}=\mathsf{poly}(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}).$
-Footnote 3 informs us that the leading term of the excess cost is given by the parameter estimation error weighted by the model-task Hessian, $\left\|\hat{\phi}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}$.
+Lemma 3.1 informs us that the leading term of the excess cost is given by the parameter estimation error weighted by the model-task Hessian, $\left\|\hat{\phi}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}$.
 Asymptotically, the distribution of the parameter estimation error is normally distributed with mean zero, and covariance given by the inverse Fisher information matrix under the data collection policy $\pi$ evaluated at the true parameter value (cf.
 Theorem 1 of Ljung and Caines (1980)):
-$\lim_{N\to\infty}\sqrt{N}\mathbf{FI}^{\pi}(\phi^{\star})^{1/2}(\hat{\phi}-\phi_{\star})\sim\mathcal{N}(0,I),$
-where the Fisher information for the parameter $\phi$ under the policy $\pi$ for the model (1) is given as
-$\mathbf{FI}^{\pi}(\phi)\!=\!\frac{\operatorname{\mathbb{E}}_{\pi}^{\phi}\left[\sum_{t=1}^{T}\!Df(X_{t},U_{t},\phi)^{\top}\!Df(X_{t},U_{t},\phi)\right]}{\sigma_{w}^{2}},$ (9)
-and measures the signal-to-noise ratio of the data collected from an episode of interaction with the system under the exploration policy $\pi$.
+$\lim_{N\to\infty}\sqrt{N}\mathbf{FI}^{\pi}(\phi^{\star})^{1/2}(\hat{\phi}-\phi_{\star})\sim\mathcal{N}(0,I).$
 We provide a novel non-asymptotic result which characterizes the $H$-norm of the parameter error for a positive definite matrix $H$ in terms of the Fisher information matrix.
@@
 Suppose Assumption 2.1 holds.
-Let $H$ be a positive definite matrix and let $\delta\in(0,\frac{1}{4}]$.
-Consider the least squares estimate $\hat{\phi}$ determined from (3) using data collected from $N$ episodes via an exploration policy $\pi$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz for some $\alpha\in(\frac{1}{4},\frac{1}{2}]$, and satisfies $\lambda_{\min}\left(\Sigma^{\pi}\right)>0$, with $\Sigma^{\pi}$ as defined in (7).
-There exists a polynomial $\texttt{poly}_{\alpha}$ depending only on $\alpha$ such that the following condition holds.
+Consider the least squares estimate $\hat{\phi}$ determined from (3) using data collected from $N$ episodes via an exploration policy $\pi$ which is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz for some $\alpha\in(\frac{1}{4},\frac{1}{2}]$, and satisfies $\lambda_{\min}\left(\mathbf{FI}^{\pi}(\phi^{\star})\right)>0$, with $\mathbf{FI}^{\pi}(\phi^{\star})$ as defined in (7).
+Let $H$ be a positive definite matrix, $\beta$ a positive number satisyfing $\beta\leq\sigma_{w}^{2}\frac{\lambda_{\min}(\mathbf{FI}^{\pi}(\phi^{\star}))}{4},$ and $\delta\in(0,\frac{1}{4}]$.
+Then there exists a polynomial $\mathsf{poly}_{\alpha}$ depending on $\alpha$ such that the following condition holds.
 With probability at least $1-\delta$,
-$\left\|\hat{\phi}-\phi^{\star}\right\|_{H}^{2}$ $\leq 40\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H\mathbf{FI}^{\pi}(\phi^{\star})^{-1})}{N}$ (10)
-as long as
-$N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},d_{\phi},d_{\mathsf{X}},\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
-By substituting the inequality (10) into the leading term of (8), one would expect that the the excess cost of deploying the certainty equivalent policy synthesized on a least squares estimate determined from data collected using policy $\pi$ is characterized by
+$\left\|\hat{\phi}-\phi^{\star}\right\|_{H}^{2}\leq 2(1+\xi)$ (9)
+×$\times\left(\frac{\operatorname{\mathrm{tr}}(H\mathbf{FI}^{\pi}(\phi^{\star})^{-1})}{N}+2\frac{\left\|H\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|}{N}\log\frac{4}{\delta}\right),$
+where $\xi=4\beta\left(\frac{1}{\sigma_{w}^{2}\lambda_{\min}(\mathbf{FI}^{\pi}(\phi^{\star}))}+d_{\phi}\right)$ as long as
+$N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},\!d_{\phi},\!d_{\mathsf{X}},\!\sigma_{w},\!\log N,\!\log\frac{1}{\delta},\!\log\frac{B}{\sigma_{w}},\!C_{\mathsf{Loja}},\!\frac{1}{\beta}\bigg{)}.$
+By substituting the inequality (LABEL:eq:_id_error) into the leading term of (8) and bounding $\left\|\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|\leq\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})$, one would expect that the the excess cost of deploying the certainty equivalent policy synthesized on a least squares estimate determined from data collected using policy $\pi$ is characterized by
 $\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}).$ (10)
@@
 It then proceeds to show that we can find an exploration policy approximately solving (11), even though the parameter $\phi^{\star}$ defining the exploration objective is unknown prior to experimentation.
+It is also necessary to address the fact that the model-task Hessian may not be positive definite.
+Thus optimizing the above objective could result in exploration policies which are not persistently exciting.
 To circumvent the issue of the unknown parameter $\phi^{\star}$, we consider a two step approach in which we first obtain a crude parameter estimate, and then refine it by playing a targeted exploration policy.
@@
 This parameter can be used to search for a policy that approximately solves the optimization problem in (11).
-The straightforward approach to do so is to solve the problem under the estimated parameter:
+A straightforward approach to do so is to solve the problem under the estimated parameter:
 $\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}(H(\hat{\phi}^{-})\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}).$ (12)
-While such a strategy may be effective if the coarse estimate is sufficiently close to the ground truth, it contends with the challenge that there is a distribution shift between the setting that the exploration policy is optimized for, and that which it is deployed in.
-In particular, $\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-1})$ pertains to the statistics from rolling out the dynamics under a parameter value different from the ground truth.
-We therefore instead consider the following problem:
-$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\right)\mbox{ where }$ (14)
-$\Lambda_{\tilde{\pi}}\!=\!\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\!\left[\sum_{t=1}^{T}Df(X_{t},U_{t};\hat{\phi}^{-})^{\top}Df(X_{t},U_{t};\hat{\phi}^{-})\right]$ (15)
-Note the difference between (13) and (14).
-In (13), the Fisher information is determined using an expectation under the the dynamics parameter estimate $\hat{\phi}^{-}$ (see (9)).
-In contrast, the expectation in (14) is evaluated under the true parameter $\phi^{\star}$.
-If (14) can be optimized over the exploration policies, it alleviates the aforementioned distribution shift issue.
-A minimizer for the above objective (14) may be approximated via an online learning procedure, as was recently proposed by Wagenmaker et al.
-(2023) for models that are linear in the parameter.
-In particular, we leverage the dynamic optimal experiment design (DOED) procedure proposed in the aforementioned work.
-For the reader’s convenience, we state the algorithm and highlight several modifications in Section C.1.
-We refer to our modified version of the algorithm as $\texttt{DOED}_{+}$.
+To address the issue of a model-task Hessian which is not positive definite, we introduce regularization into the exploration design. In particular, we set the exploration policy $\pi$ as
+$\pi=\operatorname*{argmin}_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}((H(\hat{\phi}^{-})+\nu I)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}),$ (13)
+for an appropriately chosen regularization parameter $\nu$.
 The above discussion motivates Algorithm 1, named Active Learning for Control-Oriented Identification (ALCOI).
-The algorithm takes as input an initial policy satisfying the Lojasiewicz condition (Assumption 2.5), the exploration policy class, the target policy, the number of exploration rounds, and the desired confidence parameter $\delta$.
+The algorithm takes as input an initial policy satisfying the Lojasiewicz condition (Assumption 2.5), the exploration policy class, the target policy, the number of exploration rounds, a parameter $\gamma\in(0,1)$ which controls the ratio of the exploration budget that the initial loja policy is played, the level of regularization $\nu$, and the precision of the optimization for the exploration policy $\varepsilon$.
 Given these components, the algorithm proceeds in three stages.
-The first stage begins in 2 by playing the initial policy for one fourth of the exploration budget.
+The first stage begins in 2 by playing the initial policy for a portion of the exploration budget controlled by $\gamma$.
 In 3, it uses the collected data to derive a coarse estimate $\hat{\phi}^{-}$ for the unknown parameters by solving a least squares problem.
-Then, in 4 it constructs a confidence ball $\mathcal{B}^{-}$ centered at $\hat{\phi}^{-}$:
-$\mathcal{B}^{-}$ =$=\bigg{\{}\phi:\left\|\phi-\hat{\phi}^{-}\right\|\leq\frac{R}{N^{\alpha}}\bigg{\}},$ (16)
-where
-$R=\left(\frac{2048\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(L_{f}TN\right)+\log\frac{1}{\delta}\right)\right)^{\alpha}.$ (17)
-Next, the estimate $\hat{\phi}^{-}$ is used to construct the model-task Hessian at $\hat{\phi}^{-}$ as $\mathcal{H}(\hat{\phi}^{-})$.
-This is in turn used to define an exploration objective, $\Phi:\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}$ to be passed to the $\texttt{DOED}_{+}$ algorithm.
-This exploration objective is a regularized version of that in (14) which takes as input a Gram matrix $\Lambda$, and outputs $\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$.
-Here $\lambda>0$ is a regularizer chosen from the estimate of the model-task Hessian in 5.
-The quantity $B_{\Phi}$ is then defined from the model-task Hessian and the regularizer to characterize the smoothness of the objective $\Phi$.
-The exploration objective is passed to $\texttt{DOED}_{+}$ along with the coarse estimate $\hat{\phi}^{-}$, the ball $\mathcal{B}^{-}$ and the smoothness bound $B_{\Phi}$ in 8.
-This call to $\texttt{DOED}_{+}$ returns a collection of $\tilde{K}$ policies, where $\tilde{K}$ is a polynomial function of the supplied budget; see Section C.1 for further details.
-The mixture of the returned policies approximately minimizes the exploration objective $\Phi$ over Gram matrices of the form (15), thereby approximating a solution to (14).
-We therefore use them to design a mixture policy which plays any of the returned policies with equal probability, and plays the initial exploration policy $\pi^{0}$ with probability $\frac{1}{2}$.
-[^4] The initial policy $\pi^{0}$ included in $\pi_{\mathsf{mix}}$ to ensure that $\pi_{\mathsf{mix}}$ is a $(2C_{\mathsf{Loja}},\alpha)$-Lojasiewicz policy.
+Next, the estimate $\hat{\phi}^{-}$ is used to construct the model-task Hessian as $\hat{\phi}^{-}$ as $\mathcal{H}(\hat{\phi}^{-})$ and define an exploration objective $\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\lambda I\right)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\right)$.
+Optimizing this objective to precision $\varepsilon$ over the class of exploration policies provides the policy $\pi_{\exp}$.
 This policy is run to collect data from the system, and obtain a fresh estimate $\hat{\phi}^{+}$ for $\phi^{\star}$.
-[^5] In practice, one would use all of the the data to estimate $\hat{\phi}^{+}$.
-However, using only the data from an independent rollout contributes only a constant factor to the rate, and simplifies the analysis.
 Finally, the estimate is used to synthesize the certainty equivalent policy as in (5).
-Algorithm 1 $\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$
-1:Input: Initial policy $\pi^{0}$, exploration policy class $\Pi_{\exp}$, target policy class $\Pi^{\star}$, episodes $N$, confidence $\delta$.
-2: Play $\pi^{0}$ for $\lfloor\frac{N}{4}\rfloor$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},x_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{4}\rfloor}$.
+Algorithm 1 $\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\gamma,\nu,\varepsilon)$
+1:Input: Initial policy $\pi^{0}$, exploration policy class $\Pi_{\exp}$, target policy class $\Pi^{\star}$, initial policy ratio $\gamma$, regularization parameter $\nu$, optimization precision $\varepsilon$.
+2: Play $\pi^{0}$ for $\lfloor\gamma N\rfloor$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\gamma N\rfloor}$.
 3: Fit $\hat{\phi}^{-}$ from the collected data by solving (3).
-4: Set $\mathcal{B}^{-}$ according to (16).
-5: Set $\lambda\leftarrow\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}$.
-6: Define objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\hat{\phi}^{-})(\Lambda+\lambda I)^{-1})$.
-7: Set objective smoothness bound $B_{\Phi}\leftarrow\frac{TL_{f}^{2}\left\|\mathcal{H}(\hat{\phi}^{-})\right\|}{\lambda^{2}}$.
-8: Determine exploration policies as $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}\leftarrow\texttt{DOED}_{+}(\Phi,\lfloor\frac{N}{4}\rfloor,\pi^{0},\hat{\phi}^{-},\mathcal{B}^{-},B_{\Phi})$.
-9: Define $\pi_{\mathsf{mix}}$ which at the start of each episode plays $\pi^{0}$ with probability $\frac{1}{2}$, and for $i=1,\dots,\tilde{K}$ plays $\pi^{i}$ with probability $\frac{1}{2\tilde{K}}$.
-10:Play $\pi_{\mathsf{mix}}$ for $\lfloor N/2\rfloor$ episodes, collecting data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
-11: Fit $\hat{\phi}^{+}$ by solving (3) with the data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor\frac{N}{2}\rfloor}$.
-12:Return: certainty equivalent policy $\hat{\pi}=\pi^{\star}(\hat{\phi}^{+})$
+4: Determine exploration policy as
+$\pi_{\exp}\in\Bigg{\{}\pi\in\Pi_{\exp}|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\right)$
+$\leq(1+\varepsilon)\inf_{\tilde{\pi}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\right)\Bigg{\}}.$
+5: Define $\pi_{\mathsf{mix}}$ which at the start of each episode plays $\pi^{0}$ with probability $\gamma$, and $\pi_{\exp}$ with probability $1-\gamma$.
+6: Play $\pi_{\mathsf{mix}}$ for $\lfloor(1-\gamma)N\rfloor$ episodes, collecting data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor(1-\gamma)N\rfloor}$.
+7: Fit $\hat{\phi}^{+}$ by solving (3) with the data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n},X_{t+1}^{n}\right\}\mathclose{}_{t,k=1}^{T,\lfloor(1-\gamma)N\rfloor}$.
+8:Return: certainty equivalent policy $\hat{\pi}=\pi^{\star}(\hat{\phi}^{+})$.
 Our main result is a finite sample bound characterizing the excess cost of the policy return by Algorithm 1.
 Theorem 3.2 (Main Result).
-Suppose $f$, $\pi^{0}$, $\Pi_{\exp}$, $\Pi^{\star}$ satisfy Assumptions 2.1-2.7.
-Let $\delta\in(0,\frac{1}{4}]$.
-Consider running Algorithm 1 to generate a control policy $\hat{\pi}$ as $\hat{\pi}=\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\delta)$.
-There exists a universal positive constant $c$ and polynomial function $\texttt{poly}_{\alpha}$ depending only on the Lojasiewicz parameter $\alpha$ such that the following holds true.
-With probability at least $1-\delta$ we have that
-$\mathcal{J}(\hat{\pi},\phi^{\star})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\leq c\frac{\log\frac{d_{\phi}}{\delta}}{N}\!\min_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\right),$
-as long as
-$N\!$ $\geq\!\mathsf{poly}_{\alpha}\bigg{(}\!T,L_{f},L_{\mathsf{cost}},L_{\pi^{\star}},d_{\phi},d_{\mathsf{X}},\frac{1}{\lambda_{\min}^{\star}},r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})$
-$\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))},\left\|\mathcal{H}(\phi^{\star})\right\|,\sigma_{w},\log N,\log\frac{1}{\delta},\log B_{f},C_{\mathsf{Loja}}\bigg{)}.$
+Suppose $f$, $\pi^{0}$, $\Pi_{\exp}$, $\Pi^{\star}$ satisfy Assumptions 2.1-2.6.
+Let $\nu$ be a non-negative regularization parameter such that $\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu>0$.
+Let the optimization tolerance $\varepsilon\in(0,1/2)$ and the initial policy ratio $\gamma\in(0,1/2)$.
+Consider running Algorithm 1 to generate a control policy $\hat{\pi}$ as $\hat{\pi}=\texttt{ALCOI}(\pi^{0},\Pi_{\exp},\Pi^{\star},N,\gamma,\nu,\varepsilon)$.
+Let $\delta\in\big{(}0,\frac{1}{4}]$ be the failure probability, and $\beta\in\bigg{(}0,\frac{\mu\left(\lambda_{\min}(H(\phi^{\star}))+\nu\right)}{512d_{\phi}(\left\|H(\phi^{\star})\right\|+\nu)}\bigg{)}$ [^3] Recall that $\mu$ is the persistance of excitation parameter defined in Assumption 2.6.
+be a free parameter in the bound.
+There exists a polynomial function $\texttt{poly}_{\alpha}$ depending on the Lojasiewicz parameter $\alpha$ such that the following holds true.
+With probability at least $1-\delta$, it holds that
+$\mathcal{J}(\hat{\pi},\phi^{\star})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\!\leq\!(1\!+\!4\gamma)(1\!+\!\varepsilon)(1\!+\!\xi)\left(2+4\log\frac{4}{\delta}\right)$
+×$\times\frac{\inf_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\right)}{N},$
+where $\xi=\beta\left(3+16\left(\frac{128d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}+d_{\phi}\right)\right)$ as long as
+$N\!$ $\geq\!\mathsf{poly}_{\alpha}\bigg{(}\!T,L_{f},L_{\mathsf{cost}},L_{\pi^{\star}},d_{\phi},d_{\mathsf{X}},\frac{1}{\mu},r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})$
+$\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\sigma_{w},\sigma_{w}^{-1},\log N,$
+$\log\frac{1}{\delta},\log B,C_{\mathsf{Loja}},\frac{1}{\gamma},\frac{1}{\beta}\bigg{)}.$
 The above result characterizes the excess control cost in terms of three key quantities.
 First, the term $\mathcal{H}(\phi^{\star})$ is the model-task Hessian, which describes how error in identification of the dynamics model parameters translates to the control cost.
-Second, is the inverse Fisher Information of the optimal exploration policy term, which measures a signal-to-noise ratio quantifying the hardness of parameter identification.
+Second is the inverse Fisher information of the optimal exploration policy term, which measures a signal-to-noise ratio quantifying the hardness of parameter identification.
 Finally, the number of exploration episodes $N$ on the denominator captures the rate of decay from increasing the experimental budget.
-We note that in the setting where the dynamics model has linear dependence on the parameters, Wagenmaker et al.
+In the setting where the dynamics model has linear dependence on the parameters, Wagenmaker et al.
 (2023) present a lower bound on the excess control cost achieved by any learner following the model based interaction protocol described in Section 2.
-The upper bound of Theorem 3.2 matches this lower bound up to universal constants, and the log term $\log\left(d_{\phi}/\delta\right)$.
+If we choose the free parameter in the upper bound as $\beta\leq\left(\frac{128d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}+d_{\phi}\right)^{-1}$, then the upper bound of Theorem 3.2 matches this lower bound up to universal constants, and the term $\log\left(4/\delta\right)$.
 Future work will pursue general lower bounds that hold for dynamics models with a nonlinear dependence on the unknown parameter.
-We conjecture that in general, the upper bounds of Theorem 3.2 are tight up to universal constants and the term $\log\left(d_{\phi}/\delta\right)$.
-We note that the burn-in time is currently polynomial in the relevant system parameters; however, we do not pursue tight burn-in times in this work.
+The burn-in time is currently polynomial in the relevant system parameters; however, we do not pursue tight burn-in times in this work.
 It may be possible to improve the dependence of the burn-in on various system quantities, e.g.
 by leveraging stability or reachability to obtain optimal dependence of the burn-in on $T$.
+We additionally draw attention to the utility of the parameter $\beta$ and the algorithm hyperparameters $\nu$ and $\gamma$ for navigating the tradeoff between a good burn-in time, and optimal rates.
+One can take $\beta$, $\nu$, $\gamma$ and $\varepsilon$ arbitrarily close to zero, meaning that the coefficient characterizing the excess cost can become arbitrarily close to $2+4\log\frac{8}{\delta}$.
+The cost of doing so is an increase in the burn-in time.
+A notable exception is the situation where $\mathcal{H}(\phi^{\star})\succ 0$, i.e.
+the setting where all parameters are necessary for control.
+In this case, one can take $\nu=0$, while the burn-in time must exceed a polynomial in $\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}$.
 4 Proof Sketch
@@
 Our main result proceeds by demonstrating the following sub-steps.
-In these sub-steps, let $C$ be a polynomial of the problem parameters, as in the burn-in requirement of Theorem 3.2.
+In these sub-steps, let $C$ be a polynomial of the problem parameters and log of the reciprocal of the failure probability, $\delta$, as in the burn-in requirement of Theorem 3.2.
   1. 1.
-    The coarse parameter estimation error decays gracefully with the total amount of data:
+    With high probability, the coarse parameter estimation error decays gracefully with the total amount of data:
     $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{C}{(TN)^{\alpha}},$ (14)
@@
 In particular, for any exploration policy $\pi\in\Pi_{\exp}$,
-    $\bigg{|}\!\operatorname{\mathrm{tr}}\left(\!\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi}^{-1}\!\right)\!-\!\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})\bigg{|}\!\leq\!C\!\left\|\hat{\phi}^{-}\!\!-\!\phi^{\star}\right\|,$ (19)
-    with $\Lambda_{\pi}$ as in (15).
+    $\bigg{|}\!\operatorname{\mathrm{tr}}\left(\!\mathcal{H}(\hat{\phi}^{-})\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\!\right)\!-\!\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1})\bigg{|}\!$ (15)
+    $\leq\!C\!\left\|\hat{\phi}^{-}\!\!-\!\phi^{\star}\right\|.$
   3. 3.
-    The call to the $\texttt{DOED}_{+}$ algorithm in 8 of Algorithm 1 returns a collection of policies whose mixture approximately minimizes the exploration objective (14):
-    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\Lambda_{\pi_{\mathsf{mix}}}^{-1}\right)\leq\frac{C}{N^{\alpha}}\!+\!\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}\mathcal{H}(\hat{\phi}^{-})\Lambda_{\tilde{\pi}}^{-1}\Bigg{)},$ (20)
-    for $\Lambda_{\pi}$ as in (15). Demonstrating this result involves specializing the dynamic optimal exploration algorithm and its proof of correctness to our setting from Wagenmaker et al. (2023).
-  4. 4.
     For $N$ sufficiently large, we may use the consistency guarantee (14) to prove Theorem 3.1.
@@
 rates we obtain.
-Using the above results, our argument proceeds according to the following series of inequalities applied to the excess cost:
+Using the above results, our argument proceeds according to the following series of inequalities applied to the excess cost. With high probability,
 $\mathcal{J}_{\phi^{\star}}(\hat{\phi}^{+})\!-\!\mathcal{J}_{\phi^{\star}}(\phi^{\star})\leq\left\|\hat{\phi}^{+}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}\!+\!C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{3}$ (16)
-$\leq 40\log\left(\frac{24d_{\phi}}{\delta}\right)\frac{\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})}{N}+\frac{C}{N^{3/2}},$
-where the first inequality follows by Footnote 3, and the second inequality follows by applying Theorem 3.1 with $H=\mathcal{H}(\phi^{\star})$ for the first term, and $H=I$ for the second term [^7] The condition that $\Sigma^{\pi_{\mathsf{mix}}}\succ 0$ in Theorem 3.1 can be shown from the success of DOED in (20)).
-.
+$\leq\xi(\delta)\frac{\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})}{N}+\frac{C}{N^{3/2}},$
+where the first inequality follows by Lemma 3.1, and the second inequality follows by applying Theorem 3.1 with $H=\mathcal{H}(\phi^{\star})+\nu I$ for the first term, and $H=I$ for the second term.
+The quantity $\xi(\delta)$ is a trades off the burn-in time and the final bound.
+In our analysis, it can become as small as $2+4\log\frac{1}{\delta}$.
 Next, it follows that
-$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})$ $\overset{(i)}{\leq}\frac{C}{N^{\alpha}}+\operatorname{\mathrm{tr}}\left(H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\right)$
-$\overset{(ii)}{\leq}2\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\hat{\phi}^{-})\operatorname{\mathbb{E}}_{\tilde{\pi}}^{\phi^{\star}}\left[(Df(x_{t},u_{t};\hat{\phi}^{-}))^{\top}Df(x_{t},u_{t};\hat{\phi}^{-})\right]^{-1}\Bigg{)}$
-$\overset{(iii)}{\leq}3\frac{C}{N^{\alpha}}+\min_{\tilde{\pi}\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\Bigg{(}H(\phi^{\star})\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\Bigg{)},$
-where inequality $(i)$ follows from (19) and (18), inequality $(ii)$ follows from (20), and inequality $(iii)$ follows from (19) and (18).
+$\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})$
+$\overset{(i)}{\leq}\frac{C}{N^{\alpha}}+\operatorname{\mathrm{tr}}\left(\left(H(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-}){-1}\right)$
+$\overset{(ii)}{\leq}\!2\frac{C}{N^{\alpha}}\!+\!\frac{1}{1-\gamma}\inf_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\Bigg{(}\!\left(H(\hat{\phi}^{-})\!+\!\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1}\!\Bigg{)}$
+$\overset{(iii)}{\leq}\!3\frac{C}{N^{\alpha}}\!+\!\frac{1}{1\!-\!\gamma}\!\inf_{\tilde{\pi}\in\Pi_{\exp}}\!\operatorname{\mathrm{tr}}\Bigg{(}\left(H(\phi^{\star})\!+\!\nu I\right)\mathbf{FI}^{\tilde{\pi}}(\phi^{\star})^{-1}\Bigg{)},$
+where inequality $(i)$ follows from (15) and (14), inequality $(ii)$ follows from the definition of the policy $\pi_{\mathsf{mix}}$ and inequality $(iii)$ follows from (15) and (14).
 The main result then follows by substituting the above bound into (16), and taking $N$ to exceed a polynomial burn-in time so the higher order terms become negligible.
@@
 We deploy ALCOI on an illustrative example to illustrate the benefits of active control-oriented exploration.
-For more experiments, and further details, see Appendix E.
+For more experiments, and further details, see Appendix D.
 Consider the two dimensional system
@@
 ACKNOWLEDGMENT
-Ingvar Ziemann is supported by a Swedish Research Council international postdoc grant.
-George J.
-Pappas is supported in part by NSF Award SLES-2331880.
-Nikolai Matni is supported in part by NSF Award SLES-2331880 and NSF CAREER award ECCS-2045834.
+BL and NM are supported by NSF Award SLES-2331880, NSF CAREER award ECCS-2045834 and AFOSR Award FA9550-24-1-0102.
+IZ is supported by a Swedish Research Council international postdoc grant.
+GP is supported in part by NSF Award SLES-2331880.
 Appendix A System Identification Results
@@
 Our first estimation bound instantiates these results in our setting under the assumption that the data collection policy is $(C_{\mathsf{Loja}},\alpha)$-Lojasiewicz.
-To express these bounds we define $\mathcal{G}=\mathopen{}\left\{f(\cdot,\cdot,\phi):\phi\in\mathbb{R}^{d_{\phi}}\right\}\mathclose{}$ and $\mathcal{G}_{\star}=\mathcal{G}-\mathopen{}\left\{f(\cdot,\cdot,\phi^{\star})\right\}\mathclose{}$.
+To express these bounds we define $\mathcal{G}=\mathopen{}\left\{f(\cdot,\cdot,\phi):\phi\in\mathbb{R}^{d_{\phi}},\left\|\phi\right\|\leq B\right\}\mathclose{}$ and $\mathcal{G}_{\star}=\mathcal{G}-\mathopen{}\left\{f(\cdot,\cdot,\phi^{\star})\right\}\mathclose{}$.
 Using the Lipschitz bound on $f$, we find that the size $\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)$ of the the minimal cardinality $\varepsilon$-net covering $\mathcal{G}_{\star}$ is bounded as
-$\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)\leq d_{\phi}\log\left(\frac{2L_{f}}{\varepsilon}+1\right).$ (22)
+$\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon)\leq d_{\phi}\log\left(\frac{3BL_{f}}{\varepsilon}+1\right).$ (17)
 We will present an estimation error bound that holds as long as the number of episodes exceeds some value (a burn-in time) depending on the confidence with which we want it to hold.
 To state this requirement, define
-$\tau_{\mathbf{Err}}(\delta)=\max\bigg{\{}\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{4\alpha-1}},$ (23)
-$\Bigg{(}4(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}\left(\log\left(4BK^{2\alpha}\right)+d_{\phi}\log\left(2L_{f}K^{2\alpha}+1\right)\right)+\log\frac{1}{\delta}\Bigg{)}^{\frac{1}{1-8\alpha+16\alpha^{2}}}\bigg{\}}.$
+$\tau_{\mathbf{Err}}(\delta)=\max\bigg{\{}\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{4\alpha-1}},$ (18)
+$\Bigg{(}8(L_{f}C_{\mathsf{Loja}}\alpha^{-1})^{4}\left(d_{\phi}\log\left(4BL_{f}K^{2\alpha}+1\right)\right)+\log\frac{1}{\delta}\Bigg{)}^{\frac{1}{1-8\alpha+16\alpha^{2}}}\bigg{\}}.$
 Lemma A.1.
@@
 Then with probability at least $1-\delta$
-$\mathbf{Err}_{\pi,\phi^{\star}}(\hat{\phi})\leq\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(\frac{3L_{f}TK}{\delta}\right)\right).$
+$\mathbf{Err}_{\pi,\phi^{\star}}(\hat{\phi})\leq\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(\frac{4L_{f}TK}{\sigma_{w}\delta}\right)\right).$
 To prove this result, we first state several preliminary results.
 In order to do so, we introduce additional notation.
-Let $\mathcal{G}_{\star}^{s}=\mathopen{}\left\{bg:b\in[0,1],g\in\mathcal{G}_{\star}\right\}\mathclose{}$ be the star-shaped hull of $\mathcal{G}_{\star}$.
-[^8] A function class $\mathcal{G}$ is star-shaped if for all $b\in[0,1]$, $g\in\mathcal{G}\implies bg\in\mathcal{G}$.
-Since all elements of $\mathcal{G}_{\star}^{s}$ satisfy $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$ for all $\phi\in\mathbb{R}^{d_{\phi}}$, we have from Lemma 4.5 of Mendelson (2002) that the size $\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)$ of a minimum cardinality $\varepsilon$-net of $\mathcal{G}_{\star}^{s}$ in the $\left\|\cdot\right\|_{\infty}$-norm satisfies
+Let $\mathcal{G}_{\star}^{s}=\mathopen{}\left\{bg:b\in[0,1],g\in\mathcal{G}_{\star}\right\}\mathclose{}$.
+This set is called the star-hull of $\mathcal{G}_{\star}$ [^5] The star-hull of $\mathcal{G}$ is the smallest set containing $\mathcal{G}$ that is star-shaped..
+Since all elements of $\mathcal{G}_{\star}^{s}$ satisfy $\left\|f(\cdot,\cdot,\phi)\right\|_{\infty}\leq B$, we have from Lemma 4.5 of Mendelson (2002) that the size $\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)$ of a minimum cardinality $\varepsilon$-net of $\mathcal{G}_{\star}^{s}$ in the $\left\|\cdot\right\|_{\infty}$-norm satisfies
 $\log\mathcal{N}(\mathcal{G}_{\star}^{s},\left\|\cdot\right\|_{\infty},\varepsilon)\leq\log\left(4B/\varepsilon\right)+\log\mathcal{N}(\mathcal{G}_{\star},\left\|\cdot\right\|_{\infty},\varepsilon).$ (19)
@@
 The above bound for a single member of the function class may be extended to a uniform bound over function in $\mathcal{G}_{\star}^{s}$, but outside the ball $B(r)$ using a covering argument.
-Lemma A.4 (Lower isometry).
+Lemma A.4 (Lower isometry, Multi-trajectory Modification of Theorem 5.2 of Ziemann and Tu (2022)).
 Suppose there exists a $r/\sqrt{8}$-net $\mathcal{G}_{r}$ of $\partial B(r)$ in the norm $\left\|\cdot\right\|_{\infty}$ such that $(\mathcal{G}_{r},\pi)$ satisfies $(C_{\mathsf{hc}},a)$-hypercontractivity.
 Then the following lower isometry holds
-$\operatorname{\mathbb{P}}\left[\sup_{g\in\mathcal{G}_{\star}^{s}\setminus B(r)}\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
+$\operatorname{\mathbb{P}}\left[\forall g\in\mathcal{G}_{\star}^{s}\setminus B(r),\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}-\frac{1}{8}\operatorname{\mathbb{E}}\left[\frac{1}{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 0\right]\leq{\left|\mathcal{G}_{r}\right|}\exp\left(\frac{-Kr^{4-2a}}{4C_{\mathsf{hc}}}\right).$
 Proof.
@@
 $\operatorname{\mathbb{E}}\left[\frac{1}{T}\sum_{t=1}^{T}\left\|g(X_{t}^{1},U_{t}^{1})\right\|^{2}\right]\leq 8\frac{1}{TK}\sum_{t,k=1}^{T,K}\left\|g(X_{t}^{k},U_{t}^{k})\right\|^{2}+r^{2}.$
-We may bound $\log{\left|\mathcal{G}_{r}\right|}$ using (24) and (22) as $\log{\left|\mathcal{G}_{r}\right|}\leq\log\left(\frac{4B}{r}\right)+d_{\mathsf{X}}\log\left(\frac{2L_{f}}{r}+1\right).$ Then choosing $r^{2}=\frac{1}{K^{a}}$, we find that by the requirement that
-$K\geq\left(4C_{\mathsf{hc}}\left(\log\left(4BK^{a/2}\right)+d_{\phi}\log\left(2L_{f}K^{a/2}+1\right)\right)+\log\frac{1}{\delta}\right)^{\frac{1}{1-2a+a^{2}}},$
+We may bound $\log{\left|\mathcal{G}_{r}\right|}$ using (19) and (17) as $\log{\left|\mathcal{G}_{r}\right|}\leq\log\left(\frac{4B}{r}\right)+d_{\phi}\log\left(\frac{3BL_{f}}{r}+1\right).$ Then choosing $r^{2}=\frac{1}{K^{a}}$, we find that by the requirement that
+$K\geq\left(8C_{\mathsf{hc}}\left(d_{\phi}\log\left(4L_{f}K^{a/2}\right)\right)+\log\frac{1}{\delta}\right)^{\frac{1}{1-2a+a^{2}}},$
 the inequality
@@
 We may then in turn bound the offset complexity over a finite hypothesis class using Proposition F.2 of Ziemann et al. (2023). With probability at least $1-\delta$,
-$\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$ $\leq 16\sigma_{w}^{2}\log\left(\frac{{\left|\mathcal{G}_{\gamma}\right|}}{\delta}\right)\leq 16\sigma_{w}^{2}\left(d_{\phi}\log\left(\frac{2L_{f}}{\gamma}+1\right)+\log\frac{1}{\delta}\right),$ (29)
-where the final inequality followed from (22). Setting $\gamma=\frac{\sigma_{w}}{TK}$, we may combine (28) with (29) to conclude
-$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)\leq\frac{12\sigma_{w}^{2}d_{\mathsf{X}}}{TK}+\frac{\sigma_{w}^{2}}{(TK)^{2}}+16\sigma_{w}^{2}\left(d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\leq\frac{32\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right).$
+$\sup_{g\in\mathcal{G}_{\gamma}}\frac{1}{TK}\left(\sum_{t,k=1}^{T,K}\left(4\langle g(X_{t}^{n},U_{t}^{n}),W_{t}^{k}\rangle-\frac{1}{2}\left\|g(X_{t}^{n},U_{t}^{n})\right\|^{2}\right)\right)$ (24)
+$\leq 16\sigma_{w}^{2}\log\left(\frac{{\left|\mathcal{G}_{\gamma}\right|}}{\delta}\right)\leq 16\sigma_{w}^{2}\left(d_{\phi}\log\left(\frac{3BL_{f}}{\gamma}+1\right)+\log\frac{1}{\delta}\right),$
+where the final inequality followed from (17). Setting $\gamma=\frac{\sigma_{w}}{TK}$, we may combine (23) with (LABEL:eq:_finite_offset_complexity_bound) to conclude
+$\sup_{g\in\mathcal{G}_{\star}}M_{T,K}(g)$ $\leq\frac{12\sigma_{w}^{2}d_{\mathsf{X}}}{TK}+\frac{\sigma_{w}^{2}}{(TK)^{2}}+16\sigma_{w}^{2}\left(d_{\phi}\log\left(3BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right)$
+$\leq\frac{32\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3BL_{f}TK/\sigma_{w}+1\right)+\log\frac{1}{\delta}\right).$
 By the burn-in requirement
-$K\geq\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{a-1}},$
-we may combing the above inequality with (26) and (27) to achieve the desired result. ∎
-With this result in hand, we may immediately determine a high probability bound on the parameter estimation error as it appears in the bound in Footnote 3.
+$K\geq\left(\frac{256\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4L_{f}TK/\sigma_{w}\right)+\log\frac{1}{\delta}\right)\right)^{\frac{1}{a-1}},$
+we may combine the above inequality with (21) and (22) to achieve the desired result. ∎
+With this result in hand, we may immediately determine a high probability bound on the parameter estimation error as it appears in the bound in Lemma 3.1.
 Corollary A.1.
 Under the setting of Lemma A.1, it follows that with probability at least $1-\delta$,
-$\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}\!\leq\!\left(\frac{512\sigma_{w}^{2}}{TN}\left(d_{\mathsf{X}}\!+\!d_{\phi}\log\left(\frac{3L_{f}TN}{\delta}\right)\right)\right)^{2\alpha}.$
+$\left\|\hat{\phi}\!-\!\phi^{\star}\right\|^{2}\!\leq\!\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}\!+\!d_{\phi}\log\left(4BL_{f}TK\right)\right)+\log\frac{1}{\delta}\right)^{2\alpha}.$
 For $\alpha<\frac{1}{2}$, the above bound decays at a sub-optimal rate. However, the dependence can be improved by leveraging the Delta method, outlined in the subsequent section.
@@
 $\hat{\Sigma}_{K}^{\pi}=\frac{1}{K}\sum_{k=1}^{K}\sum_{t=1}^{T}(Df_{t}^{k})^{\top}Df_{t}^{k}.$ (25)
-The superscript $\pi$ denotes the policy that is rolled out to generate $U_{t}^{k}$.
-Given sufficiently many trajectories, this estimate converges to the expected value $\Sigma^{\pi}$ (defined in (7)).
+The superscript $\pi$ denotes the policy that is rolled out to generate $U_{t}^{k}$. Given sufficiently many trajectories, this estimate converges to the expected value
+$\Sigma^{\pi}\triangleq\mathbb{E}_{\pi}^{\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right].$ (26)
 As long as this covariance matrix is positive definite, $\Sigma^{\pi}\succ 0$, we can show that the delta method is successful in improving the estimation rate.
 In order to do so, we require sufficiently many trajectories.
-In addition to the burn-in for the excess risk bound, we require that $N$ is large enough that Corollary A.1 guarantees $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ with high probability.
+In addition to the burn-in for the excess risk bound, we require that $K$ is large enough that Corollary A.1 guarantees $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ with high probability.
 We refer to the burn-in required for this to hold as $K\geq\tau_{\mathsf{small\,error}}(\delta)$, where
-$\tau_{\mathsf{small\,error}}(\delta)=\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right).$
+$\tau_{\mathsf{small\,error}}(\delta)=\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right).$
 Furthermore, we require that the higher order terms ($\mathsf{h.o.t.}$) from the taylor expansion involved in the Delta method are dominated by the lower order terms.
-For downstream use of the delta method, we allow this to depend on the norm of an arbitrary positive definite matrix $H$.
-We refer to the required burn-in as $K\geq\tau_{\mathsf{h.o.t.}}(\delta,H)$, where
-$\tau_{\mathsf{h.o.t}}(\delta)$ $=\Bigg{(}\frac{8T^{1-\alpha}L_{f}^{2}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
-$+\frac{8d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}{\lambda_{\min}(\Sigma^{\pi})\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}\sqrt{32T\log\frac{4}{\delta}}\Bigg{)}^{\frac{1}{\alpha}}.$
+For downstream use of the delta method, we allow this to depend on the norm of an arbitrary positive definite matrix $H$ and an arbitrary positive number $\beta$.
+We refer to the required burn-in as $K\geq\tau_{\mathsf{h.o.t.}}(\delta,\beta)$, where
+$\tau_{\mathsf{h.o.t}}(\delta,H)$ $=\Bigg{(}\frac{8T^{1-\alpha}L_{f}^{2}}{\beta}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4BL_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}+\frac{8d_{\mathsf{X}}\sigma_{w}L_{f}}{\beta}\sqrt{32T\log\frac{4}{\delta}}\Bigg{)}^{\frac{1}{\alpha}}.$
 With these burn-in times defined, we may state the result which expresses the parameter estimation error
@@
 Let Assumption 2.1 hold and fix $\delta\in(0,\frac{1}{2}]$.
-Let $H$ be a positive semi-definite matrix.
+Let $H$ be a positive definite matrix and $\beta$ a positive number.
 Fix some $\delta\in(0,\frac{1}{4}]$.
 Consider setting $\hat{\phi}$ as the least squares solution from $K$ episodes of data collected with policy $\pi$ for
-$K$ $\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/2),\tau_{\mathsf{small\,error}}(\delta),\tau_{\mathsf{h.o.t}}(\delta,H)\right\}\mathclose{}$
+$K$ $\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/2),\tau_{\mathsf{small\,error}}(\delta),\tau_{\mathsf{h.o.t}}(\delta,\beta)\right\}\mathclose{}$
 Suppose $\hat{\Sigma}_{K}^{\pi}\succeq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})>0$. Then with probability at least $1-\delta$,
 $\hat{\phi}-\phi^{\star}$ $=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k},$
-where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}$.
+where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$.
 As long as $\Sigma^{\pi}\succ 0$ and $K$ is sufficiently large that $\Sigma_{K}^{\pi}$ concentrates to $\Sigma^{\pi}$, we may use the above result to show that $\hat{\phi}_{K}-\phi^{\star}$ behaves like a self-normalized martingale.
-We may then invoke Theorem 7.2 of Wagenmaker et al.
-(2021) to bound this self-normalized martingale, leading to Theorem 3.1.
+We may then bound this self-normalized martingale, leading to Theorem 3.1.
 Before proving the above lemma, we introduce several preliminary results.
@@
 Let $G=-A^{-1}U(\Sigma^{-1}+V^{\top}A^{-1}U)^{-1}V$.
-By submultiiplicativity, $\left\|G\right\|_{\mathsf{op}}\leq\frac{1}{\lambda_{\min}(A)}\left(\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)\right)^{-1}.$ We have that
+By submult iplicativity, $\left\|G\right\|_{\mathsf{op}}\leq\frac{1}{\lambda_{\min}(A)}\left(\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)\right)^{-1}.$ We have that
 $\sigma_{\min}(\Sigma^{-1}+V^{\top}A^{-1}U)$ $\geq\sigma_{\min}(\Sigma^{-1})-\left\|V^{\top}A^{-1}U\right\|=\sigma_{\min}(\Sigma^{-1})-\frac{1}{\lambda_{\min}(A)}.$
@@
 Proof.
-By the fact that $\hat{\phi}_{K}$ solves (3),
+By the fact that $K\geq\tau_{\mathbf{Err}}(\delta/2)$, the conditions on Corollary A.1 are satisfied such that with probability at least $21-\delta/$ 2,
+$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}.$ (27)
+Using the assumption $K\geq\tau_{\mathsf{small\,error}}(\delta)$, we have that this is at most $1$.
+Under this event, and by the assumption that $\left\|\phi^{\star}\right\|\leq B-1$, we have that $\hat{\phi}_{K}$ is is in the interior of $\mathopen{}\left\{\phi:\left\|\phi\right\|\leq B\right\}\mathclose{}$.
+Thus by the fact that $\hat{\phi}_{K}$ solves (3), it holds that
 00 0 $=\left(\nabla_{\phi}\sum_{k=1}^{K}\sum_{t=1}^{T}\left\|X_{t+1}^{k}-f(X_{t}^{k},U_{t}^{k};\phi)\right\|^{2}\right)|_{\phi=\hat{\phi}_{K}}\triangleq\psi(\hat{\phi}_{K}).$
-A second order taylor expansion of $\psi(\hat{\phi}_{K})$ about $\phi^{\star}$ provides
-$\psi(\hat{\phi}_{K})$ $=\psi(\phi^{\star})+D_{\phi}\psi(\phi)|_{\phi=\phi_{\star}}(\hat{\phi}_{K}-\phi^{\star})+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}],$
-for some $\tilde{\phi}$ which is a convex combination of $\phi^{\star}$ and $\hat{\phi}_{K}$.
-Here $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot,\cdot]$ is a tensor mapping $\mathbb{R}^{d_{\phi}\times d_{\phi}}\to\mathbb{R}^{d_{\phi}}$.
-We write $D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot]$ as the tensor mapping $\mathbb{R}^{d_{\phi}}\to\mathbb{R}^{d_{\phi}\times d_{\phi}}$ such that
-$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star},\hat{\phi}_{K}-\phi^{\star}]=D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}](\hat{\phi}_{K}-\phi^{\star}).$
-As long as $D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ is non-singular, then the above two equations imply
-$\hat{\phi}_{K}$ $-\phi^{\star}=\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star}).$
+A first order Taylor expansion of $\psi(\hat{\phi}_{K})$ about $\phi^{\star}$ provides
+$\psi(\hat{\phi}_{K})$ $=\psi(\phi^{\star})+D_{\phi}\psi(\phi)|_{\phi=\phi_{\star}}(\hat{\phi}_{K}-\phi^{\star})+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}],$
+where $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})$ is a trilinear operator such that the vector $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}]$ is defined by the product of the matrix $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot\hat{\phi}_{K}-\phi_{\star}]$ with $\hat{\phi}_{K}-\phi_{\star}$: $\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\hat{\phi}_{K}-\phi_{\star},\hat{\phi}_{K}-\phi_{\star}]=\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot\hat{\phi}_{K}-\phi_{\star}](\hat{\phi}_{K}-\phi_{\star})$, and the satisfies the operator norm bound below:
+$\left\|\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right\|_{\mathsf{op}}$ $\leq\sup_{\ell\in[0,1]}\left\|D_{\phi}^{2}\psi(\phi)|_{\phi=\ell\phi^{\star}+(1-\ell)\hat{\phi}_{K}}[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]\right\|_{\mathsf{op}}.$
+As long as $D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]$ is non-singular, then the above two equations imply
+$\hat{\phi}_{K}$ $-\phi^{\star}=\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right)^{-1}\psi(\phi^{\star}).$
 Evaluating the derivatives provides
 $\psi(\phi^{\star})$ $=\sum_{k=1}^{K}\sum_{t=1}^{T}(Df_{t}^{k})^{\top}W_{t}^{k}$
-$D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}$ $=K\hat{\Sigma}_{K}^{\pi}+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{2}f_{t}^{k})[W_{t}^{k}]}_{E_{1}}$
-$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]$ $=\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}]Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})}_{E_{2}}$
-$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})^{\top}(D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\hat{\phi}_{K}-\phi^{\star}])^{\top}}_{E_{3}}$
-$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{3}f(X_{t}^{k},U_{t}^{k};\tilde{\phi}))[\hat{\phi}_{K}-\phi^{\star},W_{t}^{k}]}_{E_{4}}$
-We will bound the operator norms of $E_{1},E_{2},E_{3}$ and $E_{4}$ in order to apply Lemma A.6. We may bound $E_{2}$ and $E_{3}$ directly using Assumption 2.1:
+$D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}$ $=K\hat{\Sigma}_{K}^{\pi}+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{2}f_{t}^{k})[\cdot,\cdot,W_{t}^{k}]}_{E_{1}},$
+$D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]$ $=\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]^{\top}Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})}_{E_{2}}$
+$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(Df(X_{t}^{k},U_{t}^{k};\tilde{\phi})^{\top}D^{2}f(X_{t}^{k},U_{t}^{k};\tilde{\phi})[\cdot,\cdot,\hat{\phi}_{K}-\phi^{\star}]}_{E_{3}}$
+$+\underbrace{\sum_{k=1}^{K}\sum_{t=1}^{T}(D^{3}f(X_{t}^{k},U_{t}^{k};\tilde{\phi}))[\cdot,\cdot,W_{t}^{k},\hat{\phi}_{K}-\phi^{\star}]}_{E_{4}}$
+We will bound the norms of $E_{1},E_{2},E_{3}$ and $E_{4}$ for $\tilde{\phi}$ an arbitrary convex combination of $\hat{\phi}_{K}$ and $\phi_{\star}$ in order to apply Lemma A.6.
+We may bound $E_{2}$ and $E_{3}$ directly using Assumption 2.1:
 $\left\|E_{2}\right\|_{\mathsf{op}}$ $\leq TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|\mbox{ and }\left\|E_{3}\right\|_{\mathsf{op}}\leq TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|.$
@@
 $\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq d_{\mathsf{X}}\sigma_{w}L_{f}(1+\left\|\hat{\phi}_{K}-\phi^{\star}\right\|)\sqrt{32TK\log\frac{4}{\delta}}+2TKL_{f}^{2}\left\|\hat{\phi}_{K}-\phi_{\star}\right\|.$
-By the fact that $K\geq\tau_{\mathbf{Err}}(\delta/2)$, the conditions on Corollary A.1 are satisifed such that with probability at least $21-\delta/$ 2,
-$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq\left(\frac{512\sigma_{w}^{2}}{TK}\left(d_{\mathsf{X}}+d_{\phi}\log\left(3L_{f}TK\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}.$
-Using the burn-in $K\geq\tau_{\mathsf{small\,error}}(\delta)$, we have that this is at most $1$. Therefore, with probability at least $1-\delta$,
-$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq 2d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}+2(TK)^{1-\alpha}L_{f}^{2}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(2L_{f}TK+1\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
-As $K\geq\tau_{\mathsf{h.o.t}}(\delta,H)$, we have that
-$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left((H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma^{\pi})^{-1})}K\lambda_{\min}\left(\hat{\Sigma}_{K}^{\pi}\right).$
-We may therefore invoke Lemma A.6 with $s=\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{2\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}$ to show that
-$\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\frac{1}{2}D_{\phi}^{2}\psi(\phi)|_{\phi=\tilde{\phi}}[\hat{\phi}_{K}-\phi^{\star}]\right)^{-1}\psi(\phi^{\star})=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k},$
-for some $G$ satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}\left(H(\Sigma^{\pi})^{-1}\right)}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}.$
+Using the fact that $\left\|\hat{\phi}_{K}-\phi^{\star}\right\|\leq 1$ along with the bound in (27) it holds that with probability at least $1-\delta$,
+$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq 2d_{\mathsf{X}}\sigma_{w}L_{f}\sqrt{32TK\log\frac{4}{\delta}}+2(TK)^{1-\alpha}L_{f}^{2}\left(512\sigma_{w}^{2}\left(d_{\mathsf{X}}+d_{\phi}\log\left(4L_{f}TK/\sigma_{w}\right)+\log\frac{2}{\delta}\right)\right)^{\alpha}$
+As $K\geq\tau_{\mathsf{h.o.t}}(\delta,\beta)$ and $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$, it holds that
+$\left\|E_{1}+E_{2}+E_{3}+E_{4}\right\|_{\mathsf{op}}\leq\frac{\beta}{2}K\lambda_{\min}\left(\hat{\Sigma}_{K}^{\pi}\right).$
+We may therefore invoke Lemma A.6 with $s=\frac{\beta}{2}$ to show that
+$\left(D_{\phi}\psi(\phi)|_{\phi=\phi^{\star}}+\mathsf{Remainder}_{\psi}(\hat{\phi}_{K})[\cdot,\cdot,\hat{\phi}_{K}-\phi_{\star}]\right)^{-1}\psi(\phi^{\star})=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k},$
+for some $G$ satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$.
 A.3 Proof of Theorem 3.1
-To proceed with our proof of the main result, we will leverage Lemma A.5.
-In particular, we require a bound on the self-normalized martingale that arises.
-We state one such bound due to Wagenmaker et al.
-(2021) for this purpose.
-The bound is stated in order to characterize the norm of a self-normalized martingale with respect to an arbitrary positive definite matrix $H$.
-Lemma A.8 (Theorem 7.2 of Wagenmaker et al. (2021)).
-Suppose $\Sigma^{\pi}\succ 0$ and $H\succeq 0$.
-Define the event $\mathcal{E}\triangleq\mathopen{}\left\{\left\|\Sigma_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}\right\}\mathclose{})$ for some $\beta\in(0,\frac{1}{4})$.
+To proceed with our proof of the main result, we will leverage Lemma A.5. In particular, we require a bound on the weighted norm of a self-normalized martingale.
+Lemma A.8.
+Suppose $\Sigma^{\pi}\succ 0$ and $H\succ 0$.
+Define the event $\mathcal{E}\triangleq\mathopen{}\left\{\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\right\}\mathclose{}$ for some positive $\beta$ satisfying
+$\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$
 As long as $\mathcal{E}$ holds, the following holds with probability at least $1-\delta$,
-$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 5(1+\alpha)\sigma_{w}^{2}\log\frac{6d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma^{\pi}\right)^{-1}\right)}{K},$
-where $\alpha=26\beta^{2}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right))$.
+$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 2\sigma_{w}^{2}\left(1+\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}\left(1+\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{1}{\delta}.$
 Proof.
-The proof follows from the proof of Theorem 7.2 of Wagenmaker et al. (2021) by making the following substitutions for $\bm{Z}$ and $\bm{w}$ in the proof:
+Define
 $\bm{Z}=\begin{bmatrix}Df_{1}^{1}\\ \vdots\\ Df_{T}^{1}\\ \vdots\\ Df_{1}^{K}\\ \vdots\\ Df_{T}^{K}\end{bmatrix}\quad\mbox{ and }\quad\bm{w}=\begin{bmatrix}W_{1}^{1}\\ \vdots\\ W_{t}^{1}\\ \vdots\\ W_{1}^{K}\\ \vdots\\ W_{t}^{K}\end{bmatrix}.$
-To guarantee concentration of the empirical covariance (30) to the true covariance (7), we state the following covariance concentration result.
+such that
+$\left\|\left(K\hat{\Sigma}_{K}^{\pi}\right)^{-1}\sum_{t,k=1}^{T,K}(Df_{t}^{n})^{\top}W_{t}^{k}\right\|_{H}^{2}=\left\|(\bm{Z}\bm{Z}^{\top})^{-1}\bm{Z}\bm{W}\right\|_{H}^{2}\mbox{ and }\hat{\Sigma}_{K}^{\pi}=\frac{1}{K}\bm{Z}\bm{Z}^{\top}.$
+We may write
+$\left\|\left(\bm{Z}\bm{Z}^{\top}\right)^{-1}\bm{Z}\bm{W}\right\|_{H}^{2}$ $=\left\|H^{1/2}\left(\bm{Z}\bm{Z}^{\top}\right)^{-1}\bm{Z}\bm{W}\right\|^{2}$
+$=\left\|\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+$=\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+$=2\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}+\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+We have by assumption that $\left\|\frac{1}{K}\bm{Z}\bm{Z}^{\top}-\Sigma^{\pi}\right\|\leq\beta$, and therefore that
+$H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\succeq\left(1-\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)H^{-1/2}K\Sigma^{\pi}H^{-1/2}.$
+It follows from the above inequality and the assumption that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$ that
+$2\left\|\left(\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}+\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\left\|\left(\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}+\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+$=2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left\|\left(\frac{\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}}{\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)}+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}.$
+The above quantity is a self-normalized martingale.
+In particular, we can bound it by applying the bound in Theorem 4.1 of Ziemann et al.
+(2023).
+To do so, define $\tilde{\Sigma}\triangleq\frac{\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{2}}{\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)}$.
+Then it follows that with probability at least $1-\delta$ that
+$2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left\|\left(\tilde{\Sigma}+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\right)^{-1/2}H^{-1/2}\bm{Z}\bm{W}\right\|^{2}$
+$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\log\operatorname{\textbf{det}}\left(I+H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
+$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right),$
+where the second inequality follows by noting that for any positive definite matrix matrix $M\in\mathbb{R}^{d\times d}$, $\log\operatorname{\textbf{det}}(I+M)=\sum_{i=1}^{d}\log(1+\lambda_{i}(M))\leq\sum_{i=1}^{d}\lambda_{i}(M)$.
+By the concentration event on $\bm{Z}\bm{Z}^{\top}$ and the definition of $\tilde{\Sigma}$, this quantity may be bounded as
+$2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H^{-1/2}\bm{Z}\bm{Z}^{\top}H^{-1/2}\Sigma^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
+$\leq 2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\lambda_{\min}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\right)^{-1}\left(\sigma_{w}^{2}\left(1+\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}\left(H^{-1/2}K\Sigma^{\pi}H^{-1/2}\tilde{\Sigma}^{-1}\right)+2\sigma_{w}^{2}\log\frac{1}{\delta}\right)$
+$=2\left(1+\beta\frac{2}{\lambda_{\min}(\Sigma^{\pi})}\right)\left(\sigma_{w}^{2}\left(1+\frac{\beta}{\lambda_{\min}(\Sigma^{\pi})}\right)\operatorname{\mathrm{tr}}\left(H\left(K\Sigma^{\pi}\right)^{-1}\right)+2\sigma_{w}^{2}\left\|H\left(K\Sigma^{\pi}\right)^{-1}\right\|\log\frac{1}{\delta}\right)$
+$\leq 2\left(1+\beta\frac{4}{\lambda_{\min}(\Sigma^{\pi})}\right)\left(\sigma_{w}^{2}\operatorname{\mathrm{tr}}\left(H\left(\Sigma^{\pi}\right)^{-1}\right)+2\sigma_{w}^{2}\left\|H\left(\Sigma^{\pi}\right)^{-1}\right\|\log\frac{1}{\delta}\right),$
+where the final inequality follows from the fact that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{2}$. This concludes the proof.
+To guarantee concentration of the empirical covariance (25) to the true covariance, we state the following covariance concentration result.
 Lemma A.9.
-Let $\beta\in(0,\frac{1}{4})$. Suppose
-$K\geq\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta)}{\beta^{2}\lambda_{\min}(\Sigma^{\pi})^{2}}.$
+Let $\beta\in(0,\frac{\lambda_{\min}(\Sigma^{\pi})}{4})$. Suppose
+$K\geq\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta))}{\beta^{2}}.$
 Then with probability at least $1-\delta$,
-$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}).$
+$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta.$
 Proof.
@@
 We will invoke the covariance concentration lemma above with the value of $\beta$ chosen according to Lemma A.8. In particular, we define
-$\tau_{\mathsf{cov\,conc}}(\delta)=\frac{52T^{2}L_{f}^{4}\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)(3d_{\phi}+\log(1/\delta))}{\lambda_{\min}(\Sigma^{\pi})^{2}}.$
-If $K\geq\tau_{\mathsf{cov\,conc}}(\delta)$, we may invoke Lemma A.9 to show that with probability at least $41-\delta/$ 4,
-$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta\lambda_{\min}(\Sigma^{\pi}),$
-where $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}\leq\frac{1}{2}$.
-As a result, we have that $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$.
-Given this condition and the burn-in conditions $K\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/16),\tau_{\mathsf{small\,error}}(\delta/4),\tau_{\mathsf{h.o.t.}}(\delta/4)\right\}\mathclose{}$, we satisfy the conditions of Lemma A.5.
+$\tau_{\mathsf{cov\,conc}}(\delta)=\frac{2T^{2}L_{f}^{4}(3d_{\phi}+\log(1/\delta))}{\beta^{2}}.$
+If $K\geq\tau_{\mathsf{cov\,conc}}(\delta/4)$, we may invoke Lemma A.9 to show that with probability at least $41-\delta/$ 4,
+$\left\|\hat{\Sigma}_{K}^{\pi}-\Sigma^{\pi}\right\|\leq\beta.$
+By the fact that $\beta\leq\frac{\lambda_{\min}(\Sigma^{\pi})}{4}$, we have that $\lambda_{\min}(\hat{\Sigma}_{K}^{\pi})\geq\frac{1}{2}\lambda_{\min}(\Sigma^{\pi})$.
+Given this condition and the burn-in conditions $K\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/8),\tau_{\mathsf{small\,error}}(\delta/4),\tau_{\mathsf{h.o.t.}}(\delta/4,\beta)\right\}\mathclose{}$, we satisfy the conditions of Lemma A.5.
 Therefore, conditioned on the event of covariance concentration, we have that with probability at least $41-\delta/$ 4,
 $\hat{\phi}_{K}-\phi^{\star}$ $=(I+G)(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k},$
-where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\sqrt{\left\|H\right\|_{\mathsf{op}}}\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}$.
-Then by a triangle inequality and submultiplicativity, we have that under the success events of covariance concentration and the delta method,
-$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 2\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}+2\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1})}{\operatorname{\mathrm{tr}}((\Sigma_{K}^{\pi})^{-1})}\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}$ (31)
-In particular, we have reduced the parameter recovery errort to the sum of weighted norms of a self-normalized martingale.
+where $G$ is a matrix satisfying $\left\|G\right\|_{\mathsf{op}}\leq\beta$. Then by the triangle inequality, we have that under the success events of covariance concentration and the delta method,
+$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq\left(\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}+\beta\left\|H\right\|^{1/2}\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|\right)^{2}.$ (28)
+We have reduced the parameter recovery errort to the sum of weighted norms of a self-normalized martingale.
 Calling on self-normalized martingale machinary Abbasi-Yadkori et al.
-(2011), Wagenmaker et al.
-(2021) we may bound each of these norms with high probability.
-In particular, invoking Lemma A.7 with $\beta=\frac{1}{\sqrt{26\left\|\Sigma^{\pi}\right\|\operatorname{\mathrm{tr}}\left((\Sigma^{\pi})^{-1}\right)}}$, we find that with probability at least $21-\delta/$ 2, the following bounds hold:
-$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(H\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
-$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|^{2}\leq 10\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}\left(\left(\Sigma_{K}^{\pi}\right)^{-1}\right)}{K},$
-Substituting these bounds into (31), we find that by union bounding over the success events, we have that with probability at least $1-\delta$,
-$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 40\sigma_{w}^{2}\log\frac{24d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(H(\Sigma_{K}^{\pi})^{-1}}{K}.$
+(2011).
+In particular, invoking Lemma A.8 with failure probability $\delta/4$ and union bounding provides that with probability at least $21-\delta/$ 2, the following bounds hold:
+$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|_{H}\leq\sqrt{2\sigma_{w}^{2}(1+\xi)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}(1+\xi)\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{4}{\delta}},$
+$\left\|(\hat{\Sigma}_{K}^{\pi})^{-1}\frac{1}{K}\sum_{t,k=1}^{T,K}(Df_{t}^{k})^{\top}W_{t}^{k}\right\|\leq\sqrt{2\sigma_{w}^{2}(1+\xi)\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1})+4\sigma_{w}^{2}(1+\xi)\left\|(K\Sigma^{\pi})^{-1}\right\|\log\frac{4}{\delta}},$
+where $\xi=\frac{4\beta}{\lambda_{\min}(\Sigma^{\pi})}$ Substituting these bounds into (28), we find that by union bounding over the success events, we have that with probability at least $1-\delta$,
+$\left\|\hat{\phi}_{K}-\phi^{\star}\right\|_{H}^{2}\leq 2\sigma_{w}^{2}(1+\xi+4\beta d_{\phi})\operatorname{\mathrm{tr}}(\left(K\Sigma^{\pi}\right)^{-1}H)+4\sigma_{w}^{2}(1+\xi+4\beta d_{\phi})\left\|(K\Sigma^{\pi})^{-1}H\right\|\log\frac{4}{\delta}.$
 Appendix B Results for smooth nonlinear systems
@@
 In particular, we first define $p_{w}$ to denote the density of the noise, $W_{t}$.
-Let $p_{\phi,\theta}(\cdot)$ denote the density over the trajectories $(X_{1:T+1},U_{1:T})$ induced by playing controller $\pi^{\theta}$ on the system (1) with parameter $\phi$.
+Let $p_{\phi,\theta}(\cdot)$ denote the density over the states $X_{2:T+1}$ induced by playing controller $\pi^{\theta}$ on the system (1) with parameter $\phi$.
 Observe that
-$p_{\theta,\phi}(X_{1:T+1},U_{1:T})=\prod_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi^{\theta}_{t}(X_{t});\phi^{\star})).$ (32)
+$p_{\theta,\phi}(X_{2:T+1})=\prod_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi^{\theta}_{t}(X_{t});\phi^{\star})).$ (29)
 We may express our average controller cost as
-$J(\pi^{\theta},\phi)=\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
+[…]
+with $X_{1}$ fixed as 00.
 Note that by our noise assumptions and the smoothness assumptions, we may interchange limits and integrals by appealing to the dominated convergence theorem.
@@
 To show differentiability, note that
-$\frac{d}{dx_{1}}J(\pi^{\theta_{\bm{s}}},\phi_{\bm{m}})=\int\frac{d}{dx_{1}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
-$=\int\underbrace{\left(\frac{d}{dx_{1}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)\right)}_{\mbox{Derivative of log density}}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T}).$
-Proceeding to differentiate the above quantity with respect to $x_{2}$, $x_{3}$, and $x_{4}$, provides from the product rule a sum of terms which are of the form above, but with the derivative of the log density replaced by the product of higher order derivatives of the log density.
+[…]
+[…]
+[…]
+Proceeding to differentiate the above quantity with respect to $x_{2}$, $x_{3}$, and $x_{4}$, provides from the product rule a sum of terms which are of the form above, but with the derivative of the log density and the cost gradient replaced by the product of higher order derivatives of the log density.
 Recalling the form of the density in (29), we have that
-$\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)=\sum_{t=1}^{T}-\frac{1}{2\sigma_{w}^{2}}\left\|X_{t+1}-f(X_{t},\pi^{\theta_{\bm{s}}};\phi_{\bm{m}})\right\|^{2}+T\log\frac{1}{\sqrt{2\pi\sigma_{w}^{2d_{\mathsf{X}}}}}.$
+$\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)=\sum_{t=1}^{T}-\frac{1}{2\sigma_{w}^{2}}\left\|X_{t+1}-f(X_{t},\pi^{\theta_{\bm{s}}};\phi_{\bm{m}})\right\|^{2}+T\log\frac{1}{\sqrt{2\pi\sigma_{w}^{2d_{\mathsf{X}}}}}.$
 By the differentiability of $f$ with respect to $\phi$ and $u$ from Assumption 2.1 and the differentiability of $\pi^{\theta}$ with respect to $\theta$ from Assumption 2.2, we have that the log density above is four times differentiable in $x_{1},\dots x_{4}$ for any values of $\Delta_{i}^{\theta}$ and $\Delta_{j}^{\phi}$.
+Similarly, by differentiability of the stage costs from Assumption 2.3 and the policy from Assumption 2.2, the cost gradient is three times differentiable in $x_{1},\dots,x_{4}$.
 To bound the norm of the gradients, observe that
-$\frac{d}{dm_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{\phi}f(X_{t},\pi^{\theta;\phi})\Delta_{i}^{\phi}$
-$\frac{d}{ds_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{u}f(X_{t},\pi^{\theta};\phi)D_{\theta}\pi^{\theta}\Delta_{i}^{\theta}$
+$\frac{d}{dm_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{\phi}f(X_{t},\pi^{\theta;\phi})\Delta_{i}^{\phi}$
+$\frac{d}{ds_{i}}\log\left(p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{2:T+1})\right)|_{\bm{s}=\bm{m}=0}$ $=\sum_{t=1}^{T}\frac{1}{\sigma_{w}^{2}}\left(X_{t+1}-f(X_{t},\pi^{\theta};\phi)\right)^{\top}D_{u}f(X_{t},\pi^{\theta};\phi)D_{\theta}\pi^{\theta}\Delta_{i}^{\theta}$
 Differentiating further with respect to $m_{i}$ and $s_{i}$ results only in higher order derivatives.
 Using the operator norm bounds on the derivative of the policy and the dynamics, we may conclude that for $i,j\in\mathopen{}\left\{0,1,2,3\right\}\mathclose{}$ such that $1\leq i+j\leq 4$,
-$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)$
+$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{2:T+1})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)$
 We may then in turn bound $\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}$ as
 $\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)$
-×$\times\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{4}\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})$
-$\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1}\right)\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
-×$\times\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})},$
+×$\times\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{4}\left(\sum_{t=1}^{T}c_{t}(X_{t},\pi_{\theta}(X_{t})+c_{T+1}(X_{T+1})\right)p_{\theta,\phi}(X_{1:T+1})dX_{2:T+1}$
+[…]
+[…]
 by Cauchy-Schwarz. We have from Assumption 2.3 that
-$\sqrt{\int\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+c_{T+1}(X_{T+1})\right)^{2}p_{\theta_{\bm{s}},\phi_{\bm{m}}}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}$
-$=\sqrt{\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}c_{t}(X_{t},U_{t})+C_{T+1}(X_{T})\right)^{2}\right]}\leq L_{\mathsf{cost}}.$
 […]
-$\sqrt{\int\left(\sum_{t=1}^{T}\left(1+\frac{\left\|X_{t+1}-f(X_{t},\pi^{\theta}(X_{t});\phi)\right\|}{\sigma_{w}}\right)\right)^{8}p_{\theta,\phi}(X_{1:T+1},U_{1:T})d(X_{1:T+1},U_{1:T})}\leq\mathsf{poly}(T,d_{\mathsf{X}}).$
-Therefore our derivative is bounded as
-$\left\|D_{\phi}^{i}D_{\theta}^{j}\log\left(p_{\theta,\phi}(X_{1:T+1},U_{1:T})\right)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{f},L_{\theta},\sigma_{w}^{-1},L_{\mathsf{cost}},d_{\mathsf{X}}\right).$
-Given Lemma B.1, the proof of Lemma 2.1 follows from an argument identical to that of Proposition 6 in Wagenmaker et al.
+[…]
+[…]
+We may similarly bound the cost gradient by a polynomial in $L_{\mathsf{cost}}$ and $L_{\theta}$. Therefore our derivative is bounded as
+$\left\|D_{\phi}^{(i)}D_{\theta}^{(j)}\mathcal{J}(\pi^{\theta},\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(T,L_{f},L_{\theta},\sigma_{w}^{-1},L_{\mathsf{cost}},d_{\mathsf{X}}\right).$
+Given Lemma B.1, Assumption 2.4 may be shown to hold if $\nabla_{\theta}^{2}J(\pi^{\theta},\phi^{\star})\succ 0$ using an argument identical to that of Proposition 6 in Wagenmaker et al.
 (2023), with Lemma B.1 invoked in place of their Lemma D.1.
@@
 Lemma B.2 (Lemma D.3 of Wagenmaker et al. (2023)).
-Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
+Under Assumptions 2.1, 2.2, 2.3 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
 $\left\|\mathcal{H}(\phi^{\star})-\mathcal{H}(\phi)\right\|_{\mathsf{op}}\leq C_{\mathsf{Hpert}}\left\|\phi-\phi^{\star}\right\|.$
@@
 Lemma B.3 (Lemma D.4 of Wagenmaker et al. (2023)).
-Under Assumptions 2.1, 2.3, 2.2 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
+Under Assumptions 2.1, 2.2,2.3 and 2.4 we have that for any $\phi\in\mathcal{B}(\phi^{\star},\min\mathopen{}\left\{r_{\mathsf{cost}}(\phi^{\star}),r_{\theta}(\phi^{\star})\right\}\mathclose{})$,
 $\left\|\mathcal{H}(\phi)\right\|_{\mathsf{op}}\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right).$
-The proofs of Footnote 3, Lemma B.2, and Lemma B.3 from the same arguments as the proofs of Lemma D.2, Lemma D.3 and Lemma D.4 of Wagenmaker et al.
-(2023) with Lemma B.1 taking place of their Lemma D.1, and the conclusion of Lemma 2.1 taking place of their Assumption 13.
-Appendix C Modifications to Dynamic Optimal Experiment Design
-C.1 Dynamic Optimal Experiment Design
-We consider an experiment design procedure, Algorithm 2, which is a modified version of the algorithm from Wagenmaker et al.
-(2023).
-The goal of this algorithm is to collect a dataset consisting of $N$ episodes playing exploration policies belonging to $\Pi_{\exp}$ whose Gram matrix $\Lambda$ (15) which approximately minimizes the objective $\Phi(\Lambda)$.
-It takes as input the objective $\Phi$, an episode budget $N$, an initial exploration policy $\pi^{0}$, a parameter estimate $\hat{\phi}$, and a confidence set for the parameters $\mathcal{B}$.
-The algorithm plays a sequence of $N^{1/3}$ epochs, each with $N^{2/3}$ episodes.
-For the first epoch, it plays the initial policy to set an initial value of the Gram matrix $\Lambda_{0}$.
-Starting from this value, the algorithm performs a variant of gradient descent on the objective $\Phi$, where each epoch is a gradient step.
-At each iteration, it computes the gradient $\Xi_{n}$ of this objective at the current iterate $\Lambda_{n}$.
-Rather than taking a step in the direction $\Xi_{n}$, it takes a step in the descent direction constrained to the set of Gram matrices achievable under the dynamics evolve according to (1) when the input is chosen according to some exploration policy $\pi\in\Pi_{\exp}$.
-It does so by estimating the solution to
-$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right],$
-where $c^{n}$ is defined in terms of the gradient $\Xi_{n}$ as
-$c^{n}(x,u)=\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{n}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}.$
-In the above optimization problem, the expectation is taken with respect to the true parameter, which is unknown.
-However, the confidence ball $\mathcal{B}$ centered at the estimate $\hat{\phi}$ can be used to approximate the solution.
-In particular, the algorithm applies optimism in the face of uncertainty (Abbasi-Yadkori et al., 2011; Kakade et al., 2020) to select the exploration policy which solves the problem
-$\underset{\tilde{\pi}\in\Pi_{\exp}}{\min}\min_{\phi\in\mathcal{B}}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{n}(X_{t},U_{t})\right].$
-Doing so allows us to achieve high probability bounds on the objective value obtained by the dynamic optimal experiment design algorithm (20).
-The formal version of this is modified from Theorem 3 of Wagenmaker et al.
-(2023), and is presented in Theorem C.1.
-[…]
-1:Input: Objective $\Phi$, episodes $N$, initial exploration policy $\pi^{0}$, parameter estimate $\hat{\phi}$, parameter uncertainty set $\mathcal{B}$, smoothness bound $B_{\Phi}$
-2:Set $\tilde{N}\leftarrow\lfloor N^{2/3}\rfloor$, $\tilde{K}\leftarrow\lfloor N^{1/3}\rfloor-1$, $\gamma_{n}\leftarrow\frac{1}{n+1}$.
-3:Play $\pi^{0}$ for $\tilde{N}$ episodes to collect $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$,
-4:Store $\mathcal{D}_{0}\leftarrow\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,k=1}^{T+1,\tilde{N}}$
-5:Form the Gram matrix $\Delta_{0}=\frac{1}{\tilde{N}}\sum_{{t,n=1}}^{T,\tilde{N}}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})$
-6:for $k=1,\dots,\tilde{K}$ do
-7:     Define $\Xi_{k}\leftarrow\nabla_{\Lambda}\Phi(\Lambda)|_{\Lambda=\Lambda_{k-1}}$ and $c^{k}(x,u)\!\leftarrow\!\frac{\operatorname{\mathrm{tr}}\left(Df(x,u;\hat{\phi})\Xi_{k}Df(x,u;\hat{\phi})^{\top}\right)}{B_{\Phi}}$
-8:     Set $\pi^{k}\leftarrow\underset{\tilde{\pi}\in\Pi_{\exp}}{\operatorname*{argmin}}\,\underset{\phi\in\mathcal{B}}{\min}\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c^{k}(X_{t},U_{t})\right]$
-9:     Run $\pi^{k}$ for $\tilde{N}$ episodes, collecting $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T,\tilde{N}}$, and update the dataset and Gram matrix as
-$\Lambda_{k}$ $\leftarrow(1-\gamma_{k})\Lambda_{k-1}+\frac{\gamma_{k}}{\tilde{N}}\sum_{t,n=1}^{T,\tilde{N}}Df(x,u;\hat{\phi})^{\top}\left(\Xi_{k}\right)Df(x,u;\hat{\phi}),\quad\mathcal{D}_{k}\leftarrow\mathcal{D}_{k-1}\cup\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,n=1}^{T+1,\tilde{N}}$
-10:Return: $\Lambda_{\tilde{K}}$, $\mathopen{}\left\{\pi^{k}\right\}\mathclose{}_{k=1}^{\tilde{K}}$
-C.2 Proof of correctness for $\texttt{DOED}_{+}$ Algorithm
-The objective of Algorithm 2 is to collect data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{n,t=1}^{N,T}$ from the system to optimize the some objective $\Phi(\Lambda)$, where
-$\Lambda=\sum_{n,t=1}^{N,T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}).$
-In light of this, we define the set of all possible Gram matrices corresponding to the parameter parameter $\hat{\phi}$:
-$\Omega(\phi)$ =$=\bigg{\{}\sum_{t=1}^{T}Df(x_{t},u_{t};\phi)^{\top}Df(x_{t},u_{t};\phi):x_{t}\in\mathbb{R}^{d_{\mathsf{X}}},u_{t}\in\mathbb{R}^{d_{\mathsf{U}}},\forall h\in[H]\bigg{\}}.$
-The correctness of the algorithm can be guaranteed under the following regularity assumption on the objective.
-Assumption C.1.
-[…]
-$\left\|\nabla\Phi(\Lambda)-\nabla\Phi(\Lambda^{\prime})\right\|_{\star}\leq\kappa\left\|\Lambda-\Lambda^{\prime}\right\|\,\,\,\,\forall\Lambda,\Lambda^{\prime}\in\Omega(\hat{\phi}).$
-Furthermore, suppose there exists some $B_{\Phi}<\infty$ such that for all $\Lambda\in\Omega(\hat{\phi}),x\in\mathbb{R}^{d_{\mathsf{X}}},u\in\mathbb{R}^{d_{\mathsf{U}}}$,
-$\left\|Df(x,u,\hat{\phi})\nabla\Phi(\Lambda)Df(x,u,\hat{\phi})^{\top}\right\|\leq B_{\Phi}.$
-This leads us to the following result, characterizing the optimality of the data collected via the dynamic optimal experiment design procedure.
-Theorem C.1 (Modification of Theorem 3 of Wagenmaker et al. (2023)).
-Suppose Assumptions 2.1 and C.1 hold.
-Additionally, suppose that $\phi^{\star}\in\mathcal{B}$.
-Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$.
-Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfies
-$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
-$\leq\frac{2\kappa T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}\!+\!B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}\!+\!\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right).$
-To prove the above result, we first present an extension of the simulation lemma from Kakade et al. (2020) to the setting of smooth nonlinear systems.
-Lemma C.1 (Self-Bounding, Simulation Lemma: Modified from Kakade et al. (2020)).
-Suppose Assumption 2.1 holds. For any policy $\pi$, model parameters $\phi$ and non-negative cost $c$ satisfying $c<1$, it holds that
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi-\phi^{\star}\right\|.$
+The proofs of Lemma 3.1, Lemma B.2, and Lemma B.3 from the same arguments as the proofs of Lemma D.2, Lemma D.3 and Lemma D.4 of Wagenmaker et al.
+(2023) with Lemma B.1 taking place of their Lemma D.1.
+Finally, we present a result that demonstrates that under our smoothness assumptions, the Fisher information is differentiable with respect to the dynamics parameter, and that the derivative is bounded by a quantity which is polynomial in relevant system parameters and the length of the horizon.
+Lemma B.4.
+For any exploration policy $\pi\in\Pi_{\exp}$, any $\phi\in\mathbb{R}^{d_{\phi}}$, and any unit vector $v\in\mathbb{R}^{d_{\phi}}$ we have that $v^{\top}\mathbf{FI}^{\pi}(\phi)v$ is differentiable with respect to $\phi$, and that
+$\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v\right\|\leq 3\frac{T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}.$
 Proof.
-The proof follows as in the proof of Lemma B.3 of Kakade et al.
-(2020).
-Consider a state/input pair $(X_{t},U_{t})$ from rolling out the policy $\pi$ on the system $\phi^{\star}$.
-Where Kakade et al.
-(2020) apply the bound on the difference in means between two Gaussian distributions, we use the Gaussian distributions describing the next state update from $X_{t},U_{t}$ under the model described by $\phi$ and by $\phi^{\star}$: $\mathcal{N}(f(X_{t},U_{t};\phi^{\star}),\sigma_{w}^{2})$ and $\mathcal{N}(f(X_{t},U_{t};\phi),\sigma_{w}^{2})$.
-This results in a bound on the quantity of interest as
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]$
-$\leq\sqrt{T\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]}\sqrt{\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2},1\right\}\mathclose{}\right]}.$
-Using the fact that the cost is bounded by one, we have
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]\leq T^{2}$
-For the remaining quantity, we may bound
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi\right\|,1\right\}\mathclose{}\right]$ $\leq\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2}\right]$
-$=\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|Df(X_{t},U_{t},\tilde{\phi})(\phi^{\star}-\phi)\right\|^{2}\right]$
-$\leq\frac{TL_{f}^{2}}{\sigma_{w}^{2}}\left\|\phi^{\star}-\phi\right\|^{2}.$
-The equality follows by a first order Taylor expansion of $f$ about $\phi^{\star}$.
-Here $\tilde{\phi}$ is a convex combination of $\phi$ and $\phi^{\star}$.
-The final inequality follows from Assumption 2.1.
-Combining these results concludes the proof.
-∎
-We now proceed to prove Theorem C.1.
-Proof.
-The proof follows by modifying the proof of Theorem 3 in Wagenmaker et al. (2023). In place of the regret minimization algorithm used by Wagenmaker et al. (2023) to estimate the solution to
-$\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right],$
-we use the policy synthesized directly via optimism over the parameter estimate set $\mathcal{B}$.
-Let $\phi_{m}$ be the parameter value under which the optimistic policy was synthesized and let $\pi^{m}$ be the optimistic policy.
-By our assumption that $\phi^{\star}\in\mathcal{B}$, the optimistic policy synthesis provides that
-$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right].$
-By Lemma C.1, we have that
-$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi^{\star}-\phi^{m}\right\|.$
-By the fact that $D(\mathcal{B})$ is the diameter of $\mathcal{B}$, it follows that the above quantity is bounded as
-$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}D(\mathcal{B}).$
-Substituting this bound in place of the bound form the regret minimization step in the proof of Theorem 3 of Wagenmaker et al.
-(2023), and carrying out the remaining steps from the proof yields that with probability at least $1-\delta$,
-$\Phi(\Delta_{\tilde{K}})-\operatorname*{argmin}_{\pi\in\Pi_{\exp}}\Phi(\Sigma^{\pi})\leq\frac{2\beta T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}+B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log(2\tilde{K}/\delta)}{\tilde{N}}}\right).$
-We now consider the correctness of $\texttt{DOED}_{+}$ as it is invoked in 8 by verifying that the conditions to apply the above result are met.
-Lemma C.2 (Modified from Corollary 1 of Wagenmaker et al. (2023)).
-Let Assumption 2.1 hold.
-Suppose that $\phi^{\star}\in\mathcal{B}$.
-Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$.
-Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$ for some $\lambda>0$.
-Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$.
-Then with probability at least $1-\delta$, the data collected at the final round of DOED satisfies
-$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
-$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{N}+1)}{\lambda^{3}(\tilde{N}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{N}}{\delta}}{\tilde{K}}}\right).$
-Proof.
-The result follows from Theorem C.1.
-The smoothness parameters of Assumpton C.1 under the chosen objective are shown in Corollary 1 of Wagenmaker et al.
-(2023) to be bounded as $B_{\Phi}\leq\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}$ and $\kappa\leq\frac{2T^{2}L_{f}^{4}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{3}}$.
-∎
-Lemma C.3.
-Suppose Assumptions 2.1 and 2.6 hold.
-Additionally suppose that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{4TL_{f}^{2}}$.
-Then there exists an exploration policy $\pi\in\Pi_{\exp}$ such that
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]\geq\frac{\lambda_{\min}^{\star}}{2}.$
-Proof.
-The result follows by observing that there exists a policy such that
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}.$
-by Assumption 2.6. Then
-$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]-\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]$
-$=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right].$
-By a first order Taylor expansion of $Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})$ about $\phi^{\star}$, and using the boundedness conditions of Assumption 2.1, we can bound
-$\left\|\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right]\right\|_{\mathsf{op}}\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{2}.$
-Lemma C.4.
-Assume that Assumptions 2.1 and 2.6 hold.
-Additionally, suppose that $\phi^{\star}\in\mathcal{B}$.
-Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}}\left\|\phi_{1}-\phi_{2}\right\|$.
-Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$.
-Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$.
-Let $\delta\in(0,\frac{1}{2}]$.
-Suppose $\lambda\leq\frac{\lambda_{\min}(H)\lambda_{\min}^{\star}}{8\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$, $D(\mathcal{B})\leq\frac{R}{N^{\alpha}}$ and
-$N\geq\max\mathopen{}\left\{\left\lceil\frac{4\lambda_{\min}^{\star}T^{4}L_{f}^{8}\log(N+1)}{\lambda^{3}d_{\phi}}\right\rceil^{3},\left\lceil\frac{8\lambda_{\min}^{\star}TL_{f}^{2}\sqrt{2\log(2N/\delta)}}{\lambda^{2}d_{\phi}}\right\rceil^{3},\left(\frac{4T^{3}L_{f}^{3}R\lambda_{\min}^{\star}}{\lambda^{2}\sigma_{w}d_{\phi}}\right)^{1/\alpha}\right\}\mathclose{}$
-Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfy
-$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]\right)$ (33)
-$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
-and
-$\lambda_{\min}\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)\geq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}.$
-Proof.
-By Lemma C.2, we have that with probability at leat $1-\delta$,
-$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
-$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
-$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(N+1)}{\lambda^{3}(\lfloor N^{1/3}\rfloor+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}R}{N^{\alpha}\sigma_{w}}+\sqrt{\frac{8\log\frac{2N}{\delta}}{\lfloor N^{2/3}\rfloor}}\right)$
-$\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}},$
-where the final inequality follows from the lower bound on $N$. By Assumption 2.6 paired with Lemma C.3, we have that
-$\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
-$=\operatorname{\mathrm{tr}}\left(\mathcal{H}\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]+\lambda I\right)^{-1}\right)\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}}.$
-We may lower bound
-$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)$
-$\geq\frac{\lambda_{\min}\left(\mathcal{H}\right)}{\lambda_{\min}\left(\frac{1}{\tilde{N}\tilde{K}}\sum_{t,n=1}^{T,(\tilde{N}\times\tilde{K})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)+\lambda}.$
-We conclude by combining these inequalities, and using the assumption that $\lambda\leq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$. ∎
+Express
+$\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v=\nabla_{\phi}\int\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{t});\phi)Df(X_{t},\pi(X_{t});\phi)^{\top}vp_{\phi}(X_{1:T+1})d_{X_{2:T+1}},$
+where $X_{1}=0$ and $p_{\phi}(X_{1:T+1})=\Pi_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi(X_{1:t});\phi)).$
+The gradient is then given by
+$\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v$ $=\int\left(\nabla_{\phi}\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{1:t});\phi)Df(X_{t},\pi(X_{1:t});\phi)^{\top}v\right)p_{\phi}(X_{1:T+1})d_{X_{2:T+1}}$
+$+\int\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{1:t});\phi)Df(X_{t},\pi(X_{1:t});\phi)^{\top}v\left(\nabla_{\phi}\log\left(p_{\phi}(X_{1:T+1})\right)\right)p_{\phi}(X_{1:T+1})d_{X_{2:T+1}}.$
+By Assumption 2.1, the first term is differentiable, and the norm is bounded by $2TL_{f}^{2}$. For the second term, observe that
+$\nabla_{\phi}\log\left(p_{\phi}(X_{1:T+1})\right)=\nabla_{\phi}\left(-\frac{1}{2\sigma_{w}^{2}}\sum_{t=1}^{T}\left\|X_{t+1}-f(X_{t},\pi(X_{1:T});\phi)\right\|^{2}\right)=\frac{1}{\sigma_{w}^{2}}\sum_{t=1}^{T}D_{\phi}f(X_{t},\pi(X_{1:t};\phi)W_{t}.$
+Therefore, the norm of the second term may be bounded by $\frac{T^{2}L_{f}^{3}\sqrt{d_{\mathsf{X}}}}{\sigma_{w}}$ by appealing to the triangle inequality, submultiplicativity, the bound $\mathbf{E}[\left\|g_{t}\right\|]\leq\sqrt{n}$ if $g_{t}$ is a standard multi-variate random vector of dimension $n$.
+Appendix C Proof of Main Excess Cost Bound, Theorem 3.2
+First we present a result stating that using the estimated parameter in place of the true parameter leads to an adequate approximation for the ideal exploration objective.
 Lemma C.1.
 Consider a policy $\pi\in\Pi_{\exp}$.
-Suppose $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\min\mathopen{}\left\{\frac{\tilde{\lambda}}{4TL_{f}^{2}},1\right\}\mathclose{}$ for some $\tilde{\lambda}>0$.
-Define
-$\Sigma^{\pi}_{\hat{\phi}}=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right],$
-and suppose that $\Sigma^{\pi}_{\hat{\phi}}\succeq\tilde{\lambda}I$. Then $\Sigma^{\pi}\succeq\frac{\tilde{\lambda}}{2}I$, and
+Suppose $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\tilde{\lambda}\min\mathopen{}\left\{\sigma_{w},1\right\}\mathclose{}}{6T^{2}\max\mathopen{}\left\{1,L_{f}\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}$ for some $\tilde{\lambda}>0$.
+Assume that $\mathbf{FI}^{\pi}(\hat{\phi})\succeq\tilde{\lambda}I$.
+Then $\mathbf{FI}^{\pi}(\phi^{\star})\succeq\frac{\tilde{\lambda}}{2}I$, and
   1. 1.
-    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}\right)^{-1}_{\hat{\phi}}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
+    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
   2. 2.
-    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
+    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
 Proof.
-To begin, we observe that by a first order Taylor expansion of $Df(x,u;\phi^{\star})$ about $\hat{\phi}$, we have that
-$\Sigma^{\pi}$ $\succeq\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]+G,$
-where $\left\|G\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$ by the boundedness assumptions on the derivatives of $f$ from Assumption 2.1.
-By the assumption that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\tilde{\lambda}}{4TL_{f}^{2}}$, it holds that $\left\|G\right\|\leq\frac{\tilde{\lambda}}{2}$.
-Then it follows that $\lambda_{\min}(\Sigma^{\pi})\geq\lambda_{\min}(\Sigma^{\pi}_{\hat{\phi}})-\left\|G\right\|\geq\frac{\tilde{\lambda}}{2}$.
-We proceed to bound $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|$.
-We immediately have that $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\frac{1}{\lambda_{\min}\left(\Sigma^{\pi}\right)\lambda_{\min}\left(\Sigma^{\pi}_{\hat{\phi}}\right)}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq\frac{2}{\tilde{\lambda}^{2}}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|$.
-By Taylor expanding $Df(x,u;\hat{\phi})$ about $\phi^{\star}$, we have from the boundedness assumptions of Assumption 2.1 that $\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|+TL_{f}\left\|\hat{\phi}-\phi^{\star}\right\|^{2}\leq 3TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$.
+Begin by bounding $\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|=\sup_{v\in\mathbb{R}^{d_{\theta}},\left\|v\right\|=1}v^{\top}\mathbf{FI}^{\pi}(\hat{\phi})v-v^{\top}\mathbf{FI}^{\pi}(\phi^{\star})v$.
+By a first order Taylor expansion, it holds that for any unit vector $v\in\mathbb{R}^{d_{\theta}}$,
+$v^{\top}\mathbf{FI}^{\pi}(\hat{\phi})v-v^{\top}\mathbf{FI}^{\pi}(\phi^{\star})v\leq\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\tilde{\phi})v\right\|\left\|\hat{\phi}-\phi^{\star}\right\|.$
+for some $\tilde{\phi}=\alpha\hat{\phi}+(1-\alpha)\phi^{\star}$, $\alpha\in[0,1]$. By Lemma B.4, it holds that
+$\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\tilde{\phi})v\right\|\leq 3\frac{T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}.$
+Then by the assumed bound on $\left\|\hat{\phi}-\phi^{\star}\right\|$, we find that $\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|\leq\frac{\tilde{\lambda}}{2}$.
+Therefore, $\mathbf{FI}^{\pi}(\phi_{\star})\succeq\frac{\tilde{\lambda}}{2}$.
+We also have that
+$\left\|\mathbf{FI}^{\pi}(\hat{\phi})^{-1}-\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|\leq\frac{1}{\lambda_{\min}\left(\mathbf{FI}^{\pi}(\hat{\phi})\right)\lambda_{\min}\left(\mathbf{FI}^{\pi}(\phi^{\star})\right)}\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|\leq\frac{6T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\tilde{\lambda}^{2}\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}\left\|\hat{\phi}-\phi^{\star}\right\|$
 Furthermore, by Lemma B.2, $\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
 To conclude, we bound
-${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\frac{6TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|}{\tilde{\lambda}^{2}},$
+${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\right|}$ $\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\left\|\mathbf{FI}^{\pi}(\phi^{\star})^{-1}-\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right\|$
+$\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\frac{6T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\tilde{\lambda}^{2}\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}\left\|\hat{\phi}-\phi^{\star}\right\|$
 and
-${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\frac{d_{\phi}}{\tilde{\lambda}}\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|$
+${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\right|}$ $\leq\frac{d_{\phi}}{\tilde{\lambda}}\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|$
 $\leq\frac{d_{\phi}}{\tilde{\lambda}}\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
 Combining these inequalities with the triangle inequality provides the desired result. ∎
-Appendix D Proof of Main Excess Cost Bound, Theorem 3.2
 Armed with the above results, we are ready to prove Theorem 3.2.
@@
 Proof.
-We have that $\frac{N}{4}\geq\tau_{\mathbf{Err}}(\delta/4)$.
-Then by Assumption 2.5, we may invoke Corollary A.1 to show that with probability at least $41-\delta/$ 4, $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$, for $R$ defined in (17).
-This in turn implies that ϕsuperscriptitalic-$\phi^{\star}\in\mathcal{B}^{-}$ for $B^{-}$ define in (16).
-The diameter of the ball is given as $D(\mathcal{B}^{-})=\frac{2R}{N^{\alpha}}$.
-Denote the event that ϕsuperscriptitalic-$\phi^{\star}\in\mathcal{B}^{-}$ as $\mathcal{E}_{\mathsf{coarse}}$.
-We condition on the event $\mathcal{E}_{\mathsf{coarse}}$ for the remainder of the proof.
+We have that $\gamma N\geq\tau_{\mathbf{Err}}(\delta/2)$.
+Then by Assumption 2.5, we may invoke Corollary A.1 to show that with probability at least $41-\delta/$ 4, $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{\left(\frac{512\sigma_{w}^{2}}{T}\left(d_{\mathsf{X}}+d_{\phi}\log\left(L_{f}TN\right)+\log\frac{4}{\delta}\right)\right)^{\alpha}}{(\gamma N)^{\alpha}}\triangleq\frac{R}{(\gamma N)^{\alpha}}$.
+Denote the event that this inequality holds as $\mathcal{E}_{\mathsf{coarse}}$.
+Condition on the event that this event holds for the remainder of the proof.
 Let $C_{\mathsf{Hpert}}$ be the system theoretic constant defined in Lemma B.2.
-As long as $\frac{N}{4}\geq\frac{2RC_{\mathsf{Hpert}}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}^{\frac{1}{\alpha}}$, it follows that
+As long as $\gamma N\geq\frac{2RC_{\mathsf{Hpert}}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}^{\frac{1}{\alpha}}$, it follows that
 $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}{2C_{\mathsf{Hpert}}}.$ (30)
 Then by Lemma B.2, we have $\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\geq\frac{1}{2}\lambda_{\min}(\mathcal{H}(\phi^{\star}))$ and $\left\|\mathcal{H}(\hat{\phi}^{-})\right\|\leq 2\left\|\mathcal{H}(\phi^{\star})\right\|$.
-Condition on the event $\mathcal{E}_{\mathsf{coarse}}$. By the choice of $\lambda$ in 5 in Algorithm 1, the assumption that
-$\frac{N}{4}\geq\tau_{\mathsf{PD\,data}}(\delta/4),$
-where
-$\tau_{\mathsf{PD\,data}}(\delta)=\max\mathopen{}\left\{\left\lceil\frac{4\lambda_{\min}^{\star}T^{4}L_{f}^{8}\log(N+1)}{\lambda^{3}d_{\phi}}\right\rceil^{3},\left\lceil\frac{8\lambda_{\min}^{\star}TL_{f}^{2}\sqrt{2\log(2N/\delta)}}{\lambda^{2}d_{\phi}}\right\rceil^{3},\left(\frac{4T^{3}L_{f}^{3}R\lambda_{\min}^{\star}}{\lambda^{2}\sigma_{w}d_{\phi}}\right)^{1/\alpha}\right\}\mathclose{},$
-and Assumption 2.6, we may guarantee that with probability at least $1-\frac{\delta}{4}$ the success event of Lemma C.4 holds.
-Denote this event $\mathcal{E}_{\mathsf{doed}}$.
-$\texttt{DOED}_{+}$ collects $\tilde{K}\tilde{N}$ episodes of data data $\mathopen{}\left\{X_{t}^{n},U_{t}^{n}\right\}\mathclose{}_{t,k=1}^{T+1,(\tilde{N}\times\tilde{K})}$, where $\tilde{K}=\lfloor\frac{N}{4}^{1/3}\rfloor$ and $\tilde{N}=\lfloor\frac{N}{4}^{2/3}\rfloor$.
-Under the success event, this collected data satisfies
-$\lambda_{\min}\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)\geq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{4\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}.$ (35)
-By the law of total expectation, the mixture policy $\pi^{\mathsf{mix}}$ defined in 9 of Algorithm 1 satisfies
-$\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ $=\frac{1}{2}\operatorname{\mathbb{E}}_{\pi^{0},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ (36)
-$+\frac{1}{2\tilde{K}}\sum_{i=1}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$
-$\succeq\frac{1}{2(\tilde{N}+1)}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right].$
-It follows from Lemma C.10 of Wagenmaker et al.
-(2023) that the empirical covariance of the data collected from Algorithm 2 concentrates to the true covariance.
-In particular, with probability at least $41-\delta/$ 4,
-$\left\|\frac{1}{\tilde{K}+1}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]-\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right\|$ (37)
-$\leq d_{\phi}TL_{f}^{2}\sqrt{\frac{8d_{\phi}\log\left(1+8\sqrt{\tilde{K}\tilde{N}}\right)+8\log(4/\delta)}{\tilde{K}\tilde{N}}}.$
-Denote this concentration event as $\mathcal{E}_{\mathsf{doed\,conc}}$. Under the burn-in requirement,
-$\frac{N}{4}\geq\tau_{\mathsf{doed\,conc}}(\delta)=\left\lceil\left(\frac{16\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}\right)^{2}d_{\phi}^{4}T^{2}L_{f}^{4}\left(8d_{\phi}\log\left(1+8\sqrt{N/4}\right)+8\log(4/\delta)\right)\right\rceil^{3},$ (38)
-this event implies
-$\frac{1}{\tilde{K}+1}\sum_{i=0}^{\tilde{K}}\operatorname{\mathbb{E}}_{\pi^{i},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]$ $\succeq\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})$
-$-\frac{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}d_{\phi}^{2}}I.$
-Combining this inequality with (36) and (35) and using the fact that $\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\geq\frac{1}{2}\lambda_{\min}(\mathcal{H}(\phi^{\star}))$ and that $\left\|\mathcal{H}(\hat{\phi}^{-})\right\|\leq 2\left\|\mathcal{H}(\phi^{\star})\right\|$, we find that
-$\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})\right]\succeq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{16\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}I.$ (39)
-Denote the event that Algorithm 2 collects data satisfying (35) and (37) as $\mathcal{E}_{\mathsf{doed\,covariance}}$. As long as
-$\frac{N}{4}\geq\left(\frac{64RTL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))\lambda_{\min}^{\star}}\right)^{\frac{1}{\alpha}},$
-it follows that
-$\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)}{4TL_{f}^{2}}$ (40)
-from the bound $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$. Then by Lemma C.5,
-$\Sigma^{\pi_{\mathsf{mix}}}\succeq\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{-}))\lambda_{\min}^{\star}}{32\left\|\mathcal{H}(\hat{\phi}^{-})\right\|_{\mathsf{op}}d_{\phi}}I\succ 0.$
-Additionally, the policy $\pi_{\mathsf{mix}}$ is $(2C_{\mathsf{Loja}},\alpha)$-Lojasiewicz.
+Additionally, suppose $\gamma N\geq\left(\frac{6RT^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\min\mathopen{}\left\{\sigma_{w},1\right\}\mathclose{}\mu}\right)^{1/\alpha}$.
+Then by a first order Taylor expansion combined with Lemma B.4, we have that for any policy $\pi\in\Pi_{\exp}$ $\left\|\mathbf{FI}^{\pi}(\hat{\phi}^{-})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|\leq\frac{\mu}{2}$.
+As a result, by Assumption 2.6, there exists a policy $\pi$ such that $\lambda_{\min}(\mathbf{FI}^{\pi}(\hat{\phi}^{-}))\geq\frac{\mu}{2}$.
+Therefore, if $\pi_{\exp}$ satisfies $\operatorname{\mathrm{tr}}((\mathcal{H}(\hat{\phi}^{-})+\nu I)\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1})\leq(1+\varepsilon)\operatorname{\mathrm{tr}}((\mathcal{H}(\hat{\phi}^{-})+\nu I)\mathbf{FI}^{\tilde{\pi}}(\hat{\phi}^{-})^{-1})$ for all $\tilde{\pi}\in\Pi_{\exp}$, it holds that
+$\frac{\lambda_{\min}(\mathcal{H}(\hat{\phi}^{\star}))+\nu}{2\lambda_{\min}(\mathbf{FI}^{\pi_{\exp}}(\hat{\phi}^{-}))}\leq\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\nu\right)\mathbf{FI}^{\pi_{\exp}}(\hat{\phi}^{-})^{-1}\right)\leq\frac{(1+\varepsilon)4d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu},$
+and therefore that
+$\lambda_{\min}(\mathbf{FI}^{\pi_{\exp}}(\hat{\phi}^{-}))\geq\frac{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}{8(1+\varepsilon)d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}.$
+Suppose
+$\gamma N\geq\left(\frac{(1+\varepsilon)48RT^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\min\mathopen{}\left\{\sigma_{w},1\right\}\mathclose{}\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)\mu}\right)^{1/\alpha}.$
+Then by Lemma C.1, it follows that $\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})\succeq\frac{1}{2}\mathbf{FI}^{\pi_{\exp}}(\phi^{\star})\succeq\frac{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}{32(1+\varepsilon)d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}I$.
+Additionally, the policy $\pi_{\mathsf{mix}}$ is $(\frac{1}{\gamma}C_{\mathsf{Loja}},\alpha)$-Lojasiewicz.
 This can be seen by noting that by the law of total expectation.
 In particular, for all $\phi\in\mathbb{R}^{d_{\phi}}$,
-$\mathbf{Err}_{\pi_{\mathsf{mix}},\phi^{\star}}(\phi)\geq\frac{1}{2}\mathbf{Err}_{\pi^{0},\phi^{\star}}(\phi).$
-Then under the events $\mathcal{E}_{\mathsf{coarse}}$, $\mathcal{E}_{\mathsf{doed}}$ and $\mathcal{E}_{\mathsf{doed\,covariance}}$ in conjunction with the burn-in requirement
-$\frac{N}{4}\geq\max\mathopen{}\left\{\tau_{\mathbf{Err}}(\delta/32),\tau_{\mathsf{small\,error}}(\delta/16),\tau_{\mathsf{h.o.t}}(\delta/16),\tau_{\mathsf{cov\,conc}}(\delta/16)\right\}\mathclose{},$
-the conditions are set to apply Theorem 3.1 on the data collected by rerunning $\pi^{\mathsf{mix}}$ for $\frac{N}{4}$ episodes. Doing so ensures that with probability at least $41-\delta/$ 4,
-$\left\|\hat{\phi}^{-}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})}^{2}\leq 160\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}\frac{\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})}{N}.$ (41)
+$\mathbf{Err}_{\pi_{\mathsf{mix}},\phi^{\star}}(\phi)\geq\frac{1}{\gamma}\mathbf{Err}_{\pi^{0},\phi^{\star}}(\phi).$
+Then the slow rate from Corollary A.1 ensures that playing the policy $\pi^{\mathsf{mix}}$ for $(1-\gamma)N$ episodes results in a refined estimate that satisfies $\left\|\hat{\phi}^{+}-\phi^{\star}\right\|\leq\frac{R}{((1-\gamma)N)^{\alpha}}$.
+Additionally, under the event $\mathcal{E}_{\mathsf{coarse}}$ in conjunction with the burn-in requirement
+$(1-\gamma)N$ $\geq\mathsf{poly}_{\alpha}\bigg{(}T,L_{f},d_{\phi},d_{\mathsf{X}},\sigma_{w},\log N,\log\frac{1}{\delta},\log B,C_{\mathsf{Loja}},\frac{1}{\beta}\bigg{)},$
+the conditions are set to apply Theorem 3.1 on the data collected by running $\pi^{\mathsf{mix}}$ for $(1-\gamma)N$ episodes.
+To do so, we note that for any $\beta\in\left(0,\frac{\mu\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{512d_{\phi}(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu)}\right)$, the condition $\beta\leq\frac{\sigma_{w}^{2}\lambda_{\min}(\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})}{4}$ is met by the lower bound on $\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})$.
+Therefore, we may invoke Theorem 3.1 to find that with probability at least $21-\delta/$ 2,
+$\left\|\hat{\phi}^{+}-\phi^{\star}\right\|_{\mathcal{H}(\phi^{\star})+\nu I}^{2}$ $\leq 2\frac{(1+\xi)}{1-\gamma}\left(\frac{\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)}{N}+2\frac{\left\|\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|}{N}\log\frac{8}{\delta}\right)$ (31)
+$\leq\frac{(1+\xi)}{1-\gamma}\left(2+4\log\frac{8}{\delta}\right)\frac{\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)}{N},$
+where $\xi=4\beta\left(\frac{128d_{\phi}\left(\left\|\mathcal{H}(\phi^{\star})\right\|+\nu\right)}{\mu\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}+d_{\phi}\right)$.
 Denote this event as $\mathcal{E}_{\mathsf{ID}}$.
-Then under the event $\mathcal{E}_{\mathsf{coarse}}\cap\mathcal{E}_{\mathsf{doed\,conc}}\cap\mathcal{E}_{\mathsf{ID}}$, by using the covariance bound (39) and the estimation bound (40), it follows from Lemma C.5
-$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})$ $\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)^{-1}\right)$ (42)
-$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\left\|\hat{\phi}-\phi^{\star}\right\|.$
-By the lower bound (39), and the choice of $\lambda$ in 5 in Algorithm 1, it follows that
-$\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]\right)^{-1}\right)$ (43)
-$\leq 2\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi_{\mathsf{mix}},\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$
-$\leq 4\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi}^{-})+\lambda I\right)^{-1}\right)$ ((37) combined with (38))
-$\leq 4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$
-$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)$ ( (33) holds under the event $\mathcal{E}_{\mathsf{doed}}$)
-By invoking Lemma C.5,
-$\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi}^{-})\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi}^{-})\right]+\lambda I\right)^{-1}\right)$ (44)
-[…]
-$\leq\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star})}\right)\left\|\hat{\phi}-\phi^{\star}\right\|,$
-where the final inequality follows from Lemma C.5. Combining (42) with (43) and (44) and using the bound $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{N^{\alpha}}$ yields
-$\operatorname{\mathrm{tr}}(\mathcal{H}(\phi^{\star})(\Sigma^{\pi_{\mathsf{mix}}})^{-1})$ $\leq 4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)$ (45)
-$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{R}{N^{\alpha}}$
-$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)$
-As a result of the above identfiication bounds, the excess cost of the returned policy $\pi^{\star}(\hat{\phi}^{+})$ is bounded via Footnote 3 as
+It follows from Lemma C.1 that
+$(\displaystyle\operatorname{\mathrm{tr}}($$\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})\leq\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-})^{-1}\right)$ (32)
+$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\frac{1}{\mu},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)\left\|\hat{\phi}^{-}-\phi^{\star}\right\|.$
+By the fact that $\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-})\succeq(1-\gamma)\mathbf{FI}^{\pi_{\exp}}(\hat{\phi}^{-})$, and the fact that $\pi_{\exp}$ optimizes the approximate exploration objective, it holds that $\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-})^{-1}\right)\leq\frac{1+\varepsilon}{1-\gamma}\inf_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\right)$.
+Then by again appealing to Lemma C.1, it holds that
+$\inf_{\pi\in\Pi_{\exp}}$ $\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\hat{\phi}^{-})+\nu I\right)\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1}\right)\leq\inf_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)$ (33)
+$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\frac{1}{\mu},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)\left\|\hat{\phi}^{-}-\phi^{\star}\right\|.$
+Combining (32) with (33) and using the bound $\left\|\hat{\phi}^{-}-\phi^{\star}\right\|\leq\frac{R}{(\gamma N)^{\alpha}}$ yields
+$\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\phi^{\star})^{-1})\leq\frac{1+\varepsilon}{1-\gamma}\inf_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)$ (34)
+$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\frac{1}{\mu},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)\frac{R}{(\gamma N)^{\alpha}}.$
+As a result of the above identfiication bounds, the excess cost of the returned policy $\pi^{\star}(\hat{\phi}^{+})$ is bounded via Lemma 3.1 as
 $\mathcal{J}(\phi^{+})-\mathcal{J}(\phi^{\star})$ $\leq\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})}+C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{3}$
-$\leq\left(1+\frac{C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})}$
-[…]
-$\leq\left(1+\frac{2C_{\mathsf{cost}}R}{N^{\alpha}\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{160\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}}{N}\Bigg{(}4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)$
-$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)\frac{R}{N^{\alpha}}$
-$+\frac{8T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{4TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{\tilde{K}}{2\delta}}{\tilde{N}}}\right)\Bigg{)}.$
+$\leq\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})+\nu I}+C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{3}$
+$\leq\left(1+\frac{C_{\mathsf{cost}}\left\|\hat{\phi}^{+}-\phi^{\star}\right\|}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)\left\|\hat{\phi}^{+}-\phi^{\star}\right\|^{2}_{\mathcal{H}(\phi^{\star})+\nu I}$
+$\leq\left(1+\frac{C_{\mathsf{cost}}R}{((1-\gamma)N)^{\alpha}\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}\right)\frac{1+\xi}{1-\gamma}\left(2+4\log\frac{8}{\delta}\right)\frac{1}{N}\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi_{\mathsf{mix}}}(\hat{\phi}^{-})^{-1})$
+(Events $\mathcal{E}_{\mathsf{ID}}$ and the slow rate bound.)
+$\leq\left(1+\frac{C_{\mathsf{cost}}R}{((1-\gamma)N)^{\alpha}\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}\right)\frac{\frac{(1+\varepsilon)(1+\xi)}{(1-\gamma)^{2}}\left(2+4\log\frac{8}{\delta}\right)}{N}\Bigg{(}\inf_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\hat{\phi}^{-})^{-1})$
+$+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\frac{1}{\mu},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)\frac{R}{(\gamma N)^{\alpha}}\Bigg{)}.$
 To conclude, we leverage the burn-in condition
-$\frac{N}{4}\geq\max\Bigg{\{}\Bigg{\lceil}\left(\frac{4T^{4}L_{f}^{8}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\log(N+1)\lambda^{-3}+2TL_{f}^{2}\left\|\mathcal{H}(\phi^{\star})\right\|_{\mathsf{op}}\lambda^{-2}\left(2T^{2}L_{f}R\sigma_{w}^{-1}+\sqrt{8\log\frac{2N}{\delta}}\right)+C_{\mathsf{H\,pert}}R}{4\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)}\right)\bigg{\rceil}^{3/\alpha},$
-$\left(2\frac{C_{\mathsf{cost}}R}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)^{1/\alpha}\Bigg{\}},$
+$N\geq\max\Bigg{\{}\left(\frac{C_{\mathsf{H\,pert}}R}{\beta\inf_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu I\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)}\right)^{1/\alpha},\left(\frac{C_{\mathsf{cost}}R}{\beta(1-\gamma)\left(\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu\right)}\right)^{1/\alpha}\Bigg{\}},$
 where
-$C_{\mathsf{H\,pert}}=\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\lambda_{\min}^{\star}},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))}\right)$
+$C_{\mathsf{H\,pert}}=\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},\sigma_{w},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\nu,\frac{1}{\mu},\frac{1}{\lambda_{\min}(\mathcal{H}(\phi^{\star}))+\nu}\right)$
 Under this condition, the excess cost bound simplifies to
-$\mathcal{J}(\phi^{+})-\mathcal{J}(\phi^{\star})$ $\leq\frac{2560\sigma_{w}^{2}\log\frac{92d_{\phi}}{\delta}}{N}\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right).$
-Appendix E Aditional Experiments and Experiment Details
-In Section E.1, we complement our illustrative example with an implementation of ALCOI on a toy physical system. In Section E.2 we provide further experimental details.
-E.1 Cartpole Experiments
+$\mathcal{J}(\phi^{+})-\mathcal{J}(\phi^{\star})$ $\leq\frac{\frac{(1+\beta)^{2}(1+\varepsilon)(1+\xi)}{(1-\gamma)^{2}}(2+4\log\frac{8}{\delta})}{N}\min_{\pi\in\Pi_{\exp}}\operatorname{\mathrm{tr}}\left(\left(\mathcal{H}(\phi^{\star})+\nu\right)\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right).$
+Appendix D Aditional Experiments and Experiment Details
+In Section D.1, we complement our illustrative example with an implementation of ALCOI on a toy physical system. In Section D.2 we provide further experimental details.
+D.1 Cartpole Experiments
 We additionally run our proposed algorithm on the cartpole system defined by the dynamics
@@
 Figure 3: Excess cost versus number of exploration episodes for the proposed control-oriented identification procedure, approximate $A$-optimal design, and random exploration for the cartpole swing-up task.
-The mean over $100$ runs is shown, and the standard error is shaded.
-E.2 Implementation Details
+The mean over $900$ runs is shown, and the standard error is shaded.
+D.2 Implementation Details
 For both the cartpole experiments and the illustrative example, the initial exploration policy used by ALCOI and approximate $A$-optimal exploration is set to the random exploration policy.
 We deviate from the proposed algorithm in the following respects.
-  • Rather than running the initial policy for $\lfloor N/4\rfloor$ episodes and the algorithm $\texttt{DOED}_{+}$ with a budget of $\lfloor N/4\rfloor$ before running the mixture policy for $\lfloor N/4\rfloor$ episodes, we run the initial policy for $\lfloor N/2\rfloor$ episodes and the algorithm $\texttt{DOED}_{+}$ with a budget of $\lfloor N/2\rfloor$ episodes, and use all of the collected data to fit the refined estimate.
-  • We treat the regularization parameter $\lambda$ as a hyperparameter, which we fix to $0.001$ in all experiments, rather than setting it according to 5 from Algorithm 1.
-  • The approach of optimism in the face of uncertainty from 8 of Algorithm 2 is replace by certainty equivalent synthesis of the exploration policy.
+  • We use all of the collected data to fit the refined estimate.
   • As the policy class in both experiments consists of all trajectory dependent policies with bounded input energy, we approximate this certainty equivalent synthesis via a receding horizon control procedure.
@@
     Note that a consequence of this is that there may exist policies in the policy class $\Pi_{\star}$ which achieve lower cost than the policy defined in terms of the optimal parameter $\phi_{\star}$.
-The first modification is made because discarding the half of the data is wasteful.
-It is included in the algorithm only in order to simplify the analysis, as the constant factor of $2$ is not a concern of this analysis.
-The second modification is made to avoid requiring knowledge of the parameter $\lambda_{\min}^{\star}$ by the algorithm.
-The third modification is made to avoid the computational intractability of OFU.
+The first modification is made because discarding the a portion of the data is wasteful.
+It is included in the algorithm only in order to simplify the analysis, as the constant factor of $(1-\beta)$ is not a concern of this analysis.
 The sampling based approach is used since the optimization problem defining the receding horizon controller is intractable.
 The final modification is made to avoid the computational burden of solving the policy optimization problem via a Bellman equation in both policy synthesis and in computation of the model-task Hessian.
-These modifications result in a gap between the theoretical guarantees, and the numerical validation.
-Future work could reduce this gap, e.g.
-by removing the approach of optimism in the face of uncertainty from the analysis.
+These modifications result in a small gap between the theoretical guarantees, and the numerical validation, which future work will remove.